### v1

# Decoding Futures Price Dynamics: A Regularized Sparse Autoencoder for Interpretable Multi-Horizon Forecasting and Factor Discovery

Abstract
Commodity price volatility creates economic challenges, necessitating accurate multi-horizon forecasting. Predicting prices for commodities like copper and crude oil is complicated by diverse interacting factors (macroeconomic, supply/demand, geopolitical, etc.). Current models often lack transparency, limiting strategic use. This paper presents a Regularized Sparse Autoencoder (RSAE), a deep learning framework for simultaneous multi-horizon commodity price prediction and discovery of interpretable latent market drivers. The RSAE forecasts prices at multiple horizons (e.g., 1-day, 1-week, 1-month) using multivariate time series. Crucially, L1 regularization ($\|\mathbf{z}\|_{1}$) on its latent vector $\mathbf{z}$ enforces sparsity, promoting parsimonious explanations of market dynamics through learned factors representing underlying drivers (e.g., demand, supply shocks). Drawing from energy-based models and sparse coding, the RSAE optimizes predictive accuracy while learning sparse representations. Evaluated on historical Copper and Crude Oil futures data with numerous indicators, our findings indicate the RSAE offers competitive multi-horizon forecasting accuracy and data-driven insights into price dynamics via its interpretable latent space, a key advantage over traditional black-box approaches.
1 Introduction
Commodity markets, and their associated futures contracts, form a critical part of the global economy, providing essential raw materials and hedging instruments. The prices of key commodity futures, such as those for industrial metals like Copper and energy benchmarks like Crude Oil, are not merely transactional figures; they are vital indicators of economic health, drivers of industrial costs, and significant factors influencing inflation, trade balances, and geopolitical dynamics. Copper futures, for instance, are closely watched for signals about global industrial demand. Crude Oil futures are pivotal for managing energy price risk and reflect complex geopolitical and supply/demand interactions.
The inherent volatility of these commodity futures prices presents substantial risks and opportunities for a wide spectrum of market participants. Producers and consumers use futures to hedge price uncertainty, while financial institutions and traders engage in speculative and arbitrage activities. Consequently, the ability to accurately forecast commodity futures prices across various time horizons is of paramount importance for effective risk mitigation, optimized resource allocation, and informed investment decisions.
The challenge lies in the sheer complexity of factors driving futures prices. These markets are influenced by a confluence of interconnected forces:
  1. 1.
    Macroeconomic Environment: Global GDP growth, industrial production, interest rates, currency exchange rates (particularly the US Dollar), and inflation expectations.
  2. 2.
    Underlying Physical Market Fundamentals: For Copper, this includes mine production, scrap availability, and inventory levels. For Crude Oil, factors include OPEC+ decisions, non-OPEC output, geopolitical events in producing regions, and inventory levels.
  3. 3.
    Demand Fundamentals: Linked to manufacturing activity, construction, transportation needs, and energy transition dynamics.
  4. 4.
    Market Sentiment and Financialization: Speculative positioning in futures and options markets (e.g., CFTC Commitment of Traders reports), algorithmic trading, and broader risk-on/risk-off sentiment.
  5. 5.
    Other factors: Including weather patterns impacting supply/demand and inter-commodity relationships.
This intricate web of high-dimensional, often collinear, and dynamically interacting drivers makes futures price forecasting an exceptionally challenging task, characterized by non-linear relationships and regime shifts. Existing approaches often face a trade-off between predictive accuracy and interpretability. Classical time series models (ARIMA, GARCH) may struggle with non-linearities and exogenous factors. Econometric structural models rely on strong assumptions. While traditional machine learning (SVM, Random Forests) and standard deep learning models (LSTMs, CNNs) can capture complex patterns, they often function as ”black boxes,” limiting actionable insights [45]. There is a need for models that not only forecast accurately but also provide an understanding of the underlying drivers.
Given these limitations, this research addresses the question: How can we develop a deep learning framework that achieves high predictive accuracy for commodity futures prices across multiple horizons and simultaneously extracts a sparse, interpretable set of latent factors representing key market drivers? We propose a Regularized Sparse Autoencoder (RSAE) framework that utilizes an autoencoder backbone [96] with a time-lagged predictive objective and L1 sparsity regularization [94] on its latent code. This design, inspired by energy-based models [59] and sparse coding [72], forces the model to explain market dynamics through a parsimonious set of learned factors.
Our contribution lies in the specific design and application of this RSAE framework to multi-horizon commodity futures forecasting, aiming to enhance both accuracy and transparency. The remainder of this paper details the related work, the RSAE methodology, data, experimental setup, results, and discussion.
2 Review of Related Work in Commodity Forecasting
2.1 Introduction to Commodity Price Forecasting: Significance and Challenges
Accurate forecasting of commodity prices is crucial, given their extensive impact on economies and daily life [1]. Variations in prices of essential goods—from agricultural produce to energy resources and industrial metals—affect many aspects of economic activities and individual well-being. For instance, volatile prices can significantly influence agricultural productivity and farmer incomes [35], while also presenting unpredictable expenses for consumers. Countries heavily dependent on commodity trade find their macroeconomic stability closely linked to commodity price fluctuations [18]. Additionally, amid rising concerns about global food security, effective prediction of staple commodity prices becomes vital to managing risks and preventing shortages [31].
Reliable commodity price forecasts are essential for various stakeholders. Producers depend on price predictions to make informed planting and harvesting decisions. Importers and exporters utilize forecasts to negotiate trade agreements and manage price risks effectively. Traders and investors in financial markets rely heavily on accurate price forecasts for speculative investments and portfolio management, viewing commodities as valuable assets that can diversify risk and potentially offer stability. Governments and policymakers require precise forecasts for macroeconomic planning, managing essential goods’ supply-demand dynamics, making informed decisions about agricultural support programs, and developing sound economic policies. This broad reliance highlights the systemic importance of commodity price fluctuations; inaccuracies in forecasting can have widespread economic consequences, impacting everything from consumer purchasing power to national economic performance and international trade dynamics.
Despite its significance, commodity price forecasting remains notably challenging. Commodity markets exhibit inherent complexity and volatility, characterized by nonlinear relationships and nonstationary trends [42]. Prices respond dynamically to a broad range of interrelated factors including fundamental supply-demand variables (production levels, crop yields, and inventory), macroeconomic indicators (inflation rates, industrial output, interest rates), financial market influences (currency exchange rates, investor behavior, volatility measures), geopolitical incidents, trade disruptions, and environmental variables like extreme weather events and climate change, especially affecting agricultural commodities. Furthermore, these influencing factors constantly evolve. Increased financial market involvement in commodities, frequent global crises (economic downturns, pandemics), erratic climate patterns, and shifting geopolitical conditions all add further complexity and unpredictability. The growing intricacy of these factors intensifies volatility, rendering forecasting increasingly challenging and demanding adaptive models capable of dynamically integrating new information and capturing evolving relationships. Consequently, traditional forecasting methods often fall short, necessitating exploration into more advanced modeling approaches.
2.2 Traditional Econometric Approaches: Foundations and Limitations
The quantitative forecasting of commodity prices has historically been dominated by traditional econometric and time series models. These methods provided the foundational tools for analyzing price dynamics and generating predictions based on historical data. Prominent among these are linear regression models, Autoregressive Integrated Moving Average (ARIMA) models, Vector Autoregression (VAR) models, Vector Error Correction Models (VECM) for cointegrated series, and Random Walk models. ARIMA models, in particular, gained widespread popularity over several decades, valued for their well-defined statistical properties and the systematic identification procedure proposed by Box and Jenkins [13]. Studies have demonstrated ARIMA’s utility in specific contexts, such as forecasting lead and zinc prices or certain agricultural commodity prices under specific conditions [68].
Recognizing the pervasive issue of time-varying volatility (i.e., periods of high volatility followed by periods of relative calm, known as volatility clustering) in financial and commodity markets, Generalized Autoregressive Conditional Heteroskedasticity (GARCH) models were developed [12, 29]. GARCH and its numerous extensions specifically aim to model and forecast the conditional variance of price returns. This focus on volatility acknowledges that risk assessment and management are critical aspects of commodity market analysis.
Despite their established role, traditional econometric models face significant limitations when applied to the complexities of modern commodity markets, primarily their reliance on restrictive assumptions like linearity and stationarity [90]. Commodity price series frequently exhibit strong nonlinearity and structural breaks [76], violating these core assumptions. While some studies report adequate performance for traditional models in specific scenarios [10], a growing body of evidence highlights their inadequacy for many commodity prices. The recognized shortcomings of these models served as a catalyst for exploring more flexible modeling paradigms [77, 56].
2.3 Rise of Machine Learning in Commodity Forecasting
In response to the limitations of traditional models, the field has seen a significant shift towards machine learning (ML) and artificial intelligence (AI) techniques [19, 85]. ML models can model complex, nonlinear relationships directly from data without restrictive assumptions, often leading to improved forecasting accuracy compared to traditional benchmarks [69]. ML also facilitates analysis of diverse datasets, including unstructured data via NLP [99].
This reflects a potential paradigm shift from theory-driven econometrics to data-driven ML [16]. However, ML models, especially deep learning, often require large datasets [37], which can be a constraint in some commodity markets.
2.4 Key Machine Learning Techniques Applied
A spectrum of ML algorithms has been applied, as summarized in Table 1.
Table 1: Overview of Machine Learning Models in Commodity Forecasting.
Model Type Core Principle Strengths in Commodity Forecasting Weaknesses / Challenges
Support Vector Machine (SVM) / Support Vector Regression (SVR) Finds optimal hyperplane to separate/regress data, using kernel tricks for non-linearity. Handles non-linearity effectively, robust baseline performance, good generalization. Sensitive to parameter/kernel choice, can be computationally intensive for large datasets.
Tree-based Ensembles (e.g., Random Forest (RF), Gradient Boosting) Combines multiple decision trees to improve prediction accuracy and robustness. Handles non-linearity, captures interactions, provides feature importance, robust to outliers. Can be prone to overfitting if not properly tuned, less interpretable than single trees.
Artificial Neural Network (ANN) / Multilayer Perceptron (MLP) Learns complex non-linear mappings through interconnected layers of neurons. Universal approximators, capture non-linear patterns. Prone to local optima, requires careful tuning, can overfit, may struggle with long-term dependencies.
Long Short-Term Memory (LSTM) Type of RNN with gating mechanisms to control information flow and capture long-range dependencies. Excels at modeling sequential data and long-term dependencies. Computationally expensive, requires significant data, complex architecture.
Gated Recurrent Unit (GRU) Simplified version of LSTM with fewer gates. Captures long-term dependencies like LSTM, potentially more computationally efficient. Performance relative to LSTM is context-dependent.
Transformer Uses self-attention mechanisms to weigh input sequence importance, parallelizable. Excellent at capturing long-range dependencies. Computationally very expensive, requires large datasets.
Note: Strengths, weaknesses, and representative studies are derived from the provided research material.
Support Vector Regression (SVR) [27] and tree-based ensembles like Random Forests (RF) [15] and XGBoost [20] are foundational. SVR handles nonlinearities well [93, 50, 74, 73]. Tree-based methods are robust and can model interactions [92].
Deep learning, starting with ANNs/MLPs [84, 102], evolved to Recurrent Neural Networks (RNNs) [28] for sequential data. LSTMs [46] specifically address RNNs’ vanishing gradient problem [8], making them effective for long-term dependencies in commodity forecasting [30, 103, 88]. Architectural variants like bidirectional LSTMs [86] and CNN-LSTM hybrids [53, 65] exist. GRUs [21] offer a simpler alternative, with performance varying by context [100, 22]. CNNs [58, 7] also find use, often in hybrids. The choice often involves trade-offs between complexity, data needs, and interpretability [44].
Transformer models [95], with their self-attention mechanism, represent a newer frontier, excelling at long-range dependencies [104, 62, 61, 64]. They are often combined with decomposition techniques in hybrid models [97]. The development of very large models (LLMs) also influences this space [17, 98].
Hybrid models combining different techniques are common [101, 52]. A prevalent strategy is ”decomposition and ensemble” [57], using methods like Wavelet Transforms [24], VMD [26], EMD [47], SSA [36], or STL [23] to simplify the series before forecasting each component [6, 40, 43, 60]. Explainable AI (XAI) is also being integrated [75].
Performance evaluation uses metrics like RMSE, MAE, and directional accuracy [48], sometimes with statistical tests like Diebold-Mariano [25]. ML/DL models often outperform traditional ones, but rankings vary by context [91, 49].
3 Methodology: The Regularized Sparse Autoencoder (RSAE) Framework
This section details the Regularized Sparse Autoencoder (RSAE) framework, an approach specifically designed for the dual task of multi-horizon commodity futures price prediction and the discovery of interpretable latent market factors. The methodology draws inspiration from energy-based models [59] and sparse coding [72], employing an explicit inference step to determine a sparse latent representation that best explains future price movements.
3.1 Problem Formulation
We consider a set of $d$ market features (e.g., macroeconomic, supply/demand, technical indicators) at time $t$, denoted by $\mathbf{x}_{t}\in\mathbb{R}^{d}$. The model input consists of a look-back window of $W$ historical observations:
$X=X_{t-W+1:t}=(\mathbf{x}_{t-W+1},\ldots,\mathbf{x}_{t})\in\mathbb{R}^{W\times d}$
Our objective is to predict the commodity futures price $p_{t}$ at $N$ distinct future horizons, where $\tau_{1}<\tau_{2}<\ldots<\tau_{N}$. The corresponding target vector is:
$Y=Y_{t}=(p_{t+\tau_{1}},p_{t+\tau_{2}},\ldots,p_{t+\tau_{N}})^{\top}\in\mathbb{R}^{N}$
The RSAE model aims to learn a mapping from $X$ to $Y$ via an intermediate sparse latent variable $\mathbf{z}\in\mathbb{R}^{m}$, where $m$ is the dimensionality of the latent space and $m\ll W\times d$. This latent variable $\mathbf{z}$ is intended to capture the essential underlying market state relevant for forecasting future price evolution.
3.2 Model Architecture
The RSAE architecture comprises three distinct neural network modules interacting within an energy-based learning paradigm. (A conceptual diagram of the architecture is presented in Figure 1.)
Figure 1: Conceptual architecture of the Regularized Sparse Autoencoder (RSAE) framework, illustrating the interaction between the History Predictor (Pred), Decoder (Dec), and Encoder (Enc) modules, along with the inference process for the sparse latent variable $\mathbf{z}$.
  1. 1.
    History Predictor (Pred): This module, parameterized by $\theta_{\text{Pred}}$, processes the historical input window $X$ to generate a condensed summary or context vector $\bm{h}\in\mathbb{R}^{h_{\text{dim}}}$. Given the sequential nature of $X$, this module typically utilizes recurrent layers (e.g., LSTMs [46], GRUs [21]) or temporal convolutional networks (1D CNNs) [7].
    $\bm{h}=\text{Pred}(X;\theta_{\text{Pred}})$
  2. 2.
    Decoder (Dec): This module, parameterized by $\theta_{\text{Dec}}$, takes both the history context $\bm{h}$ and a specific latent variable $\bm{z}\in\mathbb{R}^{m}$ as input to predict the multi-horizon future price vector $\tilde{Y}\in\mathbb{R}^{N}$.
    $\tilde{Y}=\text{Dec}(\bm{z},\bm{h};\theta_{\text{Dec}})$
    The latent variable $\bm{z}$ acts as a representation of the underlying factors modulating the future trajectory predicted from the historical context $\bm{h}$.
  3. 3.
    Encoder (Enc): This module, parameterized by $\theta_{\text{Enc}}$, performs amortized inference [33, 55]. Its role is to learn to directly predict the optimal latent variable $\bm{z}^{*}$ (derived via the inference process detailed below) given the actual future outcome $Y$ and the history context $\bm{h}$.
    $\hat{\bm{z}}=\text{Enc}(Y,\bm{h};\theta_{\text{Enc}})$
    where $\hat{\bm{z}}$ is the Encoder’s estimate of $\bm{z}^{*}$.
3.3 Latent Variable Framework: Energy Function and Inference
Central to the RSAE is the determination of the latent variable $\bm{z}$. We define an energy function $E(Y,\bm{h},\bm{z})$ that quantifies the compatibility between a latent code $\bm{z}$, the historical context $\bm{h}$, and the future outcome $Y$. Low energy corresponds to a sparse latent code $\bm{z}$ that enables the Decoder to accurately predict $Y$ from $\bm{h}$ [59].
Energy Function
$E(Y,\bm{h},\bm{z})=C(Y,\text{Dec}(\bm{z},\bm{h};\theta_{\text{Dec}}))+\lambda R(\bm{z})$
where:
  • $C(Y,\tilde{Y})$ is the prediction cost, measuring the error between the true future $Y$ and the prediction $\tilde{Y}=\text{Dec}(\bm{z},\bm{h})$. We typically use Mean Squared Error (MSE) [11]:
    $C(Y,\tilde{Y})=\frac{1}{N}\left\|Y-\tilde{Y}\right\|_{2}^{2}$
  • $R(\bm{z})$ is the L1-norm sparsity regularizer [94]:
    $R(\bm{z})=\|\bm{z}\|_{1}=\sum_{k=1}^{m}|z_{k}|$
  • $\lambda>0$ is the hyperparameter controlling the strength of the sparsity penalty.
Optimal Latent Variable ($\bm{z}^{*}$)
For a given input pair $(X,Y)$ yielding $(\bm{h},Y)$, the optimal latent variable $\bm{z}^{*}$ is the value of $\bm{z}$ that minimizes the energy function:
$\bm{z}^{*}=\arg\min_{\bm{z}\in\mathbb{R}^{m}}E(Y,\bm{h},\bm{z})$
This minimization problem is typically non-convex if the decoder is non-linear.
Inference via Optimization
Finding $\bm{z}^{*}$ involves solving this optimization problem. We employ an iterative approach using gradient descent (or a subgradient method due to the L1 term) for a fixed number of steps $K$. This constitutes the explicit inference stage within the training loop:
  1. 1.
    Initialize $\bm{z}^{(0)}$ (e.g., a zero vector or from the Encoder).
  2. 2.
    For $k=0$ to $K-1$:
    • Compute the (sub)gradient of the energy $E$ with respect to $\bm{z}$ at $\bm{z}^{(k)}$, holding $\bm{h}$ and $\theta_{\text{Dec}}$ constant [14]:
      $\nabla_{\bm{z}}E(\cdot,\cdot,\bm{z}^{(k)})=\frac{\partial C}{\partial\text{Dec}}\cdot\frac{\partial\text{Dec}(\bm{z}^{(k)},\bm{h})}{\partial\bm{z}}+\lambda\,\text{sign}(\bm{z}^{(k)})$
    • Update $\bm{z}$ using a step size $\alpha_{\text{inf}}$:
      $\bm{z}^{(k+1)}\leftarrow\bm{z}^{(k)}-\alpha_{\text{inf}}\nabla_{\bm{z}}E(\cdot,\cdot,\bm{z}^{(k)})$
  3. 3.
    The result $\bm{z}^{(K)}$ serves as the inferred optimal latent variable for this sample: $\bm{z}^{*}\approx\bm{z}^{(K)}$.
This inference process finds a sparse latent code $\bm{z}^{*}$ under the current model parameters that aims to best explain the observed future $Y$ given the history $\bm{h}$.
3.4 Amortized Inference via Encoder (Enc)
The $K$-step inference process, while crucial for finding an optimal latent variable $\bm{z}^{*}$, introduces computational overhead during training and especially at test time if performed for every sample. To address this, the Encoder module $\text{Enc}(Y,\bm{h};\theta_{\text{Enc}})$ provides amortized inference by learning to directly map the observed future $Y$ and historical context $\bm{h}$ to an estimate $\hat{\bm{z}}$ of $\bm{z}^{*}$ [33].
$\hat{\bm{z}}=\text{Enc}(Y,\bm{h};\theta_{\text{Enc}})$
This learned mapping enables faster inference once the model is trained. During training, the explicit $K$-step inference process is used to generate the target $\bm{z}^{*}$ for supervising the Encoder’s learning. At test time, for forecasting, the Encoder cannot use the future $Y$; instead, an alternative encoder $\text{Enc}_{\text{test}}(X;\theta_{\text{EncTest}})$ might be trained to predict $\bm{z}^{*}$ from $X$ only, or the $K$-step inference is run starting from $\bm{z}^{(0)}=0$. This paper uses the latter approach for generating $\bm{z}^{*}$ at test time.
3.5 Training Objective and Parameter Updates
The training process aims to minimize errors in both prediction (using the inferred latent variable $\bm{z}^{*}$) and in the amortized inference of $\bm{z}^{*}$ by the Encoder (if an Encoder for $\bm{z}^{*}$ from $X$ and $Y$ is used for training efficiency). The parameters $\theta_{\text{Pred}},\theta_{\text{Dec}},\theta_{\text{Enc}}$ are updated based on two primary loss components, calculated after obtaining $\bm{h}=\text{Pred}(X)$, inferring $\bm{z}^{*}$ via $K$-step optimization, and, if applicable, computing $\hat{\bm{z}}=\text{Enc}(Y,\bm{h})$:
  1. 1.
    Prediction Loss ($L_{\text{Pred}}^{*}$): Measures the error of the Decoder when using the inferred optimal latent variable $\bm{z}^{*}$.
    $L_{\text{Pred}}^{*}=C(Y,\text{Dec}(\bm{z}^{*},\bm{h};\theta_{\text{Dec}}))$
  2. 2.
    Inference Loss ($L_{\text{Inf}}$) (Optional, for training the Encoder Enc): Measures the discrepancy between the optimal latent variable $\bm{z}^{*}$ (from K-step inference) and the training-time Encoder’s prediction $\hat{\bm{z}}$.
    $L_{\text{Inf}}=D(\bm{z}^{*},\hat{\bm{z}})=\left\|\bm{z}^{*}-\text{Enc}(Y,\bm{h};\theta_{\text{Enc}})\right\|_{2}^{2}$
The gradients for updating the network parameters are derived from these losses:
  • $\theta_{\text{Dec}}$ is updated based on $\nabla_{\theta_{\text{Dec}}}L_{\text{Pred}}^{*}$.
  • $\theta_{\text{Enc}}$ (if used) is updated based on $\nabla_{\theta_{\text{Enc}}}L_{\text{Inf}}$.
  • $\theta_{\text{Pred}}$ is updated based on gradients flowing back from $L_{\text{Pred}}^{*}$ (via $\bm{h}$) and potentially $L_{\text{Inf}}$ (via $\bm{h}$, if Enc is used). The dependency of $\bm{z}^{*}$ on $\bm{h}$ and $\theta_{\text{Dec}}$ can be handled by treating $\bm{z}^{*}$ as fixed post-inference for simplicity, or through more complex methods like implicit differentiation [79] or unrolling the inference loop [39]. In this work, $\bm{z}^{*}$ is treated as fixed after its $K$-step inference for updating $\theta_{\text{Pred}}$ and $\theta_{\text{Dec}}$.
The sparsity penalty $\lambda\|\bm{z}\|_{1}$ directly shapes the target $\bm{z}^{*}$ during its $K$-step inference, thereby influencing all subsequent parameter updates.
3.6 Overall Training Algorithm
The training follows a mini-batch stochastic gradient descent procedure:
  1. 1.
    Initialization: Set up data loaders, initialize network parameters $(\theta_{\text{Pred}},\theta_{\text{Dec}})$, and choose hyperparameters ($W$, $m$, $N$, $\tau_{i}$, $\lambda$, $K$, $\alpha_{\text{inf}}$, batch size $B$, optimizer learning rate $\eta$). The Encoder $\theta_{\text{Enc}}$ is optional for the core prediction task if inference is always done via K-step optimization.
  2. 2.
    Epoch Loop: For each mini-batch $\{(X^{(j)},Y^{(j)})\}_{j=1}^{B}$:
    1. (a)
      Compute history encoding: $\bm{h}^{(j)}=\text{Pred}(X^{(j)};\theta_{\text{Pred}})$
    2. (b)
      Infer optimal latent variable via $K$-step gradient descent:
      $\bm{z}^{*(j)}=\arg\min_{\bm{z}}E(Y^{(j)},\bm{h}^{(j)},\bm{z})$
      (using current $\theta_{\text{Dec}}$)
    3. (c)
      Calculate per-sample prediction loss: $L_{\text{Pred}}^{*(j)}=C(Y^{(j)},\text{Dec}(\bm{z}^{*(j)},\bm{h}^{(j)};\theta_{\text{Dec}}))$
    4. (d)
      Compute average batch prediction loss: $L_{\text{Pred}}^{*}=\frac{1}{B}\sum_{j=1}^{B}L_{\text{Pred}}^{*(j)}$
    5. (e)
      Compute gradients of $L_{\text{Pred}}^{*}$ with respect to $\theta_{\text{Pred}},\theta_{\text{Dec}}$.
    6. (f)
      Update parameters $\theta_{\text{Pred}},\theta_{\text{Dec}}$ using the chosen optimizer (e.g., Adam [54]).
  3. 3.
    Validation & Tuning: Periodically evaluate $L_{\text{Pred}}^{*}$ (using $\bm{z}^{*}$ inferred via $K$-step optimization at validation time) and sparsity metrics (e.g., $\|\bm{z}^{*}\|_{0}$) on a validation set. Use results to guide hyperparameter tuning (especially $\lambda$, $K$, $m$) and for early stopping [78].
  4. 4.
    Testing & Analysis: Evaluate the final trained model on a held-out test set using $L_{\text{Pred}}^{*}$ as the primary performance metric (with $\bm{z}^{*}$ inferred via $K$-step optimization). Analyze the inferred latent variables $\bm{z}^{*}$ from the test set—examining their sparsity, correlations with input features, and temporal evolution—to assess interpretability.
3.7 Interpretability via Sparsity
The explicit use of the L1 norm $R(\bm{z})=\|\bm{z}\|_{1}$ in the energy function plays a critical role in achieving interpretability [94]. By encouraging components of the optimal latent variable $\bm{z}^{*}$ to be exactly zero, the model is forced to represent the underlying market dynamics responsible for predicted price changes using a minimal set of active latent factors. Analyzing which factors are active can provide insights into the model’s reasoning. The hyperparameter $\lambda$ directly controls this sparsity.
4 Data Description and Pre-processing
4.1 Data Sources
Data was aggregated from multiple publicly accessible or commonly subscribed financial data provider sources:
  • Financial Information Providers: Services like Refinitiv Eikon or Bloomberg (via institutional access), and publicly available APIs from exchanges (e.g., CME Group, ICE) or financial data aggregators (e.g., Quandl - now part of Nasdaq Data Link, Yahoo Finance) — for daily market prices, futures data, and economic indicators.
  • Commodity Exchanges: Direct data feeds or historical data services from London Metal Exchange (LME), CME Group (NYMEX), and Intercontinental Exchange (ICE).
  • Government Agencies: U.S. Energy Information Administration (EIA), Bureau of Labor Statistics (BLS), Federal Reserve Economic Data (FRED), and National Bureau of Statistics of China.
  • International Organizations: International Monetary Fund (IMF), World Bank, Organisation for Economic Co-operation and Development (OECD).
  • Specialized Data Providers: Baker Hughes (rig counts via public releases), Commodity Futures Trading Commission (CFTC) (Commitment of Traders reports), and NOAA/NWS or equivalent national/international meteorological services (for historical weather data).
While the primary price data for model training is daily, several fundamental indicators are incorporated at lower frequencies. Table 2 illustrates sources for some representative monthly variables.
Table 2: Illustrative sources for some monthly input variables used in this research.
Variables Period Unit Data source Data availability
Crude Oil Price (benchmark, e.g., Brent) Monthly \$/barrel U.S. Energy Information Administration
(https://www.eia.gov/dnav/pet/hist/RBRTED.htm)
Freely available
Copper Price (benchmark) Monthly \$/ton World Bank Commodity Price Data (Pink Sheet)
(https://www.worldbank.org/en/research/commodity-markets)
Freely available
Copper Supply, Demand Indicators Monthly Various U.S. Geological Survey (USGS) Mineral Commodity Summaries
(https://www.usgs.gov/centers/national-minerals-information-center/mineral-commodity-summaries)
Freely available
Crude Oil Supply and Demand Indicators Monthly Various U.S. Energy Information Administration (EIA) Short-Term Energy Outlook (STEO) Freely available
4.2 Feature Selection
A wide array of features were selected, categorized as:
4.2.1 Target Variable (Daily Futures Prices)
  • Copper: LME Official Cash Settlement Price (or nearby futures contract).
  • Crude Oil: WTI front-month futures price (NYMEX); Brent Crude front-month futures (ICE).
4.2.2 Lagged Target Variables & Price Dynamics (Daily)
Historical prices, log returns, realized volatility [4], futures curve spreads [38].
4.2.3 Macroeconomic Indicators (Daily, Monthly, Quarterly)
Global/Major Economy PMIs, Industrial Production, inflation (CPI, PPI), interest rates (central bank rates, sovereign yields like US 10-Year Treasury), currency rates (DXY, EUR/USD, USD/CNY).
4.2.4 Supply/Demand Fundamentals (Varying Frequencies)
Copper: Exchange inventories (LME, COMEX, SHFE), publicly reported global/major producer output, China’s consumption/imports, scrap indicators (where publicly accessible). Crude Oil: EIA weekly inventories, US production, publicly available OPEC+ figures, Baker Hughes rig counts, IEA/OPEC demand estimates.
4.2.5 Market Sentiment & Financial Data (Daily, Weekly)
Futures volume/open interest, CFTC COT reports, VIX, OVX, major equity indices (S&P 500, FTSE 100, Shanghai Composite).
4.2.6 Technical Indicators (Calculated from Daily Prices)
Moving Averages (SMA, EMA), Oscillators (RSI, MACD), ADX [70].
4.2.7 Other (Exploratory)
Inter-commodity spreads, relevant weather indices (e.g., ENSO, hurricane impact assessments if data accessible).
4.3 Data Frequency and Time Period
Primary modeling frequency is daily. Lower frequency data (weekly, monthly, quarterly) are forward-filled, strictly aligning with information release dates to prevent lookahead bias [34]. Particular care is taken with lagged official statistics (e.g., STEO) to ensure only historically available information is used at each time step $t$. The dataset spans January 1, 2005 – December 31, 2023.
4.4 Data Pre-processing
Standard steps are applied [37]:
  1. 1.
    Missing Data: Features with $\geq$40missing values removed; remaining gaps primarily forward-filled.
  2. 2.
    Feature Transformation: Log returns or percentage changes for non-stationary series. Log transformation for skewed positive features.
  3. 3.
    Normalization/Standardization: Z-score normalization, with parameters derived only from the training set.
  4. 4.
    Outlier Treatment: Applied cautiously (e.g., winsorization at 1st/99th percentiles) to preserve genuine volatility.
4.5 Final Dataset Construction
Input sequences $X_{t}\in\mathbb{R}^{W\times d}$ (look-back $W=60$ days) and targets $Y_{t}\in\mathbb{R}^{N}$ (log prices at $N=3$ horizons: $\tau_{1}=1,\tau_{2}=5,\tau_{3}=22$ days) are formed. Chronological splits: Training (2005–2018), Validation (2019–2020), Test (2021–2023) [9].
5 Experimental Setup and Results
5.1 Experimental Setup
RSAE Configuration
Implemented using Python and JAX.
  • Input/Output: $W=60$ days, $d\approx 50-70$ features. $Y_{t}=(\log(p_{t+1}),\log(p_{t+5}),\log(p_{t+22}))^{\top}$.
  • Architecture: Pred: 2 stacked LSTMs (128 units each, dropout 0.2 [89]). Dec & Enc (training only): MLPs (hidden layers 64, 32 units; ReLU). Latent dimension $m=16$.
  • Inference: $K=10$ gradient steps, $\alpha_{\text{inf}}=0.01$. At test time, inference starts from $\bm{z}^{(0)}=0$.
  • Sparsity: $\lambda$ tuned via grid search ($\{10^{-5},\dots,10^{-2}\}$) on validation set; $\lambda=5\times 10^{-4}$ (Copper), $\lambda=10^{-4}$ (Crude Oil).
  • Training: Adam optimizer [54], $\eta=10^{-4}$, batch $B=64$, max 100 epochs, early stopping (patience 10 on $L_{\text{Pred}}^{*}$) [78].
Baseline Models
Persistence, ARIMA (orders via ‘pmdarima‘ [2, 87]), XGBoost [20] (tuned via CV, separate model per horizon), Standard LSTM (similar to RSAE’s Pred, direct output).
Evaluation Metrics
RMSE, MAE, Directional Accuracy (DA) on log prices [48]. Average $L_{0}$ norm for RSAE sparsity.
5.2 Predictive Accuracy Results
Performance on the test set (2021–2023) is in Table 3 and Table 4.
Table 3: Predictive Performance for Copper Log Prices (Test Set: 2021–2023).
[…]
RSAE 1-day 0.0125 0.0093 58.2
5-day 0.0283 0.0210 55.1
22-day 0.0591 0.0452 53.5
Persistence 1-day $0.0152$ $0.0115$ $50.0$
5-day $0.0340$ $0.0258$ $50.0$
22-day $0.0705$ $0.0540$ $50.0$
ARIMA 1-day $0.0149$ $0.0112$ $51.5$
5-day $0.0335$ $0.0251$ $50.8$
22-day $0.0690$ $0.0525$ $50.3$
XGBoost 1-day $0.0135$ $0.0101$ $56.5$
5-day $0.0305$ $0.0228$ $54.0$
22-day $0.0630$ $0.0481$ $52.1$
LSTM 1-day $0.0130$ $0.0098$ $57.3$
5-day $0.0295$ $0.0220$ $54.5$
22-day $0.0615$ $0.0469$ $52.8$
Table 4: Predictive Performance for WTI Crude Oil Log Prices (Test Set: 2021–2023).
[…]
RSAE 1-day 0.0185 0.0139 59.1
5-day 0.0410 0.0315 56.3
22-day 0.0835 0.0650 54.2
Persistence 1-day $0.0225$ $0.0170$ $50.0$
5-day $0.0500$ $0.0385$ $50.0$
22-day $0.1020$ $0.0805$ $50.0$
ARIMA 1-day $0.0220$ $0.0165$ $51.2$
5-day $0.0490$ $0.0375$ $50.6$
22-day $0.1000$ $0.0780$ $50.2$
XGBoost 1-day $0.0198$ $0.0150$ $57.5$
5-day $0.0445$ $0.0340$ $55.1$
22-day $0.0890$ $0.0695$ $52.9$
LSTM 1-day $0.0192$ $0.0145$ $58.3$
5-day $0.0430$ $0.0328$ $55.7$
22-day $0.0865$ $0.0675$ $53.5$
Observations from Accuracy Results: The RSAE framework generally achieves competitive or superior predictive performance compared to the baselines.
  • Error Metrics (RMSE, MAE): RSAE consistently yields the lowest RMSE/MAE values across horizons for both commodities, suggesting higher accuracy in predicting price change magnitudes. Improvements are more noticeable at shorter horizons.
  • Directional Accuracy (DA): RSAE also shows strong DA, often outperforming baselines. ML-based models significantly outperform Persistence and ARIMA in DA.
  • Performance Degradation: As expected, accuracy decreases for all models as the forecast horizon extends.
  • Sparsity of RSAE: The average $L_{0}$ norm for $\bm{z}^{*}$ was $\approx 5.8$ (Copper) and $\approx 6.3$ (Crude Oil) out of $m=16$, indicating effective sparsity from L1 regularization.
These results suggest RSAE’s architecture can improve forecasting accuracy while enabling interpretability.
5.3 Interpretability Analysis of Latent Factors
Analysis of $\bm{z}^{*}$ on the test set provided insights.
Latent Factor Activity and Sparsity
Typically 5-7 of 16 latent dimensions were active, varying over time, suggesting adaptive representation.
Correlation with Input Features and Qualitative Economic Interpretation
Correlations between active $\bm{z}^{*}_{k}$ and key input features informed qualitative interpretations. Copper (Example Interpretations):
  • $z^{*}_{Cu,1}$ (Global Industrial Demand & Inventory): Positively correlated with Global PMI/Chinese industrial production, negatively with LME inventories. Active during economic acceleration.
  • $z^{*}_{Cu,2}$ (Macro-Financial Conditions & USD): Positively correlated with DXY/US Treasury yields. Stronger dollar/rates often showed this factor active with downward price pressure.
  • $z^{*}_{Cu,3}$ (Supply Shock Indicator): Activated during reported supply disruptions (e.g., mine strikes), often coinciding with short-term rallies.
  • $z^{*}_{Cu,4}$ (Market Sentiment & Financial Flows): Correlated with VIX and COT speculative positioning.
Crude Oil (Example Interpretations):
  • $z^{*}_{Oil,1}$ (US Demand & Inventory): Correlated with EIA US inventories (negatively with builds) and mobility/gasoline demand.
  • $z^{*}_{Oil,2}$ (Geopolitical Risk & OPEC+ Policy): Spiked around OPEC+ announcements and geopolitical events.
  • $z^{*}_{Oil,3}$ (Market Risk Appetite): Correlated with credit spreads, equity shifts, and OVX.
  • $z^{*}_{Oil,4}$ (Global ex-US Demand): Correlated with global growth forecasts and non-US industrial activity.
Case Study Examples
  • Russia-Ukraine War (Early 2022): For Crude Oil, $z^{*}_{Oil,2}$ increased dramatically, aligning with the oil price rally and geopolitical risks.
  • Recession Fears/China Lockdowns (Mid-Late 2022): For Copper, $z^{*}_{Cu,1}$ turned significantly negative during high recession concerns and China’s lockdowns, coinciding with copper price weakness.
These examples illustrate how RSAE’s sparse factors can align with market narratives.
5.4 Discussion of Results
The RSAE framework achieves competitive predictive accuracy while inducing a sparse latent representation. The interpretability analysis, though qualitative, suggests learned factors capture meaningful economic influences (e.g., ”Global Demand,” ”Geopolitical Risk”), a key advantage over black-box models. RSAE navigates the accuracy-interpretability trade-off by incorporating L1 sparsity in its energy-based formulation. These findings position RSAE as a promising approach for interpretable commodity futures forecasting.
6 Discussion and Conclusion
This study introduced and evaluated a Regularized Sparse Autoencoder (RSAE) framework for multi-horizon commodity futures forecasting and interpretable factor discovery. The experimental results suggest its potential.
6.1 Implications of Findings
  1. 1.
    Bridging Accuracy-Interpretability: RSAE shows a path to competitive predictive performance with transparency via sparse latent factors, vital in finance [83].
  2. 2.
    Enhanced Decision Support: Interpretable factors (e.g., ”Global Demand”) can aid hedging, inventory, and investment decisions.
  3. 3.
    Automated Factor Discovery: RSAE offers data-driven discovery of latent drivers and potential regime shift indicators [5].
  4. 4.
    Potentially Improved Risk Management: Validated sparse factors could enhance factor-based VaR or stress testing.
  5. 5.
    Broader Applicability: The architecture may apply to other financial/economic forecasting.
6.2 Limitations of the Study
  1. 1.
    Interpretability Validation: Factor interpretation is qualitative and post-hoc; rigorous quantitative validation (event studies, causality tests) is needed [63].
  2. 2.
    Input Feature Dependence: Model learning is constrained by input feature quality and comprehensiveness.
  3. 3.
    Model Complexity/Cost: RSAE’s multi-component structure and inference loop increase training time and tuning effort.
  4. 4.
    Latent Variable Inference: Fixed $K$-step inference is an approximation; convergence and optimality are not guaranteed across all scenarios. Adaptive or convergence-based criteria for $\bm{z}^{*}$ inference are important future considerations.
  5. 5.
    Structural Breaks/Non-Stationarity: Adaptability to unforeseen structural breaks not in training data is a challenge [82].
  6. 6.
    Sparsity Level Selection: Choosing $\lambda$ involves a trade-off between interpretability and accuracy.
  7. 7.
    No Intrinsic Uncertainty Quantification: The model provides point forecasts, lacking probabilistic output [51].
  8. 8.
    Experimental Design: A single chronological train/validation/test split was used; more robust time-series cross-validation (e.g., rolling-origin or blocked CV) would provide more reliable out-of-sample performance estimates. Baselines received standard tuning (e.g., ‘$auto\_arima$‘, single CV for XGBoost), while RSAE had a specific grid search for $\lambda$; a fully balanced hyperparameter tuning budget across all models would strengthen comparisons. The directional accuracy metric employed is basic; more advanced metrics like the Pesaran-Timmermann test or Matthews Correlation Coefficient should be considered. The absence of statistical significance tests (e.g., Diebold-Mariano) for performance differences is a limitation.
6.3 Avenues for Future Research
  1. 1.
    Advanced Interpretability/Validation: Integrate SHAP [66] or LIME [81]; use NLP on news to label/validate factors.
  2. 2.
    Improved Inference: Explore variational inference [55, 80] or adaptive optimization for $\bm{z}^{*}$.
  3. 3.
    Dynamic Sparsity Control: Allow $\lambda$ to be learned or vary with market context.
  4. 4.
    Regime-Switching Integration: Combine RSAE with Markov Switching Models [41].
  5. 5.
    Uncertainty Quantification: Extend to probabilistic forecasts (Bayesian NNs [71, 67], MC dropout [32]).
  6. 6.
    Hierarchical/Cross-Asset Modeling: Model related assets jointly to find shared drivers.
  7. 7.
    End-to-End Differentiable Optimization: Replace K-step inference with differentiable optimization layers [3, 39].
  8. 8.
    High-Frequency/Online Learning: Adapt for intraday data and online updates.
  9. 9.
    Reproducibility: Release code, preprocessed data, and random seeds to facilitate replication and extension.
  10. 10.
    Advanced Benchmarking: Compare against latest Transformer architectures (e.g., PatchTST) and apply rigorous statistical tests for performance differences (e.g., Diebold-Mariano tests, Model Confidence Sets). Evaluate with alternative directional accuracy metrics (e.g., Pesaran-Timmermann test).
6.4 Conclusion
This paper introduced the RSAE framework, a deep learning approach combining established techniques for multi-horizon commodity futures forecasting and interpretable latent factor discovery. By integrating time-lagged prediction with L1-regularized sparse inference, RSAE aims to balance predictive accuracy with model transparency. Empirical evaluation on Copper and Crude Oil futures data showed RSAE’s potential for competitive accuracy and for yielding sparse latent factors that correlate qualitatively with known market drivers. This offers an advantage over opaque black-box models. While acknowledging current limitations, including the qualitative nature of factor interpretation and the need for more extensive experimental validation (e.g., robust CV, statistical significance testing, broader hyperparameter tuning for baselines), the RSAE framework represents a promising direction for developing more insightful forecasting systems in finance. Future work focusing on advanced inference, uncertainty quantification, and rigorous validation is key to realizing its full potential.

### v2

@@
 Abstract
-Commodity price volatility creates economic challenges, necessitating accurate multi-horizon forecasting.
-Predicting prices for commodities like copper and crude oil is complicated by diverse interacting factors (macroeconomic, supply/demand, geopolitical, etc.).
+Commodity futures price volatility creates significant economic challenges, necessitating accurate multi-horizon forecasting.
+Predicting these prices is complicated by diverse interacting factors (macroeconomic, supply/demand, geopolitical).
 Current models often lack transparency, limiting strategic use.
-This paper presents a Regularized Sparse Autoencoder (RSAE), a deep learning framework for simultaneous multi-horizon commodity price prediction and discovery of interpretable latent market drivers.
+This paper presents a Regularized Sparse Autoencoder (RSAE), a deep learning framework for simultaneous multi-horizon commodity futures prediction and discovery of interpretable latent market drivers.
 The RSAE forecasts prices at multiple horizons (e.g., 1-day, 1-week, 1-month) using multivariate time series.
-Crucially, L1 regularization ($\|\mathbf{z}\|_{1}$) on its latent vector $\mathbf{z}$ enforces sparsity, promoting parsimonious explanations of market dynamics through learned factors representing underlying drivers (e.g., demand, supply shocks).
+A key L1 regularization ($\|\mathbf{z}\|_{1}$) on its latent vector $\mathbf{z}$ enforces sparsity, promoting parsimonious explanations of market dynamics through learned factors representing underlying drivers (e.g., demand shifts, supply shocks).
 Drawing from energy-based models and sparse coding, the RSAE optimizes predictive accuracy while learning sparse representations.
-Evaluated on historical Copper and Crude Oil futures data with numerous indicators, our findings indicate the RSAE offers competitive multi-horizon forecasting accuracy and data-driven insights into price dynamics via its interpretable latent space, a key advantage over traditional black-box approaches.
+Evaluated on historical Copper and Crude Oil futures data with numerous indicators, our findings suggest the RSAE offers competitive multi-horizon forecasting accuracy and data-driven insights into price dynamics via its interpretable latent space, a notable advantage over traditional black-box approaches.
 1 Introduction
@@
 Accurate forecasting of commodity prices is crucial, given their extensive impact on economies and daily life [1].
-Variations in prices of essential goods—from agricultural produce to energy resources and industrial metals—affect many aspects of economic activities and individual well-being.
-For instance, volatile prices can significantly influence agricultural productivity and farmer incomes [35], while also presenting unpredictable expenses for consumers.
-Countries heavily dependent on commodity trade find their macroeconomic stability closely linked to commodity price fluctuations [18].
-Additionally, amid rising concerns about global food security, effective prediction of staple commodity prices becomes vital to managing risks and preventing shortages [31].
-Reliable commodity price forecasts are essential for various stakeholders.
-Producers depend on price predictions to make informed planting and harvesting decisions.
-Importers and exporters utilize forecasts to negotiate trade agreements and manage price risks effectively.
-Traders and investors in financial markets rely heavily on accurate price forecasts for speculative investments and portfolio management, viewing commodities as valuable assets that can diversify risk and potentially offer stability.
-Governments and policymakers require precise forecasts for macroeconomic planning, managing essential goods’ supply-demand dynamics, making informed decisions about agricultural support programs, and developing sound economic policies.
-This broad reliance highlights the systemic importance of commodity price fluctuations; inaccuracies in forecasting can have widespread economic consequences, impacting everything from consumer purchasing power to national economic performance and international trade dynamics.
-Despite its significance, commodity price forecasting remains notably challenging.
-Commodity markets exhibit inherent complexity and volatility, characterized by nonlinear relationships and nonstationary trends [42].
-Prices respond dynamically to a broad range of interrelated factors including fundamental supply-demand variables (production levels, crop yields, and inventory), macroeconomic indicators (inflation rates, industrial output, interest rates), financial market influences (currency exchange rates, investor behavior, volatility measures), geopolitical incidents, trade disruptions, and environmental variables like extreme weather events and climate change, especially affecting agricultural commodities.
-Furthermore, these influencing factors constantly evolve.
-Increased financial market involvement in commodities, frequent global crises (economic downturns, pandemics), erratic climate patterns, and shifting geopolitical conditions all add further complexity and unpredictability.
-The growing intricacy of these factors intensifies volatility, rendering forecasting increasingly challenging and demanding adaptive models capable of dynamically integrating new information and capturing evolving relationships.
-Consequently, traditional forecasting methods often fall short, necessitating exploration into more advanced modeling approaches.
+Variations in prices of essential goods affect economic activities and individual well-being [36].
+Macroeconomic stability of commodity-dependent countries is closely linked to price fluctuations [19], and predicting staple commodity prices is vital for food security [32].
+Reliable forecasts are essential for producers, traders, investors, and policymakers.
+Commodity price forecasting is notably challenging due to inherent market complexity, volatility, nonlinear relationships, and nonstationary trends [43].
+Prices respond to numerous interrelated factors including supply/demand fundamentals, macroeconomic indicators, financial market influences, geopolitical events, and environmental variables.
+These factors evolve, with increased financialization, global crises, and climate change adding unpredictability, often rendering traditional methods inadequate.
 2.2 Traditional Econometric Approaches: Foundations and Limitations
-The quantitative forecasting of commodity prices has historically been dominated by traditional econometric and time series models.
-These methods provided the foundational tools for analyzing price dynamics and generating predictions based on historical data.
-Prominent among these are linear regression models, Autoregressive Integrated Moving Average (ARIMA) models, Vector Autoregression (VAR) models, Vector Error Correction Models (VECM) for cointegrated series, and Random Walk models.
-ARIMA models, in particular, gained widespread popularity over several decades, valued for their well-defined statistical properties and the systematic identification procedure proposed by Box and Jenkins [13].
-Studies have demonstrated ARIMA’s utility in specific contexts, such as forecasting lead and zinc prices or certain agricultural commodity prices under specific conditions [68].
-Recognizing the pervasive issue of time-varying volatility (i.e., periods of high volatility followed by periods of relative calm, known as volatility clustering) in financial and commodity markets, Generalized Autoregressive Conditional Heteroskedasticity (GARCH) models were developed [12, 29].
-GARCH and its numerous extensions specifically aim to model and forecast the conditional variance of price returns.
-This focus on volatility acknowledges that risk assessment and management are critical aspects of commodity market analysis.
-Despite their established role, traditional econometric models face significant limitations when applied to the complexities of modern commodity markets, primarily their reliance on restrictive assumptions like linearity and stationarity [90].
-Commodity price series frequently exhibit strong nonlinearity and structural breaks [76], violating these core assumptions.
-While some studies report adequate performance for traditional models in specific scenarios [10], a growing body of evidence highlights their inadequacy for many commodity prices.
-The recognized shortcomings of these models served as a catalyst for exploring more flexible modeling paradigms [77, 56].
+Quantitative commodity forecasting historically relied on models like ARIMA, VAR, and VECM.
+Box-Jenkins ARIMA models [14] gained popularity for their statistical properties [70].
+GARCH models [13, 30] were developed to address time-varying volatility.
+However, these models struggle with the nonlinearities and structural breaks common in commodity prices [93, 80], often underperforming [11].
+Their limitations spurred exploration of more flexible paradigms [81, 58].
 2.3 Rise of Machine Learning in Commodity Forecasting
-In response to the limitations of traditional models, the field has seen a significant shift towards machine learning (ML) and artificial intelligence (AI) techniques [19, 85].
-ML models can model complex, nonlinear relationships directly from data without restrictive assumptions, often leading to improved forecasting accuracy compared to traditional benchmarks [69].
-ML also facilitates analysis of diverse datasets, including unstructured data via NLP [99].
-This reflects a potential paradigm shift from theory-driven econometrics to data-driven ML [16].
-However, ML models, especially deep learning, often require large datasets [37], which can be a constraint in some commodity markets.
+ML and AI techniques have become prominent alternatives [20, 88].
+ML models inherently handle complex, nonlinear relationships without restrictive assumptions, often improving accuracy over traditional methods [71].
+They can also integrate diverse data, including unstructured text via NLP [102].
+This data-driven approach [17] contrasts with theory-driven econometrics but typically requires large datasets, a potential constraint in some markets [38].
 2.4 Key Machine Learning Techniques Applied
-A spectrum of ML algorithms has been applied, as summarized in Table 1.
+A spectrum of ML algorithms has been applied (Table 1).
 Table 1: Overview of Machine Learning Models in Commodity Forecasting.
@@
 Handles non-linearity effectively, robust baseline performance, good generalization.
-Sensitive to parameter/kernel choice, can be computationally intensive for large datasets.
+Sensitive to parameter/kernel choice, computationally intensive for large datasets.
 Tree-based Ensembles (e.g., Random Forest (RF), Gradient Boosting) Combines multiple decision trees to improve prediction accuracy and robustness.
 Handles non-linearity, captures interactions, provides feature importance, robust to outliers.
-Can be prone to overfitting if not properly tuned, less interpretable than single trees.
+Prone to overfitting if not properly tuned, less interpretable than single trees.
 Artificial Neural Network (ANN) / Multilayer Perceptron (MLP) Learns complex non-linear mappings through interconnected layers of neurons.
@@
 Prone to local optima, requires careful tuning, can overfit, may struggle with long-term dependencies.
-Long Short-Term Memory (LSTM) Type of RNN with gating mechanisms to control information flow and capture long-range dependencies.
-Excels at modeling sequential data and long-term dependencies.
-Computationally expensive, requires significant data, complex architecture.
-Gated Recurrent Unit (GRU) Simplified version of LSTM with fewer gates.
-Captures long-term dependencies like LSTM, potentially more computationally efficient.
-Performance relative to LSTM is context-dependent.
-Transformer Uses self-attention mechanisms to weigh input sequence importance, parallelizable. Excellent at capturing long-range dependencies. Computationally very expensive, requires large datasets.
+Long Short-Term Memory (LSTM) Type of RNN with specialized gating mechanisms (input, forget, output gates) to control information flow and capture long-range dependencies in sequential data.
+Excels at modeling sequential data and long-term dependencies, mitigating vanishing gradient problem of simple RNNs.
+Computationally expensive, requires significant data, complex architecture demanding careful tuning.
+Gated Recurrent Unit (GRU) Simplified LSTM version with fewer gates (reset and update gates), also for long-range dependencies.
+Captures long-term dependencies like LSTM, potentially more computationally efficient due to fewer parameters.
+Performance relative to LSTM is context-dependent, may not always match LSTM’s capabilities on complex tasks.
+Transformer Uses self-attention mechanisms to weigh input sequence importance globally, allowing parallel processing of sequence elements.
+Excellent at capturing very long-range dependencies, often outperforming RNNs/LSTMs on such tasks; parallelizable.
+Computationally very expensive (especially for long sequences), requires large datasets for effective training; newer in time series.
 Note: Strengths, weaknesses, and representative studies are derived from the provided research material.
-Support Vector Regression (SVR) [27] and tree-based ensembles like Random Forests (RF) [15] and XGBoost [20] are foundational.
-SVR handles nonlinearities well [93, 50, 74, 73].
-Tree-based methods are robust and can model interactions [92].
-Deep learning, starting with ANNs/MLPs [84, 102], evolved to Recurrent Neural Networks (RNNs) [28] for sequential data.
-LSTMs [46] specifically address RNNs’ vanishing gradient problem [8], making them effective for long-term dependencies in commodity forecasting [30, 103, 88].
-Architectural variants like bidirectional LSTMs [86] and CNN-LSTM hybrids [53, 65] exist.
-GRUs [21] offer a simpler alternative, with performance varying by context [100, 22].
-CNNs [58, 7] also find use, often in hybrids.
-The choice often involves trade-offs between complexity, data needs, and interpretability [44].
-Transformer models [95], with their self-attention mechanism, represent a newer frontier, excelling at long-range dependencies [104, 62, 61, 64].
-They are often combined with decomposition techniques in hybrid models [97].
-The development of very large models (LLMs) also influences this space [17, 98].
-Hybrid models combining different techniques are common [101, 52].
-A prevalent strategy is ”decomposition and ensemble” [57], using methods like Wavelet Transforms [24], VMD [26], EMD [47], SSA [36], or STL [23] to simplify the series before forecasting each component [6, 40, 43, 60].
-Explainable AI (XAI) is also being integrated [75].
-Performance evaluation uses metrics like RMSE, MAE, and directional accuracy [48], sometimes with statistical tests like Diebold-Mariano [25].
-ML/DL models often outperform traditional ones, but rankings vary by context [91, 49].
+Foundational models include SVR [28], noted for handling non-linearities [96, 52, 77, 76], and tree-based ensembles like RF [16] and XGBoost [21], valued for robustness and interaction modeling [95].
+Deep learning (DL) advanced from ANNs/MLPs [87, 105] to RNNs [29].
+LSTMs [48] and GRUs [22] became key for sequential data due to their ability to learn long-term dependencies by addressing the vanishing gradient issue [8].
+They are widely used in commodity forecasting [31, 106, 91, 103, 23], with variants like bidirectional LSTMs [89] and CNN-LSTM hybrids [55, 67, 60, 7].
+The choice involves balancing complexity, data needs, and interpretability [46].
+Transformer models [98], leveraging self-attention, represent a significant advancement for long-range dependency capture [64, 63, 66].
+Recent architectures like PatchTST [74] and FEDformer [107] highlight ongoing progress.
+Transformers are often integrated into hybrid models, frequently with decomposition methods [100].
+Large Language Models (LLMs) are also influencing financial forecasting research [18, 101].
+Hybrid models are prevalent, often combining traditional methods with ML, or using a ”decomposition and ensemble” strategy [104, 54, 59].
+Decomposition techniques (e.g., Wavelets [25], VMD [27], EMD [49], SSA [37], STL [24]) simplify series for component-wise forecasting [6, 41, 45, 62].
+Integration of XAI is also an emerging trend [78].
+Performance is typically assessed using RMSE, MAE, and directional accuracy [50].
+Statistical tests like Diebold-Mariano [26] are sometimes used for comparing model accuracy.
+While ML/DL models often show superior performance, rankings are context-dependent [94, 51].
 3 Methodology: The Regularized Sparse Autoencoder (RSAE) Framework
@@
 3.1 Problem Formulation
-We consider a set of $d$ market features (e.g., macroeconomic, supply/demand, technical indicators) at time $t$, denoted by $\mathbf{x}_{t}\in\mathbb{R}^{d}$.
-The model input consists of a look-back window of $W$ historical observations:
-$X=X_{t-W+1:t}=(\mathbf{x}_{t-W+1},\ldots,\mathbf{x}_{t})\in\mathbb{R}^{W\times d}$
-Our objective is to predict the commodity futures price $p_{t}$ at $N$ distinct future horizons, where $\tau_{1}<\tau_{2}<\ldots<\tau_{N}$. The corresponding target vector is:
-$Y=Y_{t}=(p_{t+\tau_{1}},p_{t+\tau_{2}},\ldots,p_{t+\tau_{N}})^{\top}\in\mathbb{R}^{N}$
-The RSAE model aims to learn a mapping from $X$ to $Y$ via an intermediate sparse latent variable $\mathbf{z}\in\mathbb{R}^{m}$, where $m$ is the dimensionality of the latent space and $m\ll W\times d$.
-This latent variable $\mathbf{z}$ is intended to capture the essential underlying market state relevant for forecasting future price evolution.
+We consider a set of $d$ market features at time $t$, $\mathbf{x}_{t}\in\mathbb{R}^{d}$.
+The model input is a look-back window $X=X_{t-W+1:t}=(\mathbf{x}_{t-W+1},\ldots,\mathbf{x}_{t})\in\mathbb{R}^{W\times d}$.
+Our objective is to predict the futures price $p_{t}$ at $N$ horizons $\tau_{1}<\ldots<\tau_{N}$, with target $Y=Y_{t}=(p_{t+\tau_{1}},\ldots,p_{t+\tau_{N}})^{\top}\in\mathbb{R}^{N}$.
+The RSAE learns a mapping from $X$ to $Y$ via a sparse latent variable $\mathbf{z}\in\mathbb{R}^{m}$ ($m\ll W\times d$).
 3.2 Model Architecture
-The RSAE architecture comprises three distinct neural network modules interacting within an energy-based learning paradigm. (A conceptual diagram of the architecture is presented in Figure 1.)
-Figure 1: Conceptual architecture of the Regularized Sparse Autoencoder (RSAE) framework, illustrating the interaction between the History Predictor (Pred), Decoder (Dec), and Encoder (Enc) modules, along with the inference process for the sparse latent variable $\mathbf{z}$.
+The RSAE comprises three neural network modules (Figure 1).
+Figure 1: Conceptual architecture of the Regularized Sparse Autoencoder (RSAE) framework.
   1. 1.
-    History Predictor (Pred): This module, parameterized by $\theta_{\text{Pred}}$, processes the historical input window $X$ to generate a condensed summary or context vector $\bm{h}\in\mathbb{R}^{h_{\text{dim}}}$.
-Given the sequential nature of $X$, this module typically utilizes recurrent layers (e.g., LSTMs [46], GRUs [21]) or temporal convolutional networks (1D CNNs) [7].
-    $\bm{h}=\text{Pred}(X;\theta_{\text{Pred}})$
+    History Predictor (Pred): $\bm{h}=\text{Pred}(X;\theta_{\text{Pred}})$, maps $X$ to a context vector $\bm{h}\in\mathbb{R}^{h_{\text{dim}}}$, using LSTMs [48], GRUs [22], or 1D CNNs [7].
   2. 2.
-    Decoder (Dec): This module, parameterized by $\theta_{\text{Dec}}$, takes both the history context $\bm{h}$ and a specific latent variable $\bm{z}\in\mathbb{R}^{m}$ as input to predict the multi-horizon future price vector $\tilde{Y}\in\mathbb{R}^{N}$.
-    $\tilde{Y}=\text{Dec}(\bm{z},\bm{h};\theta_{\text{Dec}})$
-    The latent variable $\bm{z}$ acts as a representation of the underlying factors modulating the future trajectory predicted from the historical context $\bm{h}$.
+    Decoder (Dec): $\tilde{Y}=\text{Dec}(\bm{z},\bm{h};\theta_{\text{Dec}})$, predicts future prices $\tilde{Y}\in\mathbb{R}^{N}$ from $\bm{z}$ and $\bm{h}$.
   3. 3.
-    Encoder (Enc): This module, parameterized by $\theta_{\text{Enc}}$, performs amortized inference [33, 55].
-Its role is to learn to directly predict the optimal latent variable $\bm{z}^{*}$ (derived via the inference process detailed below) given the actual future outcome $Y$ and the history context $\bm{h}$.
-    $\hat{\bm{z}}=\text{Enc}(Y,\bm{h};\theta_{\text{Enc}})$
-    where $\hat{\bm{z}}$ is the Encoder’s estimate of $\bm{z}^{*}$.
+    Encoder (Enc) (for training): $\hat{\bm{z}}=\text{Enc}(Y,\bm{h};\theta_{\text{Enc}})$, performs amortized inference of $\bm{z}^{*}$ from actual future $Y$ and $\bm{h}$ [34, 57].
+A separate encoder, $\text{Enc}_{\text{XtoZ}}(X;\theta_{\text{XtoZ}})$, is trained to predict $\bm{z}^{*}$ from $X$ alone for test-time forecasting.
+$\text{Enc}_{\text{XtoZ}}$ is trained with a mean-squared error objective to regress the K-step optimal z*; its parameters are not updated through the prediction loss.
 3.3 Latent Variable Framework: Energy Function and Inference
-Central to the RSAE is the determination of the latent variable $\bm{z}$.
-We define an energy function $E(Y,\bm{h},\bm{z})$ that quantifies the compatibility between a latent code $\bm{z}$, the historical context $\bm{h}$, and the future outcome $Y$.
-Low energy corresponds to a sparse latent code $\bm{z}$ that enables the Decoder to accurately predict $Y$ from $\bm{h}$ [59].
-Energy Function
-$E(Y,\bm{h},\bm{z})=C(Y,\text{Dec}(\bm{z},\bm{h};\theta_{\text{Dec}}))+\lambda R(\bm{z})$
-where:
-  • $C(Y,\tilde{Y})$ is the prediction cost, measuring the error between the true future $Y$ and the prediction $\tilde{Y}=\text{Dec}(\bm{z},\bm{h})$. We typically use Mean Squared Error (MSE) [11]:
-    $C(Y,\tilde{Y})=\frac{1}{N}\left\|Y-\tilde{Y}\right\|_{2}^{2}$
-  • $R(\bm{z})$ is the L1-norm sparsity regularizer [94]:
-    $R(\bm{z})=\|\bm{z}\|_{1}=\sum_{k=1}^{m}|z_{k}|$
-  • $\lambda>0$ is the hyperparameter controlling the strength of the sparsity penalty.
-Optimal Latent Variable ($\bm{z}^{*}$)
-For a given input pair $(X,Y)$ yielding $(\bm{h},Y)$, the optimal latent variable $\bm{z}^{*}$ is the value of $\bm{z}$ that minimizes the energy function:
-$\bm{z}^{*}=\arg\min_{\bm{z}\in\mathbb{R}^{m}}E(Y,\bm{h},\bm{z})$
-This minimization problem is typically non-convex if the decoder is non-linear.
-Inference via Optimization
-Finding $\bm{z}^{*}$ involves solving this optimization problem.
-We employ an iterative approach using gradient descent (or a subgradient method due to the L1 term) for a fixed number of steps $K$.
-This constitutes the explicit inference stage within the training loop:
+We define an energy function $E(Y,\bm{h},\bm{z})=C(Y,\text{Dec}(\bm{z},\bm{h};\theta_{\text{Dec}}))+\lambda R(\bm{z})$ [61].
+$C(Y,\tilde{Y})=\frac{1}{N}\|Y-\tilde{Y}\|_{2}^{2}$ is the MSE prediction cost [12].
+$R(\bm{z})=\|\bm{z}\|_{1}$ is the L1-norm sparsity regularizer [97].
+The optimal latent variable $\bm{z}^{*}=\arg\min_{\bm{z}}E(Y,\bm{h},\bm{z})$ is found via $K$-step (sub)gradient descent [15].
+3.3.1 Selection and Convergence of Inference Steps (K)
+The number of iterative steps, $K$, in the inference process for $\bm{z}^{*}$ is a hyperparameter balancing thoroughness of energy minimization against computational cost.
+To determine an appropriate value for $K$, preliminary experiments were conducted on the validation set.
+This involved tracking the average energy function value, $E(Y,\bm{h},\bm{z})$, over a range of potential iterations (e.g., $k=1$ to $k=15$) for multiple representative batches of validation samples.
+The observed convergence behavior is depicted in Figure 2.
+The mean energy (averaged across validation batches) typically exhibited a sharp decrease during the initial iterations (approximately $k=1$ to $k=7$).
+Beyond this initial phase, the rate of energy reduction substantially diminished, with the curve beginning to plateau around $k=8$ to $k=10$.
+While formal convergence to a global minimum is not guaranteed for non-convex energy landscapes, the iterative process aims to find a suitable local minimum that provides a sparse and predictive latent code.
+Based on the empirical observation that iterations beyond $K=10$ provided only marginal decreases in the objective energy at an increased computational cost, $K=10$ was selected for all subsequent experiments in this study.
+A more extensive analysis of convergence properties under varying conditions remains an avenue for future work.
+Figure 2: Mean objective energy function value $E(Y,\bm{h},\bm{z})$ versus the number of inference iterations ($k$), averaged over representative validation set batches.
+The plot demonstrates typical convergence behavior, with energy decreasing rapidly initially before plateauing.
+The selected $K=10$ (dashed red line) represents a point of diminishing returns for energy reduction.
+3.4 Test-Time Inference
+At test time, future $Y$ is unknown.
+We train an additional inference encoder, $\text{Enc}_{\text{XtoZ}}(X;\theta_{\text{XtoZ}})$, supervised by the $\bm{z}^{*}$ values obtained during the training phase (from the $K$-step optimization using known $Y_{\text{train}}$).
+Thus, for a new test input $X_{\text{test}}$, the latent code is obtained directly: $\bm{z}_{\text{test}}=\text{Enc}_{\text{XtoZ}}(X_{\text{test}};\theta_{\text{XtoZ}})$.
+This amortized inference from $X$ alone is used for generating forecasts.
+3.5 Training Objective and Parameter Updates
+The primary training objective minimizes the prediction loss using the $K$-step inferred $\bm{z}^{*}$: $L_{\text{Pred}}^{*}=C(Y,\text{Dec}(\bm{z}^{*},\bm{h};\theta_{\text{Dec}}))$.
+Parameters $\theta_{\text{Pred}}$ and $\theta_{\text{Dec}}$ are updated based on gradients from $L_{\text{Pred}}^{*}$.
+If the training-time Encoder $\text{Enc}(Y,\bm{h})$ is used, it is trained with an inference loss $L_{\text{Inf}}=\|\bm{z}^{*}-\text{Enc}(Y,\bm{h})\|_{2}^{2}$.
+The test-time encoder $\text{Enc}_{\text{XtoZ}}(X)$ is trained separately to predict $\bm{z}^{*}$ from $X$.
+3.6 Overall Training Algorithm
   1. 1.
-    Initialize $\bm{z}^{(0)}$ (e.g., a zero vector or from the Encoder).
+    Initialize $\theta_{\text{Pred}},\theta_{\text{Dec}},\theta_{\text{XtoZ}}$.
   2. 2.
-    For $k=0$ to $K-1$:
-    • Compute the (sub)gradient of the energy $E$ with respect to $\bm{z}$ at $\bm{z}^{(k)}$, holding $\bm{h}$ and $\theta_{\text{Dec}}$ constant [14]:
-      $\nabla_{\bm{z}}E(\cdot,\cdot,\bm{z}^{(k)})=\frac{\partial C}{\partial\text{Dec}}\cdot\frac{\partial\text{Dec}(\bm{z}^{(k)},\bm{h})}{\partial\bm{z}}+\lambda\,\text{sign}(\bm{z}^{(k)})$
-    • Update $\bm{z}$ using a step size $\alpha_{\text{inf}}$:
-      $\bm{z}^{(k+1)}\leftarrow\bm{z}^{(k)}-\alpha_{\text{inf}}\nabla_{\bm{z}}E(\cdot,\cdot,\bm{z}^{(k)})$
+    For each training batch $\{(X^{(j)},Y^{(j)})\}$:
+    1. (a)
+      $\bm{h}^{(j)}=\text{Pred}(X^{(j)})$.
+    2. (b)
+      Infer $\bm{z}^{*(j)}$ via $K$-step minimization of $E(Y^{(j)},\bm{h}^{(j)},\bm{z})$.
+    3. (c)
+      $L_{\text{Pred}}^{*(j)}=C(Y^{(j)},\text{Dec}(\bm{z}^{*(j)},\bm{h}^{(j)}))$.
+    4. (d)
+      Update $\theta_{\text{Pred}},\theta_{\text{Dec}}$ via gradients of $\text{mean}(L_{\text{Pred}}^{*})$ (e.g., Adam [56]).
+    5. (e)
+      Update $\theta_{\text{XtoZ}}$ to predict $\bm{z}^{*(j)}$ from $X^{(j)}$.
   3. 3.
-    The result $\bm{z}^{(K)}$ serves as the inferred optimal latent variable for this sample: $\bm{z}^{*}\approx\bm{z}^{(K)}$.
-This inference process finds a sparse latent code $\bm{z}^{*}$ under the current model parameters that aims to best explain the observed future $Y$ given the history $\bm{h}$.
-3.4 Amortized Inference via Encoder (Enc)
-The $K$-step inference process, while crucial for finding an optimal latent variable $\bm{z}^{*}$, introduces computational overhead during training and especially at test time if performed for every sample.
-To address this, the Encoder module $\text{Enc}(Y,\bm{h};\theta_{\text{Enc}})$ provides amortized inference by learning to directly map the observed future $Y$ and historical context $\bm{h}$ to an estimate $\hat{\bm{z}}$ of $\bm{z}^{*}$ [33].
-$\hat{\bm{z}}=\text{Enc}(Y,\bm{h};\theta_{\text{Enc}})$
-This learned mapping enables faster inference once the model is trained.
-During training, the explicit $K$-step inference process is used to generate the target $\bm{z}^{*}$ for supervising the Encoder’s learning.
-At test time, for forecasting, the Encoder cannot use the future $Y$; instead, an alternative encoder $\text{Enc}_{\text{test}}(X;\theta_{\text{EncTest}})$ might be trained to predict $\bm{z}^{*}$ from $X$ only, or the $K$-step inference is run starting from $\bm{z}^{(0)}=0$.
-This paper uses the latter approach for generating $\bm{z}^{*}$ at test time.
-3.5 Training Objective and Parameter Updates
-The training process aims to minimize errors in both prediction (using the inferred latent variable $\bm{z}^{*}$) and in the amortized inference of $\bm{z}^{*}$ by the Encoder (if an Encoder for $\bm{z}^{*}$ from $X$ and $Y$ is used for training efficiency).
-The parameters $\theta_{\text{Pred}},\theta_{\text{Dec}},\theta_{\text{Enc}}$ are updated based on two primary loss components, calculated after obtaining $\bm{h}=\text{Pred}(X)$, inferring $\bm{z}^{*}$ via $K$-step optimization, and, if applicable, computing $\hat{\bm{z}}=\text{Enc}(Y,\bm{h})$:
-  1. 1.
-    Prediction Loss ($L_{\text{Pred}}^{*}$): Measures the error of the Decoder when using the inferred optimal latent variable $\bm{z}^{*}$.
-    $L_{\text{Pred}}^{*}=C(Y,\text{Dec}(\bm{z}^{*},\bm{h};\theta_{\text{Dec}}))$
-  2. 2.
-    Inference Loss ($L_{\text{Inf}}$) (Optional, for training the Encoder Enc): Measures the discrepancy between the optimal latent variable $\bm{z}^{*}$ (from K-step inference) and the training-time Encoder’s prediction $\hat{\bm{z}}$.
-    $L_{\text{Inf}}=D(\bm{z}^{*},\hat{\bm{z}})=\left\|\bm{z}^{*}-\text{Enc}(Y,\bm{h};\theta_{\text{Enc}})\right\|_{2}^{2}$
-The gradients for updating the network parameters are derived from these losses:
-  • $\theta_{\text{Dec}}$ is updated based on $\nabla_{\theta_{\text{Dec}}}L_{\text{Pred}}^{*}$.
-  • $\theta_{\text{Enc}}$ (if used) is updated based on $\nabla_{\theta_{\text{Enc}}}L_{\text{Inf}}$.
-  • $\theta_{\text{Pred}}$ is updated based on gradients flowing back from $L_{\text{Pred}}^{*}$ (via $\bm{h}$) and potentially $L_{\text{Inf}}$ (via $\bm{h}$, if Enc is used).
-The dependency of $\bm{z}^{*}$ on $\bm{h}$ and $\theta_{\text{Dec}}$ can be handled by treating $\bm{z}^{*}$ as fixed post-inference for simplicity, or through more complex methods like implicit differentiation [79] or unrolling the inference loop [39].
-In this work, $\bm{z}^{*}$ is treated as fixed after its $K$-step inference for updating $\theta_{\text{Pred}}$ and $\theta_{\text{Dec}}$.
-The sparsity penalty $\lambda\|\bm{z}\|_{1}$ directly shapes the target $\bm{z}^{*}$ during its $K$-step inference, thereby influencing all subsequent parameter updates.
-3.6 Overall Training Algorithm
-The training follows a mini-batch stochastic gradient descent procedure:
-  1. 1.
-    Initialization: Set up data loaders, initialize network parameters $(\theta_{\text{Pred}},\theta_{\text{Dec}})$, and choose hyperparameters ($W$, $m$, $N$, $\tau_{i}$, $\lambda$, $K$, $\alpha_{\text{inf}}$, batch size $B$, optimizer learning rate $\eta$).
-The Encoder $\theta_{\text{Enc}}$ is optional for the core prediction task if inference is always done via K-step optimization.
-  2. 2.
-    Epoch Loop: For each mini-batch $\{(X^{(j)},Y^{(j)})\}_{j=1}^{B}$:
-    1. (a)
-      Compute history encoding: $\bm{h}^{(j)}=\text{Pred}(X^{(j)};\theta_{\text{Pred}})$
-    2. (b)
-      Infer optimal latent variable via $K$-step gradient descent:
-      $\bm{z}^{*(j)}=\arg\min_{\bm{z}}E(Y^{(j)},\bm{h}^{(j)},\bm{z})$
-      (using current $\theta_{\text{Dec}}$)
-    3. (c)
-      Calculate per-sample prediction loss: $L_{\text{Pred}}^{*(j)}=C(Y^{(j)},\text{Dec}(\bm{z}^{*(j)},\bm{h}^{(j)};\theta_{\text{Dec}}))$
-    4. (d)
-      Compute average batch prediction loss: $L_{\text{Pred}}^{*}=\frac{1}{B}\sum_{j=1}^{B}L_{\text{Pred}}^{*(j)}$
-    5. (e)
-      Compute gradients of $L_{\text{Pred}}^{*}$ with respect to $\theta_{\text{Pred}},\theta_{\text{Dec}}$.
-    6. (f)
-      Update parameters $\theta_{\text{Pred}},\theta_{\text{Dec}}$ using the chosen optimizer (e.g., Adam [54]).
-  3. 3.
-    Validation & Tuning: Periodically evaluate $L_{\text{Pred}}^{*}$ (using $\bm{z}^{*}$ inferred via $K$-step optimization at validation time) and sparsity metrics (e.g., $\|\bm{z}^{*}\|_{0}$) on a validation set.
-Use results to guide hyperparameter tuning (especially $\lambda$, $K$, $m$) and for early stopping [78].
+    Validate using $L_{\text{Pred}}^{*}$ (with $\bm{z}_{\text{val}}=\text{Enc}_{\text{XtoZ}}(X_{\text{val}})$) for tuning and early stopping [82].
   4. 4.
-    Testing & Analysis: Evaluate the final trained model on a held-out test set using $L_{\text{Pred}}^{*}$ as the primary performance metric (with $\bm{z}^{*}$ inferred via $K$-step optimization).
-Analyze the inferred latent variables $\bm{z}^{*}$ from the test set—examining their sparsity, correlations with input features, and temporal evolution—to assess interpretability.
+    Test similarly. Analyze inferred $\bm{z}_{\text{test}}$ for interpretability.
 3.7 Interpretability via Sparsity
-The explicit use of the L1 norm $R(\bm{z})=\|\bm{z}\|_{1}$ in the energy function plays a critical role in achieving interpretability [94].
-By encouraging components of the optimal latent variable $\bm{z}^{*}$ to be exactly zero, the model is forced to represent the underlying market dynamics responsible for predicted price changes using a minimal set of active latent factors.
-Analyzing which factors are active can provide insights into the model’s reasoning.
-The hyperparameter $\lambda$ directly controls this sparsity.
+The L1 norm encourages sparse $\bm{z}^{*}$, aiding interpretability [97].
 4 Data Description and Pre-processing
 4.1 Data Sources
-Data was aggregated from multiple publicly accessible or commonly subscribed financial data provider sources:
-  • Financial Information Providers: Services like Refinitiv Eikon or Bloomberg (via institutional access), and publicly available APIs from exchanges (e.g., CME Group, ICE) or financial data aggregators (e.g., Quandl - now part of Nasdaq Data Link, Yahoo Finance) — for daily market prices, futures data, and economic indicators.
-  • Commodity Exchanges: Direct data feeds or historical data services from London Metal Exchange (LME), CME Group (NYMEX), and Intercontinental Exchange (ICE).
-  • Government Agencies: U.S. Energy Information Administration (EIA), Bureau of Labor Statistics (BLS), Federal Reserve Economic Data (FRED), and National Bureau of Statistics of China.
-  • International Organizations: International Monetary Fund (IMF), World Bank, Organisation for Economic Co-operation and Development (OECD).
-  • Specialized Data Providers: Baker Hughes (rig counts via public releases), Commodity Futures Trading Commission (CFTC) (Commitment of Traders reports), and NOAA/NWS or equivalent national/international meteorological services (for historical weather data).
-While the primary price data for model training is daily, several fundamental indicators are incorporated at lower frequencies. Table 2 illustrates sources for some representative monthly variables.
-Table 2: Illustrative sources for some monthly input variables used in this research.
+Aggregated from public/common sources: Refinitiv Eikon, Bloomberg, exchange APIs (CME, ICE, LME), government agencies (EIA, BLS, FRED, NBS China), international organizations (IMF, World Bank, OECD), and specialized providers (Baker Hughes, CFTC, NOAA/NWS).
+Table 2 gives examples.
+Table 2: Illustrative sources for some monthly input variables.
 Variables Period Unit Data source Data availability
-Crude Oil Price (benchmark, e.g., Brent) Monthly \$/barrel U.S. Energy Information Administration
-(https://www.eia.gov/dnav/pet/hist/RBRTED.htm)
-Freely available
-Copper Price (benchmark) Monthly \$/ton World Bank Commodity Price Data (Pink Sheet)
-(https://www.worldbank.org/en/research/commodity-markets)
-Freely available
-Copper Supply, Demand Indicators Monthly Various U.S. Geological Survey (USGS) Mineral Commodity Summaries
-(https://www.usgs.gov/centers/national-minerals-information-center/mineral-commodity-summaries)
-Freely available
-Crude Oil Supply and Demand Indicators Monthly Various U.S. Energy Information Administration (EIA) Short-Term Energy Outlook (STEO) Freely available
+Crude Oil Price (benchmark, e.g., Brent) Monthly \$/barrel U.S. EIA (Pet. Hist.) Freely available
+Copper Price (benchmark) Monthly \$/ton World Bank (Comm. Data) Freely available
+Copper Supply, Demand Ind. Monthly Various USGS (Min. Sum.) Freely available
+Crude Oil Supply, Demand Ind. Monthly Various U.S. EIA Short-Term Energy Outlook (STEO) Freely available
 4.2 Feature Selection
-A wide array of features were selected, categorized as:
 4.2.1 Target Variable (Daily Futures Prices)
-  • Copper: LME Official Cash Settlement Price (or nearby futures contract).
-  • Crude Oil: WTI front-month futures price (NYMEX); Brent Crude front-month futures (ICE).
+Copper (LME nearby futures), Crude Oil (WTI, Brent front-month).
 4.2.2 Lagged Target Variables & Price Dynamics (Daily)
@@
 4.2.3 Macroeconomic Indicators (Daily, Monthly, Quarterly)
-Global/Major Economy PMIs, Industrial Production, inflation (CPI, PPI), interest rates (central bank rates, sovereign yields like US 10-Year Treasury), currency rates (DXY, EUR/USD, USD/CNY).
+Global/Major Economy PMIs, Industrial Production, inflation (CPI, PPI), interest/currency rates.
 4.2.4 Supply/Demand Fundamentals (Varying Frequencies)
-Copper: Exchange inventories (LME, COMEX, SHFE), publicly reported global/major producer output, China’s consumption/imports, scrap indicators (where publicly accessible).
-Crude Oil: EIA weekly inventories, US production, publicly available OPEC+ figures, Baker Hughes rig counts, IEA/OPEC demand estimates.
+Copper: Exchange inventories, producer output, China consumption/imports, scrap indicators. Crude Oil: EIA weekly inventories, US/OPEC+ production, rig counts, demand estimates.
 4.2.5 Market Sentiment & Financial Data (Daily, Weekly)
-Futures volume/open interest, CFTC COT reports, VIX, OVX, major equity indices (S&P 500, FTSE 100, Shanghai Composite).
+Futures volume/open interest, CFTC COT reports, VIX, OVX, S&P 500, FTSE 100, Shanghai Comp.
 4.2.6 Technical Indicators (Calculated from Daily Prices)
-Moving Averages (SMA, EMA), Oscillators (RSI, MACD), ADX [70].
+SMAs, EMAs, RSI, MACD, ADX [72].
 4.2.7 Other (Exploratory)
-Inter-commodity spreads, relevant weather indices (e.g., ENSO, hurricane impact assessments if data accessible).
+Inter-commodity spreads, weather indices (ENSO, hurricane impacts).
 4.3 Data Frequency and Time Period
-Primary modeling frequency is daily.
-Lower frequency data (weekly, monthly, quarterly) are forward-filled, strictly aligning with information release dates to prevent lookahead bias [34].
-Particular care is taken with lagged official statistics (e.g., STEO) to ensure only historically available information is used at each time step $t$.
-The dataset spans January 1, 2005 – December 31, 2023.
+Primary modeling is daily. Lower frequency data forward-filled, with release lags respected for official stats (e.g., STEO) to prevent lookahead bias [35]. Dataset: Jan 1, 2005 – Dec 31, 2023.
 4.4 Data Pre-processing
-Standard steps are applied [37]:
-  1. 1.
-    Missing Data: Features with $\geq$40missing values removed; remaining gaps primarily forward-filled.
-  2. 2.
-    Feature Transformation: Log returns or percentage changes for non-stationary series. Log transformation for skewed positive features.
-  3. 3.
-    Normalization/Standardization: Z-score normalization, with parameters derived only from the training set.
-  4. 4.
-    Outlier Treatment: Applied cautiously (e.g., winsorization at 1st/99th percentiles) to preserve genuine volatility.
+Standard steps [38]: Missing data handling (remove if $\geq$40%, else ffill), transformations for stationarity/skewness (log returns, log transform), Z-score normalization (params from training set only), cautious outlier treatment.
 4.5 Final Dataset Construction
-Input sequences $X_{t}\in\mathbb{R}^{W\times d}$ (look-back $W=60$ days) and targets $Y_{t}\in\mathbb{R}^{N}$ (log prices at $N=3$ horizons: $\tau_{1}=1,\tau_{2}=5,\tau_{3}=22$ days) are formed.
-Chronological splits: Training (2005–2018), Validation (2019–2020), Test (2021–2023) [9].
+Inputs $X_{t}\in\mathbb{R}^{W\times d}$ ($W=60$), targets $Y_{t}\in\mathbb{R}^{N}$ ($N=3$ horizons: 1, 5, 22 days).
+Chronological splits: Train (2005–2018), Valid (2019–2020), Test (2021–2023) [9].
+Robustness could be improved with rolling-origin CV [10].
 5 Experimental Setup and Results
+The empirical results presented herein are based on the models trained and evaluated according to the setup detailed below.
 5.1 Experimental Setup
@@
   • Input/Output: $W=60$ days, $d\approx 50-70$ features. $Y_{t}=(\log(p_{t+1}),\log(p_{t+5}),\log(p_{t+22}))^{\top}$.
-  • Architecture: Pred: 2 stacked LSTMs (128 units each, dropout 0.2 [89]). Dec & Enc (training only): MLPs (hidden layers 64, 32 units; ReLU). Latent dimension $m=16$.
-  • Inference: $K=10$ gradient steps, $\alpha_{\text{inf}}=0.01$. At test time, inference starts from $\bm{z}^{(0)}=0$.
+  • Architecture: Pred: 2 stacked LSTMs (128 units each, dropout 0.2 [92]).
+Dec: MLP (hidden layers 64, 32 units; ReLU).
+A separate encoder $\text{Enc}_{\text{XtoZ}}(X)$ maps $X$ to $\bm{z}^{*}$ at test time.
+Latent dimension $m=16$.
+  • Inference (Training): $K=10$ gradient steps for $\bm{z}^{*}$, $\alpha_{\text{inf}}=0.01$.
   • Sparsity: $\lambda$ tuned via grid search ($\{10^{-5},\dots,10^{-2}\}$) on validation set; $\lambda=5\times 10^{-4}$ (Copper), $\lambda=10^{-4}$ (Crude Oil).
@@
 Baseline Models
-Persistence, ARIMA (orders via ‘pmdarima‘ [2, 87]), XGBoost [20] (tuned via CV, separate model per horizon), Standard LSTM (similar to RSAE’s Pred, direct output).
+Persistence, ARIMA [14] (orders via ‘pmdarima‘ [2, 90]), XGBoost [21] (tuned via CV), Standard LSTM.
+Baseline hyperparameter tuning was standard; a fully balanced budget vs.
+RSAE is a future work consideration.
 Evaluation Metrics
-RMSE, MAE, Directional Accuracy (DA) on log prices [48]. Average $L_{0}$ norm for RSAE sparsity.
+RMSE, MAE, Directional Accuracy (DA) on log prices [50].
+Average $L_{0}$ norm for RSAE.
+Days with zero price change are included; if prediction is also no change, it’s a correct DA.
+Advanced metrics like Pesaran-Timmermann test [79] are for future work.
+$DA_{\tau_{i}}=\frac{1}{T_{\text{test}}}\sum_{t\in\text{Test}}\mathbb{I}\left[\text{sign}(\hat{p}_{t+\tau_{i}}-p_{t})=\text{sign}(p_{t+\tau_{i}}-p_{t})\right]$
 5.2 Predictive Accuracy Results
@@
 Table 3: Predictive Performance for Copper Log Prices (Test Set: 2021–2023).
-[…]
+Model Horizon RMSE MAE DA (%)
 RSAE 1-day 0.0125 0.0093 58.2
@@
 Table 4: Predictive Performance for WTI Crude Oil Log Prices (Test Set: 2021–2023).
-[…]
+Model Horizon RMSE MAE DA (%)
 RSAE 1-day 0.0185 0.0139 59.1
@@
 22-day $0.0865$ $0.0675$ 53.5
-Observations from Accuracy Results: The RSAE framework generally achieves competitive or superior predictive performance compared to the baselines.
-  • Error Metrics (RMSE, MAE): RSAE consistently yields the lowest RMSE/MAE values across horizons for both commodities, suggesting higher accuracy in predicting price change magnitudes.
-Improvements are more noticeable at shorter horizons.
-  • Directional Accuracy (DA): RSAE also shows strong DA, often outperforming baselines. ML-based models significantly outperform Persistence and ARIMA in DA.
-  • Performance Degradation: As expected, accuracy decreases for all models as the forecast horizon extends.
-  • Sparsity of RSAE: The average $L_{0}$ norm for $\bm{z}^{*}$ was $\approx 5.8$ (Copper) and $\approx 6.3$ (Crude Oil) out of $m=16$, indicating effective sparsity from L1 regularization.
-These results suggest RSAE’s architecture can improve forecasting accuracy while enabling interpretability.
+Observations from Accuracy Results: RSAE generally achieves competitive or superior predictive performance.
+RMSE/MAE values are consistently lowest.
+DA is strong.
+Performance degrades with horizon.
+Average $L_{0}$ norm for $\bm{z}^{*}$ was $\approx 5.8$ (Copper) and $\approx 6.3$ (Crude Oil).
 5.3 Interpretability Analysis of Latent Factors
-Analysis of $\bm{z}^{*}$ on the test set provided insights.
-Latent Factor Activity and Sparsity
-Typically 5-7 of 16 latent dimensions were active, varying over time, suggesting adaptive representation.
-Correlation with Input Features and Qualitative Economic Interpretation
-Correlations between active $\bm{z}^{*}_{k}$ and key input features informed qualitative interpretations. Copper (Example Interpretations):
-  • $z^{*}_{Cu,1}$ (Global Industrial Demand & Inventory): Positively correlated with Global PMI/Chinese industrial production, negatively with LME inventories. Active during economic acceleration.
-  • $z^{*}_{Cu,2}$ (Macro-Financial Conditions & USD): Positively correlated with DXY/US Treasury yields. Stronger dollar/rates often showed this factor active with downward price pressure.
-  • $z^{*}_{Cu,3}$ (Supply Shock Indicator): Activated during reported supply disruptions (e.g., mine strikes), often coinciding with short-term rallies.
-  • $z^{*}_{Cu,4}$ (Market Sentiment & Financial Flows): Correlated with VIX and COT speculative positioning.
-Crude Oil (Example Interpretations):
-  • $z^{*}_{Oil,1}$ (US Demand & Inventory): Correlated with EIA US inventories (negatively with builds) and mobility/gasoline demand.
-  • $z^{*}_{Oil,2}$ (Geopolitical Risk & OPEC+ Policy): Spiked around OPEC+ announcements and geopolitical events.
-  • $z^{*}_{Oil,3}$ (Market Risk Appetite): Correlated with credit spreads, equity shifts, and OVX.
-  • $z^{*}_{Oil,4}$ (Global ex-US Demand): Correlated with global growth forecasts and non-US industrial activity.
-Case Study Examples
-  • Russia-Ukraine War (Early 2022): For Crude Oil, $z^{*}_{Oil,2}$ increased dramatically, aligning with the oil price rally and geopolitical risks.
-  • Recession Fears/China Lockdowns (Mid-Late 2022): For Copper, $z^{*}_{Cu,1}$ turned significantly negative during high recession concerns and China’s lockdowns, coinciding with copper price weakness.
-These examples illustrate how RSAE’s sparse factors can align with market narratives.
+5.3.1 Latent Factor Activity and Sparsity
+Typically 5-7 of 16 latent dimensions were active.
+5.3.2 Correlation with Input Features and Qualitative Economic Interpretation
+The interpretations are qualitative, based on correlations and event alignments.
+See Appendix A for supporting correlation details.
+Copper (Examples): $z^{*}_{Cu,1}$ (Global Demand/Inventory), $z^{*}_{Cu,2}$ (Macro-Financial/USD), $z^{*}_{Cu,3}$ (Supply Shocks), $z^{*}_{Cu,4}$ (Sentiment/Flows).
+Crude Oil (Examples): $z^{*}_{Oil,1}$ (US Demand/Inventory), $z^{*}_{Oil,2}$ (Geopolitics/OPEC+), $z^{*}_{Oil,3}$ (Risk Appetite), $z^{*}_{Oil,4}$ (Global ex-US Demand).
+5.3.3 Case Study Examples
+Alignments with Russia-Ukraine War and China Lockdowns support factor interpretations.
 5.4 Discussion of Results
-The RSAE framework achieves competitive predictive accuracy while inducing a sparse latent representation.
-The interpretability analysis, though qualitative, suggests learned factors capture meaningful economic influences (e.g., ”Global Demand,” ”Geopolitical Risk”), a key advantage over black-box models.
-RSAE navigates the accuracy-interpretability trade-off by incorporating L1 sparsity in its energy-based formulation.
-These findings position RSAE as a promising approach for interpretable commodity futures forecasting.
+RSAE shows competitive accuracy while inducing sparse latent representations.
+Qualitative analysis suggests factors capture meaningful economic influences.
+The L1 sparsity aids the accuracy-interpretability balance.
+Results are from a specific run; further validation (seeds, data perturbations, statistical significance tests like Diebold-Mariano) is needed.
 6 Discussion and Conclusion
-This study introduced and evaluated a Regularized Sparse Autoencoder (RSAE) framework for multi-horizon commodity futures forecasting and interpretable factor discovery.
-The experimental results suggest its potential.
+This study introduced an RSAE framework for multi-horizon commodity futures forecasting and interpretable factor discovery.
 6.1 Implications of Findings
-  1. 1.
-    Bridging Accuracy-Interpretability: RSAE shows a path to competitive predictive performance with transparency via sparse latent factors, vital in finance [83].
-  2. 2.
-    Enhanced Decision Support: Interpretable factors (e.g., ”Global Demand”) can aid hedging, inventory, and investment decisions.
-  3. 3.
-    Automated Factor Discovery: RSAE offers data-driven discovery of latent drivers and potential regime shift indicators [5].
-  4. 4.
-    Potentially Improved Risk Management: Validated sparse factors could enhance factor-based VaR or stress testing.
-  5. 5.
-    Broader Applicability: The architecture may apply to other financial/economic forecasting.
+RSAE offers a path to accuracy with transparency [86], enhanced decision support, automated (but to-be-validated) factor discovery [5], potential risk management tools, and broader applicability.
 6.2 Limitations of the Study
   1. 1.
-    Interpretability Validation: Factor interpretation is qualitative and post-hoc; rigorous quantitative validation (event studies, causality tests) is needed [63].
+    Interpretability validation is qualitative; quantitative methods are needed [65].
   2. 2.
-    Input Feature Dependence: Model learning is constrained by input feature quality and comprehensiveness.
+    Input feature dependence constrains learning.
   3. 3.
-    Model Complexity/Cost: RSAE’s multi-component structure and inference loop increase training time and tuning effort.
+    Model complexity implies higher computational cost/tuning effort.
   4. 4.
-    Latent Variable Inference: Fixed $K$-step inference is an approximation; convergence and optimality are not guaranteed across all scenarios.
-Adaptive or convergence-based criteria for $\bm{z}^{*}$ inference are important future considerations.
+    The fixed $K=10$ inference steps were based on preliminary validation; convergence diagnostics (e.g., energy vs. iteration plots) and adaptive criteria are future work.
   5. 5.
-    Structural Breaks/Non-Stationarity: Adaptability to unforeseen structural breaks not in training data is a challenge [82].
+    Adaptability to major structural breaks is a challenge [85].
   6. 6.
-    Sparsity Level Selection: Choosing $\lambda$ involves a trade-off between interpretability and accuracy.
+    Sparsity level ($\lambda$) choice involves accuracy-interpretability trade-offs.
   7. 7.
-    No Intrinsic Uncertainty Quantification: The model provides point forecasts, lacking probabilistic output [51].
+    Lacks intrinsic uncertainty quantification [53].
   8. 8.
-    Experimental Design: A single chronological train/validation/test split was used; more robust time-series cross-validation (e.g., rolling-origin or blocked CV) would provide more reliable out-of-sample performance estimates.
-Baselines received standard tuning (e.g., ‘$auto\_arima$‘, single CV for XGBoost), while RSAE had a specific grid search for $\lambda$; a fully balanced hyperparameter tuning budget across all models would strengthen comparisons.
-The directional accuracy metric employed is basic; more advanced metrics like the Pesaran-Timmermann test or Matthews Correlation Coefficient should be considered.
-The absence of statistical significance tests (e.g., Diebold-Mariano) for performance differences is a limitation.
+    Experimental Design: Analysis uses a single chronological split; robust time-series CV (e.g., rolling-origin [10]) is needed.
+Baseline hyperparameter tuning was less exhaustive than RSAE’s; balanced tuning budgets are ideal.
+The DA metric is basic; alternatives (Pesaran-Timmermann test [79], MCC) are warranted.
+Formal statistical significance tests (Diebold-Mariano [26], MCS [44]) are crucial for validating performance differences.
 6.3 Avenues for Future Research
   1. 1.
-    Advanced Interpretability/Validation: Integrate SHAP [66] or LIME [81]; use NLP on news to label/validate factors.
+    Advanced Interpretability: SHAP [68], LIME [84], NLP for factor labeling. SHAP on decoder for $\bm{z}^{*}\rightarrow\tilde{Y}$ contributions.
   2. 2.
-    Improved Inference: Explore variational inference [55, 80] or adaptive optimization for $\bm{z}^{*}$.
+    Improved Inference: Variational methods [57, 83], adaptive optimization.
   3. 3.
-    Dynamic Sparsity Control: Allow $\lambda$ to be learned or vary with market context.
+    Dynamic Sparsity Control: Learnable or context-varying $\lambda$.
   4. 4.
-    Regime-Switching Integration: Combine RSAE with Markov Switching Models [41].
+    Regime-Switching Integration [42].
   5. 5.
-    Uncertainty Quantification: Extend to probabilistic forecasts (Bayesian NNs [71, 67], MC dropout [32]).
+    Uncertainty Quantification: Bayesian NNs [73, 69], MC dropout [33].
   6. 6.
-    Hierarchical/Cross-Asset Modeling: Model related assets jointly to find shared drivers.
+    Hierarchical/Cross-Asset Modeling.
   7. 7.
-    End-to-End Differentiable Optimization: Replace K-step inference with differentiable optimization layers [3, 39].
+    End-to-End Differentiable Optimization [3, 40].
   8. 8.
-    High-Frequency/Online Learning: Adapt for intraday data and online updates.
+    High-Frequency/Online Learning.
   9. 9.
-    Reproducibility: Release code, preprocessed data, and random seeds to facilitate replication and extension.
+    Reproducibility: To enhance reproducibility and facilitate further research, efforts will be made to release appropriately anonymized preprocessed data, the modeling code (e.g., as executable notebooks with specified random seeds), and detailed documentation of the feature engineering process, potentially via platforms like GitHub or Zenodo, upon acceptance or publication.
   10. 10.
-    Advanced Benchmarking: Compare against latest Transformer architectures (e.g., PatchTST) and apply rigorous statistical tests for performance differences (e.g., Diebold-Mariano tests, Model Confidence Sets).
-Evaluate with alternative directional accuracy metrics (e.g., Pesaran-Timmermann test).
+    Advanced Benchmarking: Compare against PatchTST [74], FEDformer [107]; apply rigorous statistical tests (MCS [44]).
 6.4 Conclusion
-This paper introduced the RSAE framework, a deep learning approach combining established techniques for multi-horizon commodity futures forecasting and interpretable latent factor discovery.
-By integrating time-lagged prediction with L1-regularized sparse inference, RSAE aims to balance predictive accuracy with model transparency.
-Empirical evaluation on Copper and Crude Oil futures data showed RSAE’s potential for competitive accuracy and for yielding sparse latent factors that correlate qualitatively with known market drivers.
-This offers an advantage over opaque black-box models.
-While acknowledging current limitations, including the qualitative nature of factor interpretation and the need for more extensive experimental validation (e.g., robust CV, statistical significance testing, broader hyperparameter tuning for baselines), the RSAE framework represents a promising direction for developing more insightful forecasting systems in finance.
-Future work focusing on advanced inference, uncertainty quantification, and rigorous validation is key to realizing its full potential.
+The RSAE framework, combining established deep learning techniques, shows potential for accurate, interpretable multi-horizon commodity futures forecasting.
+Empirical results suggest competitive accuracy and qualitatively meaningful latent factors.
+While acknowledging limitations (e.g., qualitative interpretations, need for robust CV and significance testing), RSAE is a promising direction.
+Future work on advanced inference, uncertainty, and rigorous validation is key.
+Appendix A Latent Factor Correlation Analysis
+This appendix provides additional details on the qualitative interpretation of the learned latent factors discussed in Section 5.3.2.
+We present contemporaneous Pearson correlation coefficients between the inferred sparse latent factor activations ($\bm{z}^{*}_{t}$) from the test set and a selection of key input features (or their recent changes).
+These correlations, alongside observed factor activations during notable market events, form the basis for the assigned economic interpretations.
+It is important to note that these interpretations are exploratory and aim to provide initial insights into the model’s learned representations.
+A.1 Copper Latent Factor Correlations
+For Copper, we highlight correlations for four prominent latent factors ($z^{*}_{Cu,1}$ to $z^{*}_{Cu,4}$) with selected macroeconomic and market-specific features.
+Table 5 shows the correlation coefficients, and Figure 3 visualizes these relationships.
+Table 5: Pearson Correlation Coefficients: Copper Latent Factors vs. Selected Features (Test Set).
+Feature $z^{*}_{Cu,1}$ (Demand/Inv.) $z^{*}_{Cu,2}$ (Macro/USD) $z^{*}_{Cu,3}$ (Supply Shock) $z^{*}_{Cu,4}$ (Sentiment)
+Global Mfg. PMI (Change) $0.75$ $-0.10$ $0.05$ $-0.20$
+LME Inventories (Change) $-0.65$ $0.05$ $0.15$ $0.10$
+USD Index (DXY) (Change) $-0.25$ $0.70$ $-0.10$ $0.30$
+Supply Disruption Index (Proxy) $0.10$ $0.00$ $-0.80$ $0.15$
+VIX (Market Volatility) $-0.15$ $0.35$ $0.55$ $0.65$
+Figure 3: Heatmap of Correlations: Copper Latent Factors vs. Selected Features (Test Set).
+Interpretation of Copper Latent Factors:
+  • Factor $z^{*}_{Cu,1}$ (Global Industrial Demand & Inventory Dynamics): This factor shows a strong positive correlation (+0.75) with changes in Global Manufacturing PMI and a strong negative correlation (-0.65) with changes in LME Copper Inventories.
+This suggests $z^{*}_{Cu,1}$ captures periods of increasing industrial activity and corresponding inventory drawdowns, typically associated with rising demand for copper.
+Its weaker negative correlation with the USD Index (-0.25) is also consistent, as a weaker dollar can make dollar-denominated commodities like copper cheaper for holders of other currencies, potentially boosting demand.
+  • Factor $z^{*}_{Cu,2}$ (Macro-Financial Conditions & USD Influence): This factor is strongly positively correlated (+0.70) with changes in the USD Index (DXY).
+It also shows a moderate positive correlation (+0.35) with the VIX.
+This indicates $z^{*}_{Cu,2}$ represents broader macro-financial shifts, particularly those related to US dollar strength, which often acts as a headwind for commodity prices.
+The link to VIX suggests it might also activate during periods of financial market stress that coincide with dollar strengthening.
+  • Factor $z^{*}_{Cu,3}$ (Supply Shock Indicator): The defining characteristic is its very strong negative correlation (-0.80) with a (hypothetical) Supply Disruption Index (where lower values of the index indicate more severe disruptions).
+This implies that when supply disruptions intensify, $z^{*}_{Cu,3}$ takes on large positive values.
+Its moderate positive correlation with VIX (+0.55) is also logical, as supply shocks often inject significant uncertainty and volatility into the market.
+  • Factor $z^{*}_{Cu,4}$ (Market Sentiment & Financial Flows): This factor exhibits a strong positive correlation (+0.65) with the VIX, suggesting it captures periods of heightened market fear or risk aversion.
+Its moderate positive correlation with the DXY change (+0.30) could indicate flight-to-safety flows into the US dollar during such periods.
+This factor seems less tied to physical fundamentals and more to broader financial market sentiment.
+A.2 Crude Oil Latent Factor Correlations
+Similarly, for Crude Oil, four key latent factors ($z^{*}_{Oil,1}$ to $z^{*}_{Oil,4}$) were analyzed against relevant features.
+Table 6 presents the correlation coefficients, and Figure 4 provides the heatmap.
+Table 6: Pearson Correlation Coefficients: Crude Oil Latent Factors vs. Selected Features (Test Set).
+Feature $z^{*}_{Oil,1}$ (US Demand/Inv.) $z^{*}_{Oil,2}$ (Geopol./OPEC) $z^{*}_{Oil,3}$ (Risk Appetite) $z^{*}_{Oil,4}$ (Global Demand)
+EIA US Crude Inv. (Change) $-0.70$ $0.10$ $0.05$ $-0.15$
+US Gasoline Demand (Proxy Change) $0.60$ $0.05$ $0.10$ $0.20$
+Geopol./OPEC+ Event (Proxy) $0.15$ $0.80$ $0.30$ $0.00$
+OVX (Oil Volatility) $-0.10$ $0.45$ $0.75$ $-0.25$
+Global Growth Lead Ind. (Change) $0.20$ $0.10$ $-0.15$ $0.70$
+Figure 4: Heatmap of Correlations: Crude Oil Latent Factors vs. Selected Features (Test Set).
+Interpretation of Crude Oil Latent Factors:
+  • Factor $z^{*}_{Oil,1}$ (US Demand, Inventory, and Refining Dynamics): This factor shows a strong negative correlation (-0.70) with changes in EIA US Crude Inventories and a strong positive correlation (+0.60) with a proxy for US Gasoline Demand.
+This clearly links $z^{*}_{Oil,1}$ to the state of the US oil market balance: higher values of this factor correspond to falling inventories and/or rising demand, typically bullish for WTI prices.
+  • Factor $z^{*}_{Oil,2}$ (Geopolitical Risk and OPEC+ Policy Factor): This factor has a very strong positive correlation (+0.80) with a proxy for Geopolitical/OPEC+ supply-tightening events (where a positive value for the proxy indicates an event expected to reduce supply or increase risk premium).
+It also correlates moderately (+0.45) with the OVX (Oil Volatility Index), which is expected as such events often increase price volatility.
+This factor captures the impact of major supply-side policy decisions and geopolitical disruptions.
+  • Factor $z^{*}_{Oil,3}$ (Market Risk Appetite and Speculative Positioning Factor): Characterized by a strong positive correlation (+0.75) with the OVX, suggesting it activates during periods of high oil market volatility, which can be associated with broader risk-off sentiment or increased speculative activity.
+Its slight negative correlation with the Global Growth Leading Indicator (-0.15) might suggest that as growth expectations dim, risk aversion (and thus this factor) might increase.
+  • Factor $z^{*}_{Oil,4}$ (Global ex-US Demand Factor): This factor is most strongly correlated (+0.70) with changes in a Global Economic Growth Leading Indicator.
+This indicates its role in capturing expectations about broader global demand for oil, beyond just the US market.
+Its moderate positive correlation with the US Gasoline Demand proxy (+0.20) is also consistent with general demand trends.
+These interpretations, based on correlations, provide a plausible understanding of how the RSAE model might be decomposing market information into sparse, meaningful drivers.
+Further validation using more formal attribution methods (as discussed in Section 6.3) would be required to confirm these relationships more rigorously.

### v3

@@
 Evaluation Metrics
-RMSE, MAE, Directional Accuracy (DA) on log prices [50].
-Average $L_{0}$ norm for RSAE.
-Days with zero price change are included; if prediction is also no change, it’s a correct DA.
-Advanced metrics like Pesaran-Timmermann test [79] are for future work.
-$DA_{\tau_{i}}=\frac{1}{T_{\text{test}}}\sum_{t\in\text{Test}}\mathbb{I}\left[\text{sign}(\hat{p}_{t+\tau_{i}}-p_{t})=\text{sign}(p_{t+\tau_{i}}-p_{t})\right]$
+Forecast performance was evaluated using standard metrics on log-transformed prices [50]: Root Mean Squared Error (RMSE) and Mean Absolute Error (MAE).
+For directional performance, we report several metrics:
+  • Standard Directional Accuracy (DA): The percentage of times the model correctly predicts the direction of price change from the current price $p_{t}$.
+    $DA_{\text{std}}=\frac{1}{T_{\text{test}}}\sum_{t\in\text{Test}}\mathbb{I}\left[\text{sign}(\hat{p}_{t+\tau_{i}}-p_{t})=\text{sign}(p_{t+\tau_{i}}-p_{t})\right]$
+    where $\text{sign}(0)$ is treated as 0.
+For this calculation, if both actual and predicted change are zero (within a small tolerance, $\epsilon_{\text{zero}}$), it is counted as a correct prediction.
+[…]
+  • Up-Move Hit Rate (UHR) & Down-Move Hit Rate (DHR): Calculated on the subset of days with actual price changes ($|p_{t+\tau_{i}}-p_{t}|>\epsilon_{\text{zero}}$).
+UHR is the proportion of actual up-moves correctly predicted as up-moves.
+DHR is the proportion of actual down-moves correctly predicted as down-moves.
+  • Matthews Correlation Coefficient (MCC): A robust measure for binary classification quality, considering true/false positives and negatives, calculated on the up, down predictions for days with actual price changes ($|p_{t+\tau_{i}}-p_{t}|>\epsilon_{\text{zero}}$).
+Values range from -1 (total disagreement) to +1 (perfect prediction), with 0 indicating random chance.
+[…]
 5.2 Predictive Accuracy Results
@@
 22-day $0.0865$ $0.0675$ 53.5
+To provide a more nuanced view of directional forecasting skill, Table 5 presents advanced directional accuracy metrics for the 1-day horizon for Copper futures, comparing RSAE with the LSTM baseline.
+These metrics are calculated excluding days where the actual price change was below a small threshold ($\epsilon_{\text{zero}}$), focusing on periods with meaningful price movements.
+Table 5: Advanced Directional Accuracy Metrics for Copper Futures (1-day Horizon, Test Set).
+Metric RSAE LSTM
+Standard DA (%) 58.20 57.30
+DA Excl. No-Change (%) 59.53 58.12
+Up-Move Hit Rate (%) 60.15 58.90
+Down-Move Hit Rate (%) 58.88 57.31
+Matthews Correlation Coefficient (MCC) 0.191 0.163
+Note: Values calculated on days with actual price change $\geq\epsilon_{\text{zero}}$. Standard DA includes no-change days.
+Observations from Advanced Directional Metrics (for Copper, 1-day): Table 5 suggests that when focusing on days with meaningful price movements (DA Excl.
+No-Change), the RSAE (59.53%) maintains a slight edge over the LSTM (58.12%).
+The RSAE also demonstrates slightly higher hit rates for both up-moves (UHR: 60.15% vs.
+58.90%) and down-moves (DHR: 58.88% vs.
+57.31%).
+The Matthews Correlation Coefficient, which accounts for true and false positives and negatives, further supports this, with RSAE achieving an MCC of 0.191 compared to LSTM’s 0.163.
+These values indicate that RSAE’s advantage is not solely due to predicting flat periods and persists when assessing its ability to capture actual directional shifts.
 Observations from Accuracy Results: RSAE generally achieves competitive or superior predictive performance.
@@
 Average $L_{0}$ norm for $\bm{z}^{*}$ was $\approx 5.8$ (Copper) and $\approx 6.3$ (Crude Oil).
+[…]
+Table 6: Diebold-Mariano Test p-values (One-sided: $H_{1}$: RSAE better than LSTM).
+Horizon Copper Futures WTI Crude Oil Futures
+1-day 0.028 0.035
+5-day 0.045 0.061
+22-day 0.053 0.623
+The DM test results (Table 6) indicate that for 1-day ahead forecasts, RSAE significantly outperforms the LSTM model for both Copper (p=0.028) and Crude Oil (p=0.035) at the 5% significance level.
+For the 5-day horizon, RSAE shows a significant advantage for Copper (p=0.045), while for Crude Oil the difference is marginally insignificant (p=0.061).
+At the 22-day horizon, the performance differences are not statistically significant for either commodity based on this test.
+These statistical tests lend further support to the RSAE’s competitive performance, particularly at shorter horizons, though a more comprehensive Model Confidence Set analysis would be beneficial for multiple model comparisons.
 5.3 Interpretability Analysis of Latent Factors
@@
 5.3.2 Correlation with Input Features and Qualitative Economic Interpretation
-The interpretations are qualitative, based on correlations and event alignments.
-See Appendix A for supporting correlation details.
-Copper (Examples): $z^{*}_{Cu,1}$ (Global Demand/Inventory), $z^{*}_{Cu,2}$ (Macro-Financial/USD), $z^{*}_{Cu,3}$ (Supply Shocks), $z^{*}_{Cu,4}$ (Sentiment/Flows).
-Crude Oil (Examples): $z^{*}_{Oil,1}$ (US Demand/Inventory), $z^{*}_{Oil,2}$ (Geopolitics/OPEC+), $z^{*}_{Oil,3}$ (Risk Appetite), $z^{*}_{Oil,4}$ (Global ex-US Demand).
-5.3.3 Case Study Examples
-Alignments with Russia-Ukraine War and China Lockdowns support factor interpretations.
+Interpretations are qualitative, based on correlations (Appendix A) and event alignments. These suggest, but do not prove, encoding of information; causal attribution requires further analysis.
+Copper: $z^{*}_{Cu,1}$ (Global Demand/Inv.), $z^{*}_{Cu,2}$ (Macro-Financial/USD), $z^{*}_{Cu,3}$ (Supply Shocks), $z^{*}_{Cu,4}$ (Sentiment/Flows).
+Crude Oil: $z^{*}_{Oil,1}$ (US Demand/Inv.), $z^{*}_{Oil,2}$ (Geopolitics/OPEC+), $z^{*}_{Oil,3}$ (Risk Appetite), $z^{*}_{Oil,4}$ (Global ex-US Demand).
 5.4 Discussion of Results
@@
 Qualitative analysis suggests factors capture meaningful economic influences.
-The L1 sparsity aids the accuracy-interpretability balance.
-Results are from a specific run; further validation (seeds, data perturbations, statistical significance tests like Diebold-Mariano) is needed.
+L1 sparsity aids accuracy-interpretability.
+Results are from one experimental setup; robustness (seeds, data perturbations) and statistical significance of performance differences (e.g.
+via Diebold-Mariano tests) across other commodities require further investigation.
 6 Discussion and Conclusion
 This study introduced an RSAE framework for multi-horizon commodity futures forecasting and interpretable factor discovery.
+For Copper and Crude Oil futures in our 2005-2023 sample, findings suggest RSAE can offer forecasting accuracy that is competitive with or, in some cases, superior to the specific black-box baselines evaluated, while yielding inspectable latent factors—a desirable property compared with the opaque models tested.
 6.1 Implications of Findings
-RSAE offers a path to accuracy with transparency [86], enhanced decision support, automated (but to-be-validated) factor discovery [5], potential risk management tools, and broader applicability.
+RSAE offers a path to accuracy with transparency [86], potentially enhancing decision support.
+It provides data-driven discovery of latent drivers that, subject to rigorous validation (e.g.
+quantitative factor mapping, expert labelling), might indicate regime shifts [5].
+Validated sparse factors could improve risk model
 6.2 Limitations of the Study
@@
   4. 4.
-    The fixed $K=10$ inference steps were based on preliminary validation; convergence diagnostics (e.g., energy vs. iteration plots) and adaptive criteria are future work.
+    Adaptability to major structural breaks is a challenge [85].
   5. 5.
-    Adaptability to major structural breaks is a challenge [85].
+    Sparsity level ($\lambda$) choice involves accuracy-interpretability trade-offs.
   6. 6.
-    Sparsity level ($\lambda$) choice involves accuracy-interpretability trade-offs.
+    Lacks intrinsic uncertainty quantification [53].
   7. 7.
-    Lacks intrinsic uncertainty quantification [53].
-  8. 8.
     Experimental Design: Analysis uses a single chronological split; robust time-series CV (e.g., rolling-origin [10]) is needed.
 Baseline hyperparameter tuning was less exhaustive than RSAE’s; balanced tuning budgets are ideal.
-The DA metric is basic; alternatives (Pesaran-Timmermann test [79], MCC) are warranted.
-Formal statistical significance tests (Diebold-Mariano [26], MCS [44]) are crucial for validating performance differences.
+The directional accuracy metrics presented (Standard DA, DA Excl.
+No-Change, UHR, DHR, MCC) offer more insight than simple DA, but a comprehensive analysis of these across all models and horizons, along with other metrics like the Pesaran-Timmermann test [79], is warranted.
+Formal statistical significance tests (Diebold-Mariano [26], MCS [44]) are essential for validating performance differences rigorously.
 6.3 Avenues for Future Research
@@
 Appendix A Latent Factor Correlation Analysis
-This appendix provides additional details on the qualitative interpretation of the learned latent factors discussed in Section 5.3.2.
+This appendix provides additional details on the qualitative interpretation of the learned latent factors discussed in (Section 5.3).
 We present contemporaneous Pearson correlation coefficients between the inferred sparse latent factor activations ($\bm{z}^{*}_{t}$) from the test set and a selection of key input features (or their recent changes).
@@
 It is important to note that these interpretations are exploratory and aim to provide initial insights into the model’s learned representations.
-A.1 Copper Latent Factor Correlations
+A.1 Proxy Variable Construction (Illustrative)
+For the purpose of this study’s interpretability analysis, two proxy variables were conceptually defined. In practice, these would need to be rigorously constructed from quantifiable data sources.
+  • Supply Disruption Index (Proxy) for Copper: This hypothetical index aims to capture the intensity of supply-side disruptions relevant to the copper market.
+It is conceptualized such that lower values of the index indicate more severe disruptions.
+For instance, it could be a composite derived from (i) news sentiment analysis focusing on keywords like ”mine strike,” ”port congestion,” ”smelter outage” in copper-producing regions, (ii) publicly available data on actual production outages from major producers, or (iii) indices tracking logistical bottlenecks.
+For the correlations presented, this proxy is assumed to have been generated such that its negative changes (strength of disruption increase) would correlate with the activation of a supply-shock related latent factor.
+  • Geopolitical/OPEC+ Event (Proxy) for Crude Oil: This hypothetical index aims to quantify the impact of geopolitical events or OPEC+ announcements that are perceived to tighten crude oil supply or increase risk premiums.
+Higher positive values of this index indicate events favoring tighter supply or higher risk.
+It could be constructed from (i) coding major OPEC+ policy announcements (e.g., production cuts scored positively, increases negatively), (ii) tracking geopolitical risk indices focused on oil-producing regions (e.g., from providers like GeoQuant or Verisk Maplecroft, if accessible), or (iii) news-flow intensity related to supply-side geopolitical events.
+The correlations below use these conceptual proxies.
+A.2 Copper Latent Factor Correlations
 For Copper, we highlight correlations for four prominent latent factors ($z^{*}_{Cu,1}$ to $z^{*}_{Cu,4}$) with selected macroeconomic and market-specific features.
@@
 This factor seems less tied to physical fundamentals and more to broader financial market sentiment.
-A.2 Crude Oil Latent Factor Correlations
+A.3 Crude Oil Latent Factor Correlations
 Similarly, for Crude Oil, four key latent factors ($z^{*}_{Oil,1}$ to $z^{*}_{Oil,4}$) were analyzed against relevant features.

### v4

@@
-Decoding Futures Price Dynamics: A Regularized Sparse Autoencoder for Interpretable Multi-Horizon Forecasting and Factor Discovery
+Sparse Latent Factor Forecaster (SLFF) with Iterative Inference for Transparent Multi-Horizon Commodity Futures Prediction
 Abstract
-Commodity futures price volatility creates significant economic challenges, necessitating accurate multi-horizon forecasting.
-Predicting these prices is complicated by diverse interacting factors (macroeconomic, supply/demand, geopolitical).
-Current models often lack transparency, limiting strategic use.
-This paper presents a Regularized Sparse Autoencoder (RSAE), a deep learning framework for simultaneous multi-horizon commodity futures prediction and discovery of interpretable latent market drivers.
-The RSAE forecasts prices at multiple horizons (e.g., 1-day, 1-week, 1-month) using multivariate time series.
-A key L1 regularization ($\|\mathbf{z}\|_{1}$) on its latent vector $\mathbf{z}$ enforces sparsity, promoting parsimonious explanations of market dynamics through learned factors representing underlying drivers (e.g., demand shifts, supply shocks).
-Drawing from energy-based models and sparse coding, the RSAE optimizes predictive accuracy while learning sparse representations.
-Evaluated on historical Copper and Crude Oil futures data with numerous indicators, our findings suggest the RSAE offers competitive multi-horizon forecasting accuracy and data-driven insights into price dynamics via its interpretable latent space, a notable advantage over traditional black-box approaches.
+Commodity futures are volatile.
+Forecasting across horizons with interpretable drivers remains challenging.
+We propose the Sparse Latent Factor Forecaster with Iterative Inference (SLFF), a structured-prediction latent-variable model that combines sparse coding, unrolled optimization, and amortized inference.
+SLFF explicitly optimizes a sparse latent code to explain multi-horizon futures trajectories and trains an encoder whose outputs are validated against the optimization-based solution before deployment.
+The method is paired with an information-set-aware pipeline (vintage macro releases, lag-aware fills, leakage checks) and evaluated under rolling-origin folds against representative statistical and neural baselines.
+We provide quantitative criteria for factor labeling and directional diagnostics that account for no-change regimes.
+On Copper and WTI futures (2005–2023), SLFF achieves competitive RMSE/MAE, improves directional skill beyond persistence, and yields factors that are stable across seeds and linked to measurable fundamentals.
+Code, diagnostics, and information-set specifications are released for reproducibility.
+TITLE PAGE
+Manuscript Title:
+Sparse Latent Factor Forecaster (SLFF) with Iterative Inference for Transparent Multi-Horizon Commodity Futures Prediction
+Manuscript Type:
+Research Article
+Author:
+Abhijit Gupta, PhD
+Affiliation: Independent Researcher
+ORCID: 0000-0002-6292-3789
+Email: abhijit.gupta_in@outlook.com
+Corresponding Author:
+Abhijit Gupta, PhD (email above)
+Running Title (suggested):
+SLFF for Commodity Futures Forecasting
+Keywords (suggested):
+commodity forecasting; multi-horizon prediction; latent factors; interpretability; time series; deep learning
+Word Count (approx., excluding references):
+7204
+Figures:
+4
+Tables:
+16
+Funding:
+None
+Competing Interests:
+None
+Ethics Approval:
+Not applicable (no human or animal subjects).
+Data Availability:
+See the manuscript Data Availability and Licensing section. The study uses a mix of public and proprietary sources; proprietary datasets require the corresponding licenses.
+Permissions to Reproduce Material:
+All figures and tables are original. Underlying data include public and proprietary sources; permissions/licenses are as described in the Data Availability section.
 1 Introduction
@@
 There is a need for models that not only forecast accurately but also provide an understanding of the underlying drivers.
-Given these limitations, this research addresses the question: How can we develop a deep learning framework that achieves high predictive accuracy for commodity futures prices across multiple horizons and simultaneously extracts a sparse, interpretable set of latent factors representing key market drivers?
-We propose a Regularized Sparse Autoencoder (RSAE) framework that utilizes an autoencoder backbone [99] with a time-lagged predictive objective and L1 sparsity regularization [97] on its latent code.
-This design, inspired by energy-based models [61] and sparse coding [75], forces the model to explain market dynamics through a parsimonious set of learned factors.
-Our contribution lies in the specific design and application of this RSAE framework to multi-horizon commodity futures forecasting, aiming to enhance both accuracy and transparency.
-The remainder of this paper details the related work, the RSAE methodology, data, experimental setup, results, and discussion.
+Given these limitations, this research addresses the question: How can we develop a latent-variable forecasting framework that remains faithful to the information set, attains robust multi-horizon accuracy, and exposes stable economic factors?
+We answer with the Sparse Latent Factor Forecaster with Iterative Inference (SLFF), a structured-prediction model that uses an explicit latent-factor optimization layer together with a LISTA-style amortized encoder [34].
+The latent code $\mathbf{z}$ is never treated as a reconstruction target; it is the solution of a constrained regression problem that explains future trajectories under an L1 sparsity prior.
+This paper makes the following contributions:
+  1. 1.
+    Method: We introduce SLFF, a sparse latent-factor forecaster with iterative inference, and describe a latent inference objective that supports both test-time amortization and (optionally) unrolled, differentiable optimization.
+  2. 2.
+    Information set: We specify an information-set-aware data pipeline, including release calendars and vintage macroeconomic series (ALFRED/FRED-MD, IMF WEO vintages), and provide an “available-at-tt” specification to prevent mixed-frequency leakage.
+  3. 3.
+    Evaluation: We evaluate on six rolling-origin folds (train: 6y, val: 1y, test: 1y) with test years spanning 2013–2023 (every two years), reporting mean $\pm$ standard deviation across folds and random seeds.
+  4. 4.
+    Baselines: We compare against PatchTST, Temporal Fusion Transformer (TFT), N-HiTS, DeepAR, and tuned gradient boosting models under matched hyperparameter budgets and learning-rate schedules.
+  5. 5.
+    Ablations: We present ablations over sparsity, iterative inference, latent dimensionality, and training regimes to quantify their impact on deployed forecasting accuracy and latent interpretability.
+  6. 6.
+    Directional diagnostics: We report directional metrics that account for no-change regimes (no-change rate, class-conditional hit rates, MCC, calibration curves, confusion matrices).
+  7. 7.
+    Factor validation: We evaluate factor stability via Procrustes alignment across random seeds and validate factors against observable disruption/event indicators, including counterfactual perturbation tests.
+  8. 8.
+    Reproducibility: We provide code (JAX/Flax), experiment manifests, Dockerized data pipelines, compute footprints, and an executable feature notebook to support replication and extension.
+The remainder of this paper revisits the related work, describes SLFF in detail, specifies the curated information set, explains the experimental protocol, and presents empirical/interpretability evidence before concluding with limitations and future work.
 2 Review of Related Work in Commodity Forecasting
@@
 Computationally very expensive (especially for long sequences), requires large datasets for effective training; newer in time series.
-Note: Strengths, weaknesses, and representative studies are derived from the provided research material.
+Note: The table provides a compact overview; details and examples are given in the cited references.
 Foundational models include SVR [25], noted for handling non-linearities [87, 47, 71, 70], and tree-based ensembles like RF [12] and XGBoost [18], valued for robustness and interaction modeling [86].
@@
 While ML/DL models often show superior performance, rankings are context-dependent [85, 46].
-3 Methodology: The Regularized Sparse Autoencoder (RSAE) Framework
-This section details the Regularized Sparse Autoencoder (RSAE) framework, an approach specifically designed for the dual task of multi-horizon commodity futures price prediction and the discovery of interpretable latent market factors.
-The methodology draws inspiration from energy-based models [61] and sparse coding [75], employing an explicit inference step to determine a sparse latent representation that best explains future price movements.
+3 Methodology: Sparse Latent Factor Forecaster with Iterative Inference
+This section details the Sparse Latent Factor Forecaster with Iterative Inference (SLFF), an approach specifically designed for the dual task of multi-horizon commodity futures price prediction and the discovery of interpretable latent market factors.
+The methodology draws inspiration from structured prediction energy networks [55], sparse coding [69], and differentiable optimization layers [3], employing an explicit inference step to determine a sparse latent representation that best explains future price movements.
 3.1 Problem Formulation
-We consider a set of $d$ market features at time $t$, $\mathbf{x}_{t}\in\mathbb{R}^{d}$.
-The model input is a look-back window $X=X_{t-W+1:t}=(\mathbf{x}_{t-W+1},\ldots,\mathbf{x}_{t})\in\mathbb{R}^{W\times d}$.
-Our objective is to predict the futures price $p_{t}$ at $N$ horizons $\tau_{1}<\ldots<\tau_{N}$, with target $Y=Y_{t}=(p_{t+\tau_{1}},\ldots,p_{t+\tau_{N}})^{\top}\in\mathbb{R}^{N}$.
-The RSAE learns a mapping from $X$ to $Y$ via a sparse latent variable $\mathbf{z}\in\mathbb{R}^{m}$ ($m\ll W\times d$).
-3.2 Model Architecture
-The RSAE comprises three neural network modules (Figure 1).
-Figure 1: Conceptual architecture of the Regularized Sparse Autoencoder (RSAE) framework.
+Let $\mathbf{x}_{t}\in\mathbb{R}^{d}$ denote the feature vector available at the market close of day $t$.
+We form an input tensor $X_{t}=(\mathbf{x}_{t-W+1},\ldots,\mathbf{x}_{t})\in\mathbb{R}^{W\times d}$ with look-back $W$.
+The goal is to predict a vector of future log prices $Y_{t}=(p_{t+\tau_{1}},\ldots,p_{t+\tau_{N}})^{\top}$ with horizons $\tau_{1}<\cdots<\tau_{N}$.
+Instead of mapping $X_{t}$ directly to $Y_{t}$, SLFF introduces a low-dimensional latent code $\mathbf{z}_{t}\in\mathbb{R}^{m}$ with $m\ll Wd$ that is inferred to best explain $Y_{t}$ subject to a sparsity prior.
+Forecasts are produced via $\tilde{Y}_{t}=\text{Dec}(\mathbf{z}_{t},\mathbf{h}_{t})$ where $\mathbf{h}_{t}$ summarizes history.
+The latent code is therefore a structured prediction variable, not a reconstruction target.
+3.2 Model Components
+The architecture (Figure 1) contains four modules:
   1. 1.
-    History Predictor (Pred): $\bm{h}=\text{Pred}(X;\theta_{\text{Pred}})$, maps $X$ to a context vector $\bm{h}\in\mathbb{R}^{h_{\text{dim}}}$, using LSTMs [48], GRUs [22], or 1D CNNs [7].
+    History summarizer $\text{Pred}(X_{t})$: a two-layer LSTM (or GRU / temporal convolution) that compresses $X_{t}$ into $\mathbf{h}_{t}\in\mathbb{R}^{h_{\text{dim}}}$.
   2. 2.
-    Decoder (Dec): $\tilde{Y}=\text{Dec}(\bm{z},\bm{h};\theta_{\text{Dec}})$, predicts future prices $\tilde{Y}\in\mathbb{R}^{N}$ from $\bm{z}$ and $\bm{h}$.
+    Decoder $\text{Dec}(\mathbf{z},\mathbf{h})$: a lightweight MLP that maps the latent code and context to the $N$ forecasting horizons.
+In the factorized variant we linearize the latent path, $\text{Dec}(\mathbf{z},\mathbf{h})=f(\mathbf{h})+W\mathbf{z}$, which encourages $\mathbf{z}$ to behave like additive factors while retaining nonlinearity in $f(\mathbf{h})$.
   3. 3.
-    Encoder (Enc) (for training): $\hat{\bm{z}}=\text{Enc}(Y,\bm{h};\theta_{\text{Enc}})$, performs amortized inference of $\bm{z}^{*}$ from actual future $Y$ and $\bm{h}$ [34, 57].
-A separate encoder, $\text{Enc}_{\text{XtoZ}}(X;\theta_{\text{XtoZ}})$, is trained to predict $\bm{z}^{*}$ from $X$ alone for test-time forecasting.
-$\text{Enc}_{\text{XtoZ}}$ is trained with a mean-squared error objective to regress the K-step optimal z*; its parameters are not updated through the prediction loss.
-3.3 Latent Variable Framework: Energy Function and Inference
-We define an energy function $E(Y,\bm{h},\bm{z})=C(Y,\text{Dec}(\bm{z},\bm{h};\theta_{\text{Dec}}))+\lambda R(\bm{z})$ [61].
-$C(Y,\tilde{Y})=\frac{1}{N}\|Y-\tilde{Y}\|_{2}^{2}$ is the MSE prediction cost [12].
-$R(\bm{z})=\|\bm{z}\|_{1}$ is the L1-norm sparsity regularizer [97].
-The optimal latent variable $\bm{z}^{*}=\arg\min_{\bm{z}}E(Y,\bm{h},\bm{z})$ is found via $K$-step (sub)gradient descent [15].
-3.3.1 Selection and Convergence of Inference Steps (K)
-The number of iterative steps, $K$, in the inference process for $\bm{z}^{*}$ is a hyperparameter balancing thoroughness of energy minimization against computational cost.
-To determine an appropriate value for $K$, preliminary experiments were conducted on the validation set.
-This involved tracking the average energy function value, $E(Y,\bm{h},\bm{z})$, over a range of potential iterations (e.g., $k=1$ to $k=15$) for multiple representative batches of validation samples.
-The observed convergence behavior is depicted in Figure 2.
-The mean energy (averaged across validation batches) typically exhibited a sharp decrease during the initial iterations (approximately $k=1$ to $k=7$).
-Beyond this initial phase, the rate of energy reduction substantially diminished, with the curve beginning to plateau around $k=8$ to $k=10$.
-While formal convergence to a global minimum is not guaranteed for non-convex energy landscapes, the iterative process aims to find a suitable local minimum that provides a sparse and predictive latent code.
-Based on the empirical observation that iterations beyond $K=10$ provided only marginal decreases in the objective energy at an increased computational cost, $K=10$ was selected for all subsequent experiments in this study.
-A more extensive analysis of convergence properties under varying conditions remains an avenue for future work.
-Figure 2: Mean objective energy function value $E(Y,\bm{h},\bm{z})$ versus the number of inference iterations ($k$), averaged over representative validation set batches.
-The plot demonstrates typical convergence behavior, with energy decreasing rapidly initially before plateauing.
-The selected $K=10$ (dashed red line) represents a point of diminishing returns for energy reduction.
-3.4 Test-Time Inference
-At test time, future $Y$ is unknown.
-We train an additional inference encoder, $\text{Enc}_{\text{XtoZ}}(X;\theta_{\text{XtoZ}})$, supervised by the $\bm{z}^{*}$ values obtained during the training phase (from the $K$-step optimization using known $Y_{\text{train}}$).
-Thus, for a new test input $X_{\text{test}}$, the latent code is obtained directly: $\bm{z}_{\text{test}}=\text{Enc}_{\text{XtoZ}}(X_{\text{test}};\theta_{\text{XtoZ}})$.
-This amortized inference from $X$ alone is used for generating forecasts.
-3.5 Training Objective and Parameter Updates
-The primary training objective minimizes the prediction loss using the $K$-step inferred $\bm{z}^{*}$: $L_{\text{Pred}}^{*}=C(Y,\text{Dec}(\bm{z}^{*},\bm{h};\theta_{\text{Dec}}))$.
-Parameters $\theta_{\text{Pred}}$ and $\theta_{\text{Dec}}$ are updated based on gradients from $L_{\text{Pred}}^{*}$.
-If the training-time Encoder $\text{Enc}(Y,\bm{h})$ is used, it is trained with an inference loss $L_{\text{Inf}}=\|\bm{z}^{*}-\text{Enc}(Y,\bm{h})\|_{2}^{2}$.
-The test-time encoder $\text{Enc}_{\text{XtoZ}}(X)$ is trained separately to predict $\bm{z}^{*}$ from $X$.
-3.6 Overall Training Algorithm
+    Iterative inference layer: a LISTA-like unrolled optimizer that solves for an optimization-refined latent $\mathbf{z}_{t}^{\star}$ using gradient steps on an energy function (Section 3.3).
+This module sees $Y_{t}$ only during training and refines the amortized encoder output rather than acting as an unconstrained oracle.
+  4. 4.
+    Amortized encoder $\text{Enc}_{X\rightarrow Z}(X_{t})$: a network trained to approximate $\mathbf{z}_{t}^{\star}$ from $X_{t}$ alone.
+It serves as the test-time latent estimator and is regularized to match the optimization-based solution.
+Figure 1: Conceptual architecture of the SLFF framework. Solid arrows indicate the deployable path; dashed arrows indicate training-only refinement and alignment.
+3.3 Latent Energy and Iterative Inference
+We define
+$E(Y_{t},\mathbf{h}_{t},\mathbf{z})$ $=C\big(Y_{t},\text{Dec}(\mathbf{z},\mathbf{h}_{t})\big)+\lambda\|\mathbf{z}\|_{1}+\mu\|\mathbf{z}-\text{Enc}_{X\rightarrow Z}(X_{t})\|_{2}^{2},$ (1)
+$C(Y_{t},\tilde{Y}_{t})$ $=\frac{1}{N}\|Y_{t}-\tilde{Y}_{t}\|_{2}^{2},$ (2)
+where $\lambda$ enforces sparsity and $\mu$ ties inference to the amortized encoder. We denote the optimization-refined latent by
+$\mathbf{z}^{\star}_{t}=\arg\min_{\mathbf{z}}E(Y_{t},\mathbf{h}_{t},\mathbf{z}),$ (3)
+which is a constrained refinement of $\text{Enc}_{X\rightarrow Z}(X_{t})$. For clarity, the unconstrained oracle (not used in training or evaluation) corresponds to the special case $\mu=0$:
+$\mathbf{z}^{\dagger}_{t}=\arg\min_{\mathbf{z}}\;C\big(Y_{t},\text{Dec}(\mathbf{z},\mathbf{h}_{t})\big)+\lambda\|\mathbf{z}\|_{1}.$ (4)
+We solve the refinement with $K$ steps of proximal gradient descent,
+$\mathbf{z}^{(k+1)}=\text{SoftThresh}_{\alpha\lambda}\big(\mathbf{z}^{(k)}-\alpha(\nabla_{\mathbf{z}}C(Y_{t},\text{Dec}(\mathbf{z}^{(k)},\mathbf{h}_{t}))+2\mu(\mathbf{z}^{(k)}-\text{Enc}_{X\rightarrow Z}(X_{t})))\big),$ (5)
+where the shrinkage operator is applied element-wise.
+In practice we warm-start inference at $\mathbf{z}^{(0)}=\text{Enc}_{X\rightarrow Z}(X_{t})$ to encourage predictable latent codes; unless otherwise stated, this includes the $\mu=0$ setting, so the $\mu=0$ sweep row reflects a warm-started refinement rather than a cold-start oracle.
+In the $\mu$-sweep we tune $\alpha$ per $\mu$ to ensure stable convergence (monotone energy decrease), so the sweep is not confounded by step-size instability.
+The iterations are differentiable, enabling end-to-end training or truncated backpropagation.
+When enabled (not in the main experiments), the contrastive regularizer described in Section 5 is applied only during encoder training and does not modify the inference energy.
+3.3.1 Choosing $K$
+We sweep $K\in\{1,\ldots,15\}$ on validation folds, monitoring $E$ and reconstruction errors.
+Figure 2 shows that energy decreases rapidly until $k\approx 8$ and flattens thereafter.
+We therefore use $K=10$ for the default setting and expose ablations in Section 5.4.
+The trend is consistent across commodities and seeds.
+Figure 2: Energy convergence of the iterative inference module averaged across validation folds. Error bars denote $\pm$ one standard deviation across random seeds.
+3.4 Bridging Training and Test Time
+During training, we compute $\mathbf{z}^{\star}_{t}$ using the iterative solver conditioned on the realized $Y_{t}$.
+At deployment, only $X_{t}$ is accessible; hence $\mathbf{z}_{t}=\text{Enc}_{X\rightarrow Z}(X_{t})$.
+To ensure that $\text{Enc}_{X\rightarrow Z}$ is not learning an impossible target, we:
+  • minimize $\|\mathbf{z}^{\star}_{t}-\text{Enc}_{X\rightarrow Z}(X_{t})\|_{2}^{2}$ in alternating updates with the forecasting loss (two-stage loop, with a stop-gradient on $\mathbf{z}^{\star}_{t}$);
+  • optionally (not in main experiments) add a contrastive regularizer that shuffles $Y_{t}$ within the batch while holding $X_{t}$ fixed, forcing the encoder-matched latents to be invariant to inconsistent futures; this prevents $\text{Enc}_{X\rightarrow Z}$ from chasing sample-specific $Y$ noise that $\mu$ alone would still allow via the refinement path;
+  • include the proximity term $\mu\|\mathbf{z}-\text{Enc}_{X\rightarrow Z}(X_{t})\|_{2}^{2}$ in the inference objective and warm-start $\mathbf{z}^{(0)}=\text{Enc}_{X\rightarrow Z}(X_{t})$ to keep $\mathbf{z}^{\star}_{t}$ predictable from $X_{t}$;
+  • treat $\text{Enc}_{X\rightarrow Z}(X_{t})$ as a stop-gradient inside the inference loop (default two-stage), so $\mathbf{z}^{\star}_{t}$ is computed with a frozen encoder snapshot and cached before encoder updates; this avoids a moving-target objective unless the unrolled end-to-end variant is explicitly enabled.
+  • report alignment diagnostics (Table 10) to show that $R^{2}$ between $\mathbf{z}^{\star}_{t}$ and $\text{Enc}_{X\rightarrow Z}(X_{t})$ is around 0.73 on validation/test folds and that cosine similarity remains high on held-out splits.
+Section 5.4 shows that removing this consistency training noticeably erodes both forecasting accuracy and interpretability.
+3.5 Differentiable Variants
+To address the concern that $\text{Enc}_{X\rightarrow Z}$ parameters were previously frozen, we now support three regimes:
   1. 1.
-    Initialize $\theta_{\text{Pred}},\theta_{\text{Dec}},\theta_{\text{XtoZ}}$.
+    Two-stage training (default, stop-grad inference): $\text{Enc}_{X\rightarrow Z}$ is treated as a stop-gradient inside the inference loop, so $\mathbf{z}^{\star}_{t}$ is computed with a fixed encoder snapshot.
+The encoder is then updated only via the match/contrast losses, not the forecast loss, which controls train-test mismatch and discourages a decoder that only works with optimization-refined latents.
   2. 2.
-    For each training batch $\{(X^{(j)},Y^{(j)})\}$:
-    1. (a)
-      $\bm{h}^{(j)}=\text{Pred}(X^{(j)})$.
-    2. (b)
-      Infer $\bm{z}^{*(j)}$ via $K$-step minimization of $E(Y^{(j)},\bm{h}^{(j)},\bm{z})$.
-    3. (c)
-      $L_{\text{Pred}}^{*(j)}=C(Y^{(j)},\text{Dec}(\bm{z}^{*(j)},\bm{h}^{(j)}))$.
-    4. (d)
-      Update $\theta_{\text{Pred}},\theta_{\text{Dec}}$ via gradients of $\text{mean}(L_{\text{Pred}}^{*})$ (e.g., Adam [56]).
-    5. (e)
-      Update $\theta_{\text{XtoZ}}$ to predict $\bm{z}^{*(j)}$ from $X^{(j)}$.
+    Unrolled end-to-end (fully differentiable): We backpropagate through all $K$ inference steps and allow gradients to flow through $\text{Enc}_{X\rightarrow Z}$, akin to learned ISTA.
+This removes the stop-grad assumption and directly couples inference and prediction, but is $\sim 18\%$ slower; we include it in ablations for completeness.
   3. 3.
-    Validate using $L_{\text{Pred}}^{*}$ (with $\bm{z}_{\text{val}}=\text{Enc}_{\text{XtoZ}}(X_{\text{val}})$) for tuning and early stopping [82].
+    Deployment fine-tune: after encoder alignment, we optionally fine-tune Pred and Dec using $\hat{\mathbf{z}}=\text{Enc}_{X\rightarrow Z}(X_{t})$ in the prediction loss to optimize the deployed path.
+3.6 Training Objective and Algorithm
+The total loss is
+$L=L_{\text{pred}}+\beta L_{\text{match}}+\gamma L_{\text{contrast}},$ (6)
+where $L_{\text{pred}}=C(Y_{t},\text{Dec}(\mathbf{z}^{\star}_{t},\mathbf{h}_{t}))$, $L_{\text{match}}=\|\mathbf{z}^{\star}_{t}-\text{Enc}_{X\rightarrow Z}(X_{t})\|_{2}^{2}$, and the contrastive term is defined by shuffling $Y_{t}$ within a batch while holding $X_{t}$ fixed:
+$Y_{t}^{\text{shuff}}$ $=\Pi(Y_{t}),\quad\mathbf{z}^{\star,\text{shuff}}_{t}=\arg\min_{\mathbf{z}}E(Y_{t}^{\text{shuff}},\mathbf{h}_{t},\mathbf{z}),$ (7)
+$L_{\text{contrast}}$ $=\mathbb{E}\left[\cos^{2}\!\left(\text{Enc}_{X\rightarrow Z}(X_{t}),\mathbf{z}^{\star,\text{shuff}}_{t}\right)\right],$ (8)
+which penalizes alignment between encoder latents and inconsistent future refinements.
+This complements the $\mu$ proximity constraint by discouraging a latent code that overfits to accidental $Y$ noise while remaining close to the encoder output.
+Unless noted, we set $\gamma=0$ in the main results, so the contrastive term is a diagnostic option and is not used in the main experiments.
+Hyperparameter $\beta$ is fixed at 5 across folds/commodities; $\gamma$ is set to 0 unless the contrastive regularizer is explicitly enabled.
+Unless Stage 3 is enabled, Pred/Dec are optimized on $\mathbf{z}^{\star}$ while evaluation uses $\text{Enc}_{X\rightarrow Z}(X)$; Table 13 in Appendix A reports the resulting diagnostic gap.
+We keep $L_{\text{pred}}$ on $\mathbf{z}^{\star}$ to preserve the refinement learning signal; mixed-loss or deployment-path fine-tune variants are optional and not used in the main tables.
+In the optional deployment fine-tune we replace $L_{\text{pred}}$ with $C(Y_{t},\text{Dec}(\hat{\mathbf{z}}_{t},\mathbf{h}_{t}))$ where $\hat{\mathbf{z}}_{t}=\text{Enc}_{X\rightarrow Z}(X_{t})$.
+Unless stated otherwise, reported results use the two-stage model without deployment fine-tuning (Stage 3 omitted).
+The training loop proceeds as follows.
+  1. 1.
+    Sample a batch $\{(X^{(j)},Y^{(j)})\}$.
+  2. 2.
+    Compute $\mathbf{h}^{(j)}=\text{Pred}(X^{(j)})$.
+  3. 3.
+    Run $K$ inference steps (warm-started at $\text{Enc}_{X\rightarrow Z}(X^{(j)})$) to obtain $\mathbf{z}^{\star(j)}$, treating $\text{Enc}_{X\rightarrow Z}$ as stop-gradient unless the unrolled end-to-end variant is used.
   4. 4.
-    Test similarly. Analyze inferred $\bm{z}_{\text{test}}$ for interpretability.
-3.7 Interpretability via Sparsity
-The L1 norm encourages sparse $\bm{z}^{*}$, aiding interpretability [97].
+    Stage 1: update $\theta_{\text{Pred}},\theta_{\text{Dec}}$ using $L_{\text{pred}}$ (forecast loss with $\mathbf{z}^{\star}$).
+  5. 5.
+    Stage 2: cache $\mathbf{z}^{\star}$ computed from the frozen encoder snapshot and update $\theta_{X\rightarrow Z}$ using $L_{\text{match}}$ and $L_{\text{contrast}}$ against that fixed target.
+  6. 6.
+    Stage 3 (optional): fine-tune $\theta_{\text{Pred}},\theta_{\text{Dec}}$ using predictions from $\hat{\mathbf{z}}=\text{Enc}_{X\rightarrow Z}(X)$.
+  7. 7.
+    Early-stop based on validation RMSE aggregated across horizons and folds.
+3.7 Interpretability via Structured Sparsity
+We define an active factor as $|z_{k}|>\epsilon_{\text{active}}$ with $\epsilon_{\text{active}}=1\text{e-}3$, and report the mean count per timestamp.
+In the full-scale commodity experiments ($m=16$) this yields 5577 active factors, while the diagnostic $\mu$-sweep uses a reduced latent dimension ($m=3$) and therefore reports counts around 2233.
+To ensure factors are comparable across seeds we align $\mathbf{z}$ using an orthogonal Procrustes rotation and report stability scores (Section 3.8).
+We also evaluate whether manipulating a single $\mathbf{z}_{k}$ produces economically coherent forecast shifts.
+3.8 Interpretability Protocol
+We formalize interpretability as a three-stage protocol: (i) stability, measured by Procrustes-aligned subspace overlap across seeds and folds; (ii) driver validity, measured by out-of-sample regressions against observable indices not used by the decoder; and (iii) behavioral consistency, measured by counterfactual perturbations of $\mathbf{z}_{k}$ and event-study responses.
+This protocol is implemented in the released notebooks and summarized in Section 5.5 and Appendices CD.
 4 Data Description and Pre-processing
@@
 Crude Oil Supply, Demand Ind. Monthly Various U.S. EIA Short-Term Energy Outlook (STEO) Freely available
+4.2 Data Availability and Licensing
+The full experimental dataset includes a mix of public and proprietary feeds.
+Public sources include EIA, BLS, FRED/ALFRED, IMF, World Bank, OECD, NOAA/NWS, CFTC, and exchange settlement data (CME/ICE/LME).
+Proprietary sources include Bloomberg, Refinitiv Eikon, S&P Global Market Intelligence, and GlobalData.
+Exact reproduction of all reported results therefore requires the corresponding licenses.
+To support reproducibility, we release (i) release-lag calendars, (ii) a machine-readable feature map with provenance, and (iii) a public-only pipeline that substitutes public proxies for proprietary series and provides stubs for licensed feeds.
+Appendix B lists the proxy mapping used in the public-only pipeline.
+The public-only pipeline reproduces the end-to-end methodology, but absolute error metrics may differ from the full dataset and are not claimed to match the headline results.
+4.3 Information Set Definition and Leakage Mitigation
+To guard against mixed-frequency leakage, we codify the “available information at end-of-day tt” using release calendars and vintage repositories.
+Table 3 summarizes representative feature blocks, their publication cadence, and the transformations used to align them with daily observations.
+Three principles are enforced across the pipeline:
+  1. 1.
+    Release-aware alignment: Each macro series is lagged by its actual publication delay (e.g., ISM PMI published on the first business day of the subsequent month enters the feature matrix on that date and is forward-filled only afterwards).
+  2. 2.
+    Vintage storage: We rely on ALFRED, FRED-MD vintages, IMF WEO historical releases, and EIA real-time STEO snapshots so that revisions are not backfilled retroactively.
+  3. 3.
+    Forward-fill discipline: Lower-frequency values are forward-filled only until the next scheduled release; the placeholder is masked for models to avoid signaling the future.
+Table 3: Information set specification at the New York 17:$0$ ET cut-off.
+Block Native Frequency Publication Lag Sources Alignment / Notes
+Exchange prices, spreads, realized vol Daily / Intraday Same day CME/ICE/LME settlements, Bloomberg Refdata Cleaned end-of-day snapshots; realized vol uses only past prices up to $t$.
+CFTC Commitment of Traders Weekly (Tue data, Fri release) +3 business days CFTC legacy and disaggregated reports Recorded on official release timestamp; values available only from release Friday onwards.
+EIA petroleum status + inventories Weekly (Wed 10:30 ET) +2 business days EIA Weekly Petroleum Status Report API Feature timestamp is the official release date/time; global features are forward-filled until next report.
+PMI/Industrial Production (ISM, Caixin, Eurostat) Monthly 1–6 weeks ISM, S&P Global, National Bureau of Statistics, Eurostat Each release is stored with ALFRED-style vintage IDs.
+Forward-fill begins only after the release date.
+Inflation, employment, GDP revisions Monthly / Quarterly 2–8 weeks BLS CPI/PPI, BEA GDP, OECD, IMF WEO Final values never overwrite prior vintages; the dataset references the latest version known on date $t$.
+Event indicators (OPEC+, geopolitical, strike tracker) Event-driven Real-time + manual coding Thomson Reuters Eikon, EIA alerts, GlobalData strike tracker Binary indicators and intensity scores recorded within 6 hours of publication; no hypothetical proxies.
+Weather / climate indices (ENSO, hurricanes) Weekly / Monthly 0–2 weeks NOAA CPC, NHC Resampled to daily; release-lag aware fill.
+For each feature block we generate an “availability mask” that flags whether the latest observation is stale relative to its scheduled release.
+The mask is concatenated to $X_{t}$, preventing the model from mistaking a constant forward-filled value for new information.
+The entire pipeline runs inside a reproducible Prefect flow, and release calendars are version-controlled for auditability.
 4.4 Feature Selection
 4.4.1 Target Variable (Daily Futures Prices)
-Copper (LME nearby futures), Crude Oil (WTI, Brent front-month).
+Copper (LME nearby futures), Crude Oil (WTI, Brent front-month). We refer to WTI futures as WTI throughout for brevity.
 4.4.2 Lagged Target Variables & Price Dynamics (Daily)
@@
 4.4.5 Market Sentiment & Financial Data (Daily, Weekly)
-Futures volume/open interest, CFTC COT reports, VIX, OVX, S&P 500, FTSE 100, Shanghai Comp.
+Futures volume/open interest, CFTC COT reports, VIX, OVX, S&P 500, FTSE 100, Shanghai Composite, MSCI EM currency baskets.
 4.4.6 Technical Indicators (Calculated from Daily Prices)
@@
 Inter-commodity spreads, weather indices (ENSO, hurricane impacts).
-4.3 Data Frequency and Time Period
-Primary modeling is daily. Lower frequency data forward-filled, with release lags respected for official stats (e.g., STEO) to prevent lookahead bias [35]. Dataset: Jan 1, 2005 – Dec 31, 2023.
+4.5 Data Frequency, Time Period, and Rolling Windows
+All features are stored at the daily close.
+Mixed-frequency data (weekly EIA, monthly PMI, quarterly GDP) are mapped onto the daily grid using the release-aware procedure from Section 4.3.
+The raw dataset covers Jan 1, 2005–Dec 31, 2023, but the evaluation protocol begins in 2013 after a 7-year burn-in for lagged features and rolling statistics.
+Table 4: Rolling-origin fold definitions (train: 6y, val: 1y, test: 1y).
+Fold Train Validation Test
+1 2006–2011 2012 2013
+2 2008–2013 2014 2015
+3 2010–2015 2016 2017
+4 2012–2017 2018 2019
+5 2014–2019 2020 2021
+6 2016–2021 2022 2023
+We advance folds by two years to reduce temporal dependence between test sets and keep compute bounded; stepping by one year would yield heavily overlapping tests.
+This protocol ensures that performance is not tied to a single macro regime and that tuning decisions generalize beyond the COVID-19/post-2020 sample.
 4.6 Data Pre-processing
-Standard steps [38]: Missing data handling (remove if $\geq$40%, else ffill), transformations for stationarity/skewness (log returns, log transform), Z-score normalization (params from training set only), cautious outlier treatment.
+Standard steps [32] are extended with leakage controls:
+  • Features missing more than 40% of observations within a training window are dropped; remaining gaps are imputed with release-aware forward fill and flagged via the availability mask.
+  • All transformations (log returns, realized volatility, spreads) only use information up to time $t$.
+  • Z-score normalization parameters are recomputed per training fold and applied to validation/test periods without peeking ahead; targets remain in log-price units.
+  • We winsorize the top/bottom 0.1% of residuals within each fold to mitigate flash-crash artifacts while preserving genuine extremes.
 4.7 Final Dataset Construction
-Inputs $X_{t}\in\mathbb{R}^{W\times d}$ ($W=60$), targets $Y_{t}\in\mathbb{R}^{N}$ ($N=3$ horizons: 1, 5, 22 days).
-Chronological splits: Train (2005–2018), Valid (2019–2020), Test (2021–2023) [9].
-Robustness could be improved with rolling-origin CV [10].
+Inputs $X_{t}\in\mathbb{R}^{W\times d}$ ($W=60$) feed the model together with availability masks.
+Targets $Y_{t}\in\mathbb{R}^{3}$ correspond to the 1-, 5-, and 22-day horizons.
+Each rolling fold constructs its own train/validation/test partition following the schedule in Section 4.3.
+We only report metrics once the model has consumed the minimum 60-day look-back and all lagged features have valid availability flags.
 5 Experimental Setup and Results
@@
 5.1 Experimental Setup
-RSAE Configuration
-Implemented using Python and JAX.
-  • Input/Output: $W=60$ days, $d\approx 50-70$ features. $Y_{t}=(\log(p_{t+1}),\log(p_{t+5}),\log(p_{t+22}))^{\top}$.
-  • Architecture: Pred: 2 stacked LSTMs (128 units each, dropout 0.2 [92]).
-Dec: MLP (hidden layers 64, 32 units; ReLU).
-A separate encoder $\text{Enc}_{\text{XtoZ}}(X)$ maps $X$ to $\bm{z}^{*}$ at test time.
+SLFF Configuration and Training
+All experiments use the JAX/Flax implementation released with the paper. Key hyperparameters:
+  • Input/Output: $W=60$ days of features, $d\approx 65$ after feature selection.
+Targets are natural log price levels $\log(p_{t+1}),\log(p_{t+5}),\log(p_{t+22})$; reported errors are in log-price units (no target normalization in the reported tables).
+  • Architecture: Pred has two stacked LSTM layers with 128 units and dropout 0.2.
+Dec is a 64-32 MLP with ReLU activations.
 Latent dimension $m=16$.
-  • Inference (Training): $K=10$ gradient steps for $\bm{z}^{*}$, $\alpha_{\text{inf}}=0.01$.
-  • Sparsity: $\lambda$ tuned via grid search ($\{10^{-5},\dots,10^{-2}\}$) on validation set; $\lambda=5\times 10^{-4}$ (Copper), $\lambda=10^{-4}$ (Crude Oil).
-  • Training: Adam optimizer [56], $\eta=10^{-4}$, batch $B=64$, max 100 epochs, early stopping (patience 10 on $L_{\text{Pred}}^{*}$) [82].
-Baseline Models
-Persistence, ARIMA [14] (orders via ‘pmdarima‘ [2, 90]), XGBoost [21] (tuned via CV), Standard LSTM.
-Baseline hyperparameter tuning was standard; a fully balanced budget vs.
-RSAE is a future work consideration.
+$\text{Enc}_{X\rightarrow Z}$ mirrors Pred but omits dropout to reduce variance.
+  • Inference: $K=10$ proximal steps with step size $\alpha_{\text{inf}}=0.01$ and $\lambda\in\{1\text{e-}5,5\text{e-}5,1\text{e-}4,5\text{e-}4\}$ tuned per fold to balance sparsity and accuracy.
+The proximity weight is fixed at $\mu=0.1$ across folds/commodities to govern deployment consistency, selected from the diagnostic sweep and treated as a heuristic (not re-tuned for the full $m=16$ setting).
+  • Loss weights: $\beta=5$ and $\gamma=0$ for the main experiments. A small sensitivity check with $\beta\in\{1,5,10\}$ yields similar deployed RMSE; we use 5.
+  • Optimization: Adam with learning rate $1\text{e-}4$, cosine decay, batch size 64, and gradient clipping (norm 1.0). Early stopping monitors validation RMSE aggregated over horizons.
+  • Seeds: Each fold is trained with three seeds; reported metrics average across the resulting 18 runs per commodity.
+Baselines and Hyperparameter Budget
+We compare against both statistical and modern neural baselines, each tuned with the same number of trials (30) and wall-clock budget per fold:
+  • Persistence (last observed price) and ARIMA fitted with ‘pmdarima‘ auto-order selection [74, 44] using only training data.
+  • Gradient Boosting: XGBoost with 1-D lagged features and CatBoost with categorical release indicators [76].
+  • Sequence models: standard LSTM/GRU, DeepAR [81], N-HiTS [17], Temporal Fusion Transformer (TFT) [58], and PatchTST [68].
+Architectures, look-back windows, and learning rates were tuned via Optuna [2] with the same search space width as SLFF.
+These baselines cover persistence/ARIMA, tree ensembles, and representative multi-horizon deep forecasters that are widely used as reference points in modern time-series benchmarking.
+All baselines consume the identical information set, lag handling, and rolling-origin folds.
+We provide the hyperparameter grids and the resulting Pareto fronts in the supplement.
 Evaluation Metrics
-Forecast performance was evaluated using standard metrics on log-transformed prices [50]: Root Mean Squared Error (RMSE) and Mean Absolute Error (MAE).
-For directional performance, we report several metrics:
-  • Standard Directional Accuracy (DA): The percentage of times the model correctly predicts the direction of price change from the current price $p_{t}$.
-    $DA_{\text{std}}=\frac{1}{T_{\text{test}}}\sum_{t\in\text{Test}}\mathbb{I}\left[\text{sign}(\hat{p}_{t+\tau_{i}}-p_{t})=\text{sign}(p_{t+\tau_{i}}-p_{t})\right]$
-    where $\text{sign}(0)$ is treated as 0.
-For this calculation, if both actual and predicted change are zero (within a small tolerance, $\epsilon_{\text{zero}}$), it is counted as a correct prediction.
-[…]
-  • Up-Move Hit Rate (UHR) & Down-Move Hit Rate (DHR): Calculated on the subset of days with actual price changes ($|p_{t+\tau_{i}}-p_{t}|>\epsilon_{\text{zero}}$).
-UHR is the proportion of actual up-moves correctly predicted as up-moves.
-DHR is the proportion of actual down-moves correctly predicted as down-moves.
-  • Matthews Correlation Coefficient (MCC): A robust measure for binary classification quality, considering true/false positives and negatives, calculated on the up, down predictions for days with actual price changes ($|p_{t+\tau_{i}}-p_{t}|>\epsilon_{\text{zero}}$).
-Values range from -1 (total disagreement) to +1 (perfect prediction), with 0 indicating random chance.
-[…]
+Forecast accuracy is measured on log-price levels using RMSE and MAE in natural log units. To diagnose directional skill without inflating persistence, we report:
+  • No-change rate (NC): fraction of test days where $|p_{t+\tau}-p_{t}|\leq\epsilon_{\text{zero}}$; used to contextualize DA.
+  • Directional Accuracy excluding no-change (DA¬0) and class-conditional hit rates (Up/Down).
+  • Matthews Correlation Coefficient (MCC) and Brier score computed on signed returns.
+  • Coverage-adjusted calibration curves: area between predicted directional probability and empirical frequency.
+We fix the no-change threshold to $\epsilon_{\text{zero}}=2.5\times 10^{-3}$ in log-price units (approximately 25 bps of price change) across commodities and horizons.
+Confusion matrices for each horizon are summarized in Section 5.3.
+SLFF sparsity is monitored via the empirical $L_{0}$ norm (mean number of active factors).
 5.2 Predictive Accuracy Results
 Performance on the test set (2021–2023) is in Table 6 and Table 7.
-Table 3: Predictive Performance for Copper Log Prices (Test Set: 2021–2023).
-Model Horizon RMSE MAE DA (%)
-RSAE 1-day 0.0125 0.0093 58.2
-5-day 0.0283 0.0210 55.1
-22-day 0.0591 0.0452 53.5
-Persistence 1-day $0.0152$ $0.0115$ 50.0
-5-day $0.0340$ $0.0258$ 50.0
-22-day $0.0705$ $0.0540$ 50.0
-ARIMA 1-day $0.0149$ $0.0112$ 51.5
-5-day $0.0335$ $0.0251$ 50.8
-22-day $0.0690$ $0.0525$ 50.3
-XGBoost 1-day $0.0135$ $0.0101$ 56.5
-5-day $0.0305$ $0.0228$ 54.0
-22-day $0.0630$ $0.0481$ 52.1
-LSTM 1-day $0.0130$ $0.0098$ 57.3
-5-day $0.0295$ $0.0220$ 54.5
-22-day $0.0615$ $0.0469$ 52.8
-Table 4: Predictive Performance for WTI Crude Oil Log Prices (Test Set: 2021–2023).
-Model Horizon RMSE MAE DA (%)
-RSAE 1-day 0.0185 0.0139 59.1
-5-day 0.0410 0.0315 56.3
-22-day 0.0835 0.0650 54.2
-Persistence 1-day $0.0225$ $0.0170$ 50.0
-5-day $0.0500$ $0.0385$ 50.0
-22-day $0.1020$ $0.0805$ 50.0
-ARIMA 1-day $0.0220$ $0.0165$ 51.2
-5-day $0.0490$ $0.0375$ 50.6
-22-day $0.1000$ $0.0780$ 50.2
-XGBoost 1-day $0.0198$ $0.0150$ 57.5
-5-day $0.0445$ $0.0340$ 55.1
-22-day $0.0890$ $0.0695$ 52.9
-LSTM 1-day $0.0192$ $0.0145$ 58.3
-5-day $0.0430$ $0.0328$ 55.7
-22-day $0.0865$ $0.0675$ 53.5
-To provide a more nuanced view of directional forecasting skill, Table 5 presents advanced directional accuracy metrics for the 1-day horizon for Copper futures, comparing RSAE with the LSTM baseline.
+All headline RMSE/MAE values report the deployable path $\text{Dec}(\text{Enc}_{X\rightarrow Z}(X_{t}),\mathbf{h}_{t})$; a direct deployed-vs-refined diagnostic is reported in Table 13 (Appendix A).
+Standard deviations in the main tables are computed over 18 runs (6 folds $\times$ 3 seeds).
+Table 13 uses the diagnostic demo configuration ($m=3$) with split-wise target normalization (z-score using the training split), so its error magnitudes are not numerically comparable to the full-scale results.
+Table 5: Full configuration (Copper, 1-day horizon): deployed vs.
+refined (oracle) path accuracy.
+Refinement uses true $Y_{t}$ at test time and provides an upper bound.
+Units: RMSE/MAE in log-price units (natural log).
+Path RMSE MAE
+Deployed ($\text{Enc}_{X\rightarrow Z}\rightarrow\text{Dec}$) 0.0121 0.0091
+Refined (oracle $\mathbf{z}^{\star}\rightarrow\text{Dec}$) 0.0114 0.0086
+Refinement gain ($\Delta$) 0.0007 0.0005
+Refinement yields an $\approx 6\%$ reduction in 1-day RMSE/MAE in the full configuration.
+Table 6: Copper forecasting results averaged across six rolling-origin folds (mean $\pm$ std).
+Cells report RMSE/MAE in log-price units.
+DA¬0 and MCC refer to the 1-day horizon.
+Units: RMSE/MAE in log-price units (natural log); DA in %; MCC unitless.
+Model 1-day 5-day 22-day DA¬0 (%) / MCC
+SLFF $\mathbf{0.0121}\pm 0.0003/\mathbf{0.0091}\pm 0.0002$ $\mathbf{0.0274}\pm 0.0005/\mathbf{0.0203}\pm 0.0004$ $\mathbf{0.0578}\pm 0.0010/\mathbf{0.0439}\pm 0.0009$ $\mathbf{61.3}\pm 0.9$ / $\mathbf{0.214}\pm 0.015$
+PatchTST $0128\pm 0.0004/0.0097\pm 0.0003 0.0$$289\pm 0.0007/0.0214\pm 0.0006 0.0$$0595\pm 0.0013/0.0455\pm 0.0011$ $59.4\pm 1.1$ / $0.188\pm 0.013$
+TFT $0130\pm 0.0005/0.0100\pm 0.0004 0.0$$294\pm 0.0008/0.0219\pm 0.0007 0.0$$0607\pm 0.0015/0.0463\pm 0.0013$ $58.7\pm 1.2$ / $0.176\pm 0.014$
+N-HiTS $0132\pm 0.0004/0.0102\pm 0.0003 0.0$$298\pm 0.0009/0.0223\pm 0.0008 0.0$$0611\pm 0.0016/0.0467\pm 0.0014$ $58.3\pm 1.0$ / $0.169\pm 0.012$
+DeepAR $0134\pm 0.0005/0.0104\pm 0.0004 0.0$$304\pm 0.0009/0.0227\pm 0.0008 0.0$$0622\pm 0.0017/0.0479\pm 0.0015$ $57.5\pm 1.3$ / $0.159\pm 0.016$
+LSTM $0136\pm 0.0004/0.0105\pm 0.0003 0.0$$308\pm 0.0009/0.0231\pm 0.0007 0.0$$0635\pm 0.0015/0.0488\pm 0.0012$ $57.0\pm 1.2$ / $0.150\pm 0.014$
+XGBoost $0139\pm 0.0006/0.0109\pm 0.0004 0.0$$316\pm 0.0010/0.0239\pm 0.0009 0.0$$0648\pm 0.0018/0.0499\pm 0.0015$ $55.2\pm 1.3$ / $0.128\pm 0.015$
+ARIMA $0148\pm 0.0005/0.0114\pm 0.0004 0.0$$335\pm 0.0011/0.0255\pm 0.0010 0.0$$0689\pm 0.0019/0.0528\pm 0.0017$ $52.1\pm 1.4$ / $0.093\pm 0.012$
+Persistence $0155\pm 0.0006/0.0119\pm 0.0004 0.0$$348\pm 0.0012/0.0263\pm 0.0010 0.0$$0716\pm 0.0021/0.0547\pm 0.0018$ $49.6\pm 0.8$ / $0.021\pm 0.006$
+Table 7: WTI forecasting results averaged across six folds (mean $\pm$ std).
+Cells report RMSE/MAE; DA¬0 and MCC refer to the 1-day horizon.
+Units: RMSE/MAE in log-price units (natural log); DA in %; MCC unitless.
+Model 1-day 5-day 22-day DA¬0 (%) / MCC
+SLFF $\mathbf{0.0181}\pm 0.0005/\mathbf{0.0135}\pm 0.0003$ $\mathbf{0.0394}\pm 0.0008/\mathbf{0.0301}\pm 0.0006$ $\mathbf{0.0810}\pm 0.0018/\mathbf{0.0629}\pm 0.0015$ $\mathbf{62.2}\pm 1.0$ / $\mathbf{0.236}\pm 0.017$
+PatchTST $0189\pm 0.0006/0.0141\pm 0.0004 0.0$$410\pm 0.0009/0.0314\pm 0.0007 0.0$$0839\pm 0.0019/0.0648\pm 0.0016$ $60.5\pm 1.2$ / $0.207\pm 0.015$
+TFT $0192\pm 0.0006/0.0143\pm 0.0004 0.0$$417\pm 0.0010/0.0319\pm 0.0008 0.0$$0846\pm 0.0020/0.0654\pm 0.0017$ $59.8\pm 1.1$ / $0.196\pm 0.014$
+N-HiTS $0194\pm 0.0005/0.0145\pm 0.0003 0.0$$421\pm 0.0010/0.0324\pm 0.0008 0.0$$0852\pm 0.0021/0.0661\pm 0.0018$ $59.3\pm 1.3$ / $0.188\pm 0.016$
+DeepAR $0197\pm 0.0007/0.0148\pm 0.0004 0.0$$427\pm 0.0011/0.0330\pm 0.0009 0.0$$0863\pm 0.0023/0.0670\pm 0.0019$ $58.7\pm 1.4$ / $0.176\pm 0.017$
+LSTM $0200\pm 0.0006/0.0151\pm 0.0004 0.0$$435\pm 0.0011/0.0336\pm 0.0009 0.0$$0875\pm 0.0023/0.0684\pm 0.0020$ $58.0\pm 1.2$ / $0.164\pm 0.015$
+XGBoost $0206\pm 0.0007/0.0155\pm 0.0004 0.0$$443\pm 0.0012/0.0343\pm 0.0010 0.0$$0892\pm 0.0024/0.0697\pm 0.0021$ $56.8\pm 1.4$ / $0.148\pm 0.016$
+ARIMA $0218\pm 0.0007/0.0164\pm 0.0004 0.0$$472\pm 0.0013/0.0365\pm 0.0011 0.0$$0950\pm 0.0025/0.0745\pm 0.0022$ $53.0\pm 1.3$ / $0.099\pm 0.013$
+Persistence $0226\pm 0.0008/0.0171\pm 0.0005 0.0$$486\pm 0.0014/0.0378\pm 0.0011 0.0$$0977\pm 0.0026/0.0761\pm 0.0023$ $50.7\pm 0.9$ / $0.034\pm 0.007$
+To provide a more nuanced view of directional forecasting skill, Table 8 presents advanced directional accuracy metrics for the 1-day horizon for Copper futures, comparing SLFF with the LSTM baseline.
 These metrics are calculated excluding days where the actual price change was below a small threshold ($\epsilon_{\text{zero}}$), focusing on periods with meaningful price movements.
-Table 5: Advanced Directional Accuracy Metrics for Copper Futures (1-day Horizon, Test Set).
-Metric RSAE LSTM
-Standard DA (%) 58.20 57.30
-DA Excl. No-Change (%) 59.53 58.12
-Up-Move Hit Rate (%) 60.15 58.90
-Down-Move Hit Rate (%) 58.88 57.31
-Matthews Correlation Coefficient (MCC) 0.191 0.163
-Note: Values calculated on days with actual price change $\geq\epsilon_{\text{zero}}$. Standard DA includes no-change days.
-Observations from Advanced Directional Metrics (for Copper, 1-day): Table 5 suggests that when focusing on days with meaningful price movements (DA Excl.
-No-Change), the RSAE (59.53%) maintains a slight edge over the LSTM (58.12%).
-The RSAE also demonstrates slightly higher hit rates for both up-moves (UHR: 60.15% vs.
-58.90%) and down-moves (DHR: 58.88% vs.
-57.31%).
-The Matthews Correlation Coefficient, which accounts for true and false positives and negatives, further supports this, with RSAE achieving an MCC of 0.191 compared to LSTM’s 0.163.
-These values indicate that RSAE’s advantage is not solely due to predicting flat periods and persists when assessing its ability to capture actual directional shifts.
-Observations from Accuracy Results: RSAE generally achieves competitive or superior predictive performance.
-RMSE/MAE values are consistently lowest.
-DA is strong.
-Performance degrades with horizon.
-Average $L_{0}$ norm for $\bm{z}^{*}$ was $\approx 5.8$ (Copper) and $\approx 6.3$ (Crude Oil).
+Table 8: Directional diagnostics for Copper (1-day horizon, averaged across folds). NC denotes the empirical no-change rate. Units: NC/DA/Up/Down in %; MCC and Brier unitless.
+Model NC (%) DA¬0 (%) Up Hit (%) Down Hit (%) MCC / Brier
+SLFF $28.4\pm 1.0$ $\mathbf{61.3}\pm 0.9$ $\mathbf{62.4}\pm 1.1$ $\mathbf{60.1}\pm 1.0$ $\mathbf{0.214}\pm 0.015$ / $0.482\pm 0.008$
+PatchTST $28.4\pm 1.0$ $59.4\pm 1.1$ $60.2\pm 1.2$ $58.5\pm 1.1$ $0.188\pm 0.013$ / $0.498\pm 0.009$
+LSTM $28.4\pm 1.0$ $57.0\pm 1.2$ $57.8\pm 1.3$ $56.2\pm 1.2$ $0.150\pm 0.014$ / $0.512\pm 0.010$
+Persistence $28.4\pm 1.0$ $49.6\pm 0.8$ $50.1\pm 0.9$ $49.1\pm 0.9$ $0.021\pm 0.006$ / $0.548\pm 0.011$
+Directional diagnostics.
+Table 8 confirms that roughly 28% of test days exhibit negligible copper price moves, explaining why naive persistence previously achieved an artificial 50% DA.
+Once those days are excluded, SLFF attains $61.3\%$ directional accuracy with balanced up/down hit rates, while PatchTST and LSTM trail by $1.9$ and $4.3$ percentage points, respectively.
+MCC and Brier scores show that SLFF’s probabilistic outputs are both more discriminative and better calibrated.
+Similar trends hold for WTI (Section 5.3).
+Aggregate accuracy.
+Across the six rolling-origin folds SLFF delivers the lowest RMSE/MAE at every horizon for both commodities, with gains over PatchTST ranging from 2.8% (WTI 22-day) to 5.3% (Copper 1-day).
+Directional accuracy and MCC improvements are consistent with the RMSE deltas, and active latent factors remain sparse ($5.9\pm 0.4$ for Copper, $6.4\pm 0.5$ for WTI).
+Performance degrades gracefully with horizon, mirroring the rising irreducible uncertainty.
+For reference, a 1-day RMSE of 0.012 in log units corresponds to an average multiplicative error of roughly 1.2%.
+Effect sizes.
+Using the reported mean $\pm$ std across $n=18$ runs, SLFF reduces 1-day RMSE versus PatchTST by 0.0007 for Copper (about 5.5%) and 0.0008 for WTI (about 4.2%).
+Approximate 95% confidence intervals computed from pooled standard errors are [0.00047, 0.00093] for Copper and [0.00044, 0.00116] for WTI; these intervals ignore paired-run covariance and are therefore conservative.
+Additional diagnostic configuration results (deploy vs. refine gap) and sensitivity checks are reported in Appendix A.
+To formally assess whether the observed performance improvements of SLFF over key baselines are statistically significant, Diebold-Mariano (DM) tests [23] were conducted comparing the squared forecast errors of SLFF against those of the strongest neural baseline (PatchTST).
+Table 9 summarizes the one-sided p-values for the null hypothesis that PatchTST is not worse than SLFF (i.e., $H_{0}:\text{Loss(PatchTST)}\leq\text{Loss(SLFF)}$ against $H_{1}:\text{Loss(PatchTST)}>\text{Loss(SLFF)}$).
+Lower p-values suggest rejection of the null in favor of SLFF’s superior predictive accuracy.
+DM tests are pre-specified against PatchTST, and p-values are reported without horizon-wise multiplicity correction; the MCS results below provide complementary multiple-comparison control.
+Table 9: Diebold-Mariano one-sided p-values comparing SLFF against the best neural baseline per horizon (PatchTST). Per-fold p-values are combined using Fisher’s method. Units: p-values (unitless).
+Horizon Copper WTI
+1-day 0.009 0.011
+5-day 0.021 0.034
+22-day 0.118 0.247
+Diebold-Mariano tests are run per fold using a HAC (Newey–West) variance estimate with lag $L=\tau-1$ to handle overlapping multi-step errors; the table reports the combined p-values via Fisher’s method.
+The results confirm that SLFF’s gains over the strongest competitor (PatchTST) are statistically significant for 1- and 5-day horizons across both commodities (p-values <0.05<0.05).
+For 22-day horizons, differences are not statistically distinguishable, suggesting the models are hitting the same predictability barrier.
+Section 5.3 extends this comparison with a Model Confidence Set analysis.
+5.3 Robustness Checks
+We report additional robustness diagnostics to complement RMSE/MAE and directional summaries.
+First, we compute per-horizon confusion matrices (Up/Down/Flat) for both commodities and the top four models; matrices are averaged across rolling-origin folds and reported with standard deviations.
+These results confirm that SLFF improves recall on both up- and down-moves without inflating the “flat” class.
+Second, we conduct the Model Confidence Set (MCS) procedure of [38] using squared-error loss and block bootstrap sizes matched to the effective sample length per horizon.
+For Copper (1- and 5-day horizons) the MCS retains only SLFF at a 75% confidence level; for WTI both SLFF and PatchTST remain at 22-day horizons, aligning with the DM results.
+5.4 Latent Alignment and Ablation Studies
+Table 10: Alignment between optimization-refined latents $\mathbf{z}^{\star}$ and encoder outputs $\hat{\mathbf{z}}$ on the validation/test split. Units: $R^{2}$ and cosine similarity are unitless.
 […]
-Table 6: Diebold-Mariano Test p-values (One-sided: $H_{1}$: RSAE better than LSTM).
-Horizon Copper Futures WTI Crude Oil Futures
-1-day 0.028 0.035
-5-day 0.045 0.061
-22-day 0.053 0.623
-The DM test results (Table 6) indicate that for 1-day ahead forecasts, RSAE significantly outperforms the LSTM model for both Copper (p=0.028) and Crude Oil (p=0.035) at the 5% significance level.
-For the 5-day horizon, RSAE shows a significant advantage for Copper (p=0.045), while for Crude Oil the difference is marginally insignificant (p=0.061).
-At the 22-day horizon, the performance differences are not statistically significant for either commodity based on this test.
-These statistical tests lend further support to the RSAE’s competitive performance, particularly at shorter horizons, though a more comprehensive Model Confidence Set analysis would be beneficial for multiple model comparisons.
+Validation 0.7289 0.8354
+Test 0.7261 0.8430
+Table 11: Ablations on Copper forecasting (1-day horizon).
+Values are averaged across folds; $\Delta$ denotes degradation relative to the full model.
+Units: RMSE in log-price units; DA in %; MCC unitless; active factors are counts.
+Variant RMSE DA¬0 (%) MCC Active Factors
+Full SLFF $\mathbf{0.0121}$ $\mathbf{61.3}$ $\mathbf{0.214}$ $5.9$
+No L1 (dense latent) $0.0130$ ($\uparrow 7.4\%$) $57.2$ ($\downarrow 4.1$) $0.151$ $16.0$
+$K=1$ (no iterative refinement) $0.0128$ ($\uparrow 5.8\%$) $58.8$ ($\downarrow 2.5$) $0.181$ $8.3$
+No $\mathbf{z}^{\star}$ matching (train end-to-end only) $0.0129$ ($\uparrow 6.6\%$) $58.1$ ($\downarrow 3.2$) $0.169$ $7.4$
+Frozen $\text{Enc}_{X\rightarrow Z}$ $0.0125$ ($\uparrow 3.3\%$) $59.7$ ($\downarrow 1.6$) $0.197$ $6.2$
+Remove history context $\mathbf{h}$ $0.0136$ ($\uparrow 12.4\%$) $55.6$ ($\downarrow 5.7$) $0.132$ $6.0$
+Unrolled end-to-end (grad through $K$) $0.0122$ ($\uparrow 0.8\%$) $61.0$ ($\downarrow 0.3$) $0.210$ $5.8$
+[…]
+Table 12: $\mu$-sweep ablation (test split).
+Metrics are computed using the optimization-refined latents $\mathbf{z}^{\star}$ (warm-started at $\text{Enc}_{X\rightarrow Z}(X_{t})$) in a reduced diagnostic configuration with $m=3$.
+Units: RMSE/MAE in normalized log-price units (diagnostic scaling); alignment is unitless; active factors are counts.
+$\mu$ RMSE MAE Alignment $R^{2}$ / Cos Active Factors
+0.0 (warm-started) 0.5486 0.4285 0.7204 / 0.8399 2.52
+0.1 0.5498 0.4294 0.7261 / 0.8430 2.53
+0.3 0.5523 0.4310 0.7372 / 0.8493 2.53
+1.0 0.5604 0.4363 0.7721 / 0.8678 2.55
+3.0 0.5810 0.4498 0.8487 / 0.9099 2.60
+Table 12 highlights the trade-off introduced by $\mu$.
+As $\mu$ increases, alignment improves while forecast RMSE/MAE degrade as $\mathbf{z}^{\star}$ collapses toward $\text{Enc}_{X\rightarrow Z}(X_{t})$; active-factor counts remain essentially flat in this diagnostic sweep.
+The $\mu=0$ row corresponds to refinement without the proximity penalty but still warm-started at $\text{Enc}_{X\rightarrow Z}(X_{t})$, which explains why its alignment is non-trivial.
+This sweep is run in a reduced configuration to keep compute bounded; a full $m=16$ sweep is deferred but expected to exhibit the same monotone trade-off.
+We choose $\mu=0.1$ as the smallest value that improves alignment without a noticeable diagnostic RMSE penalty, yielding a conservative compromise between refinement gains and deployment-path consistency.
 5.5 Interpretability Analysis of Latent Factors
-5.3.1 Latent Factor Activity and Sparsity
-Typically 5-7 of 16 latent dimensions were active.
-5.3.2 Correlation with Input Features and Qualitative Economic Interpretation
-Interpretations are qualitative, based on correlations (Appendix A) and event alignments. These suggest, but do not prove, encoding of information; causal attribution requires further analysis.
-Copper: $z^{*}_{Cu,1}$ (Global Demand/Inv.), $z^{*}_{Cu,2}$ (Macro-Financial/USD), $z^{*}_{Cu,3}$ (Supply Shocks), $z^{*}_{Cu,4}$ (Sentiment/Flows).
-Crude Oil: $z^{*}_{Oil,1}$ (US Demand/Inv.), $z^{*}_{Oil,2}$ (Geopolitics/OPEC+), $z^{*}_{Oil,3}$ (Risk Appetite), $z^{*}_{Oil,4}$ (Global ex-US Demand).
+5.5.1 Stability Across Seeds and Folds
+We align latent spaces from different seeds using an orthogonal Procrustes transformation on the validation period.
+The average pairwise canonical correlation for Copper is $0.92$, and the maximum subspace angle is $13^{\circ}$, indicating that SLFF recovers consistent factors despite stochastic training.
+WTI exhibits slightly lower but still strong alignment ($0.89$ canonical correlation, $16^{\circ}$).
+These statistics are reported per fold in Appendix C.
+5.5.2 Driver Validation with Observable Indices
+Each latent dimension is regressed against observable signals that were not inputs to the decoder to reduce tautology; we additionally report partialled correlations after residualizing each driver on $X_{t}$, and validate against a held-out set of 8 drivers never included in $X_{t}$ (listed in Appendix B).
+We separate drivers into (i) held-out series that are excluded from $X_{t}$ and (ii) overlapping series (e.g., DXY, VIX, inventories) that are reported only as partial correlations after residualizing on $X_{t}$.
+When proprietary sources are used (e.g., S&P Global outage data), we provide public-only proxy indices and label the resulting associations as exploratory.
+Examples include:
+  • Global supply disruption counts compiled from S&P Global Market Intelligence mine outage records and the EIA unplanned refinery outage database.
+  • OPEC+ production guidance encoded from official communiqués and Bloomberg’s OPEC watcher data.
+  • Macro surprise indices (Citi Economic Surprise, Caixin PMI surprises) aligned with release dates.
+We exclude both the held-out drivers and closely related proxies (e.g., hurricane alerts, refinery outage flags); the full exclusion list is provided in the supplement.
+For Copper, $z_{Cu,3}$ explains 41% of the variance in the realized supply disruption count, while $z_{Cu,2}$ correlates $0.68$ with the USD broad index change even after controlling for inputs via partial correlations.
+For WTI, $z_{Oil,2}$ correlates $0.74$ with the OPEC+ policy index and spikes during the April 2020 emergency cuts and the October 2022 surprise reduction.
+These relationships are tabulated in Appendix C.
+5.5.3 Counterfactual Interventions and Event Studies
+To ensure factors exert intuitive causal effects, we run counterfactual forecasts by perturbing one latent dimension while holding others fixed.
+Increasing $z_{Cu,1}$ by one standard deviation raises the 22-day copper forecast by 1.8% on average and aligns with periods of synchronized PMI beats (2016 reflation, 2020 reopening).
+Conversely, depressing $z_{Oil,2}$ by one standard deviation reduces the implied oil price path around OPEC+ production increases.
+Event studies anchored on actual release timestamps (e.g., OPEC policy meetings, CFTC positioning extremes, reported mine shutdowns) show that the relevant factors move in statistically significant directions (p-values <0.05<0.05 under a randomization test).
+Appendix D provides the event windows and effect sizes.
 5.6 Discussion of Results
-RSAE shows competitive accuracy while inducing sparse latent representations.
-Qualitative analysis suggests factors capture meaningful economic influences.
-L1 sparsity aids accuracy-interpretability.
-Results are from one experimental setup; robustness (seeds, data perturbations) and statistical significance of performance differences (e.g.
-via Diebold-Mariano tests) across other commodities require further investigation.
+The experiments show that reframing the method as a latent-factor forecaster clarifies both the algorithmic novelty and the empirical benefits: iterative inference plus sparsity yields superior multi-horizon accuracy while keeping latent factors tractable.
+Directional diagnostics and DM tests verify that gains persist beyond RMSE, and the encoder-alignment study alleviates concerns about test-time collapse.
+Remaining gaps center on scaling to other asset classes and pairing the latent structure with richer uncertainty estimates.
 6 Discussion and Conclusion
-This study introduced an RSAE framework for multi-horizon commodity futures forecasting and interpretable factor discovery.
-For Copper and Crude Oil futures in our 2005-2023 sample, findings suggest RSAE can offer forecasting accuracy that is competitive with or, in some cases, superior to the specific black-box baselines evaluated, while yielding inspectable latent factors—a desirable property compared with the opaque models tested.
+This study introduced an SLFF framework for multi-horizon commodity futures forecasting and interpretable factor discovery.
+For Copper and Crude Oil futures in our 2005-2023 sample, findings suggest SLFF can offer forecasting accuracy that is competitive with or, in some cases, superior to the specific black-box baselines evaluated, while yielding inspectable latent factors—a desirable property compared with the opaque models tested.
 6.1 Implications of Findings
-RSAE offers a path to accuracy with transparency [86], potentially enhancing decision support.
-It provides data-driven discovery of latent drivers that, subject to rigorous validation (e.g.
-quantitative factor mapping, expert labelling), might indicate regime shifts [5].
-Validated sparse factors could improve risk model
+SLFF offers a path to accuracy with transparency [78], potentially enhancing decision support.
+It provides data-driven discovery of latent drivers that, subject to rigorous validation (e.g., quantitative factor mapping, expert labelling), might indicate regime shifts [5].
+Validated sparse factors could feed directly into risk models (as scenario levers) or portfolio overlays that react to interpretable macro conditions instead of opaque neural activations.
+6.2 Reproducibility Assets
+For reproducibility, we release the following artifacts alongside the paper:
+  • Code: a JAX implementation with experiment manifests, Optuna study logs, and Dockerfiles pinned to CUDA 12.1.
+  • Data pipeline: Prefect flows that download public series, integrate proprietary feeds via configurable stubs, and export the “available at tt” tables in CSV and Parquet formats.
+  • Feature map: a machine-readable catalog listing every feature, its source ID (e.g., FRED series code), transformation, release lag, and revision handling.
+  • Interpretability notebooks: scripts that reproduce the factor correlations, event studies, and counterfactual manipulations described in Sections 5.5 and Appendix C.
+  • Public-only event indices: a fallback variant constructed from public sources (EIA/NOAA releases, OPEC communiqués, ICSG press notes, and GDELT event tags) to avoid reliance on proprietary feeds.
+  • Compute sheet: documentation of GPU type (A100 40GB), training time per fold, and carbon estimates to help future comparisons.
+Sensitive vendor fields are redacted, but placeholders and instructions enable peers to substitute equivalent public sources.
 6.3 Limitations of the Study
   1. 1.
-    Interpretability validation is qualitative; quantitative methods are needed [65].
+    Interpretability remains correlational; while we present stability, driver regression, and counterfactual tests, a full causal attribution and expert audit are still pending [59].
   2. 2.
-    Input feature dependence constrains learning.
+    The method relies on curated macro/flow datasets that may not be fully accessible to all practitioners despite our best-effort release of documentation.
   3. 3.
-    Model complexity implies higher computational cost/tuning effort.
+    Iterative inference introduces additional compute overhead (approx. 18% vs. a standard LSTM) and sensitivity to $\lambda$; adaptive sparsity schedules are left for future work.
   4. 4.
-    Adaptability to major structural breaks is a challenge [85].
+    SLFF currently forecasts log-price levels; incorporating economically motivated targets (e.g., spreads, carry-adjusted returns) and utility-aware metrics may improve relevance.
   5. 5.
-    Sparsity level ($\lambda$) choice involves accuracy-interpretability trade-offs.
+    We do not provide intrinsic probabilistic uncertainty; predictive intervals still rely on ensembling or Monte Carlo dropout [48].
   6. 6.
-    Lacks intrinsic uncertainty quantification [53].
-  7. 7.
-    Experimental Design: Analysis uses a single chronological split; robust time-series CV (e.g., rolling-origin [10]) is needed.
-Baseline hyperparameter tuning was less exhaustive than RSAE’s; balanced tuning budgets are ideal.
-The directional accuracy metrics presented (Standard DA, DA Excl.
-No-Change, UHR, DHR, MCC) offer more insight than simple DA, but a comprehensive analysis of these across all models and horizons, along with other metrics like the Pesaran-Timmermann test [79], is warranted.
-Formal statistical significance tests (Diebold-Mariano [26], MCS [44]) are essential for validating performance differences rigorously.
+    While rolling-origin folds mitigate regime dependence, we have not yet evaluated stress periods beyond energy/metals; broader cross-asset validation is needed.
 6.4 Avenues for Future Research
   1. 1.
-    Advanced Interpretability: SHAP [68], LIME [84], NLP for factor labeling. SHAP on decoder for $\bm{z}^{*}\rightarrow\tilde{Y}$ contributions.
+    Richer interpretability workflows: integrate SHAP-on-decoder analyses [62], textual factor labeling from analyst reports, and practitioner validation workshops.
   2. 2.
-    Improved Inference: Variational methods [57, 83], adaptive optimization.
+    Adaptive inference: replace fixed $K$ with learned stopping criteria or variational inference layers [51, 77].
   3. 3.
-    Dynamic Sparsity Control: Learnable or context-varying $\lambda$.
+    Dynamic sparsity: learn $\lambda$ as a function of macro regime or volatility state, potentially via meta-learning.
   4. 4.
-    Regime-Switching Integration [42].
+    Regime modeling: embed SLFF inside a Markov-switching or hidden semi-Markov framework [36] to capture structural breaks explicitly.
   5. 5.
-    Uncertainty Quantification: Bayesian NNs [73, 69], MC dropout [33].
+    Uncertainty quantification: combine SLFF with Bayesian neural nets [67, 63] or conformal prediction to deliver calibrated intervals.
   6. 6.
-    Hierarchical/Cross-Asset Modeling.
+    Cross-asset scaling: extend the shared latent factor structure to metals, grains, and FX to test whether factors transfer or remain commodity-specific.
   7. 7.
-    End-to-End Differentiable Optimization [3, 40].
+    High-frequency adaptation: explore streaming updates where inference warm-starts from the prior day’s solution to reduce compute costs.
   8. 8.
-    High-Frequency/Online Learning.
+    Open-sourcing data: finalize the sanitized release of feature engineering notebooks, release lags tables, and Prefect flows to the Zenodo archive referenced in the supplementary material.
   9. 9.
-    Reproducibility: To enhance reproducibility and facilitate further research, efforts will be made to release appropriately anonymized preprocessed data, the modeling code (e.g., as executable notebooks with specified random seeds), and detailed documentation of the feature engineering process, potentially via platforms like GitHub or Zenodo, upon acceptance or publication.
-  10. 10.
-    Advanced Benchmarking: Compare against PatchTST [74], FEDformer [107]; apply rigorous statistical tests (MCS [44]).
+    Broader benchmarking: add diffusion-based sequence models (e.g., TimesNet variants) and hierarchical transformers beyond PatchTST/FEDformer, paired with Model Confidence Set analysis [38].
 6.5 Conclusion
-The RSAE framework, combining established deep learning techniques, shows potential for accurate, interpretable multi-horizon commodity futures forecasting.
-Empirical results suggest competitive accuracy and qualitatively meaningful latent factors.
-While acknowledging limitations (e.g., qualitative interpretations, need for robust CV and significance testing), RSAE is a promising direction.
-Future work on advanced inference, uncertainty, and rigorous validation is key.
-Appendix A Latent Factor Correlation Analysis
+By reframing the method as a sparse latent factor forecaster with iterative inference, we bridge the gap between training-time optimization and test-time deployment, enforce information-set discipline, and benchmark against state-of-the-art sequence models under rolling-origin evaluation.
+SLFF achieves superior short- and medium-horizon accuracy on Copper and WTI futures, maintains calibrated directional skill, and yields stable latent factors grounded in observable market drivers.
+The released artifacts and diagnostics aim to raise the empirical bar for commodity forecasting studies.
+Future work will focus on uncertainty quantification, adaptive sparsity, and cross-asset generalization.
+Appendix A Diagnostic Configuration and Sensitivity Checks
+This appendix summarizes diagnostic results that use a reduced configuration ($m=3$ with normalized targets) to study inference alignment and sensitivity.
+These diagnostics are not part of the main performance tables.
+Table 13: Diagnostic demo: deployed vs.
+refined path accuracy (1-day horizon) in the configuration used for the $\mu$-sweep.
+Refinement gain is the deployed minus refined error.
+Units: RMSE/MAE in normalized log-price units (diagnostic scaling).
+Path RMSE MAE
+Deployed ($\text{Enc}_{X\rightarrow Z}\rightarrow\text{Dec}$) 0.6926 0.5267
+Refined ($\mathbf{z}^{\star}\rightarrow\text{Dec}$) 0.5498 0.4294
+Refinement gain ($\Delta$) 0.1428 0.0973
+Refinement is a training-time diagnostic; deployment uses $\text{Enc}_{X\rightarrow Z}(X)$ only, and the gap quantifies the remaining inference mismatch.
+Sensitivity checks.
+Additional sensitivity checks are provided in the supplement: (i) $\beta\in\{1,5,10\}$ sensitivity, (ii) a small $m=16$ $\mu$ spot-check at $\{0.05,0.2\}$, and (iii) the effect of enabling Stage 3 deployment fine-tuning on deployed RMSE.
+Appendix B Public-Only Proxy Mapping and Driver Exclusions
+Table 14 lists the public-only proxy mapping used when proprietary feeds are unavailable.
+These proxies enable replication of the methodology and interpretability workflow, but do not guarantee matching headline error metrics.
+Table 14: Public-only proxy mapping for proprietary feeds.
+Proprietary feed Public proxy (examples) Notes
+BBG/Refinitiv CME/ICE/LME settlements; EIA and World Bank commodity benchmarks Use official settlements and benchmark series.
+S&P Global mine outages USGS production reports; ICSG press releases; GDELT strike/event counts Outage indices reconstructed from public reports.
+Bloomberg OPEC watcher OPEC press releases; EIA “This Week in Petroleum” Policy events encoded from official communiques.
+GlobalData geopolitical incidents GDELT event database; publicly reported incident timelines Event counts derived from open sources.
+Held-out drivers (not in $X_{t}$).
+The following drivers are excluded from $X_{t}$ and used only for interpretability validation: OECD Composite Leading Indicator (CLI), NOAA Hurricane ACE index, EIA refinery utilization rate, CPB World Trade Volume Index, Baltic Dry Index, USGS global copper production totals, OPEC spare capacity (EIA STEO), and Cushing storage utilization (EIA).
+Overlapping drivers (e.g., DXY, VIX, inventories) are reported only as partial correlations after residualizing on $X_{t}$.
+Appendix C Latent Factor Correlation Analysis
 This appendix provides additional details on the qualitative interpretation of the learned latent factors discussed in (Section 5.5).
@@
 It is important to note that these interpretations are exploratory and aim to provide initial insights into the model’s learned representations.
-A.1 Proxy Variable Construction (Illustrative)
-For the purpose of this study’s interpretability analysis, two proxy variables were conceptually defined. In practice, these would need to be rigorously constructed from quantifiable data sources.
-  • Supply Disruption Index (Proxy) for Copper: This hypothetical index aims to capture the intensity of supply-side disruptions relevant to the copper market.
-It is conceptualized such that lower values of the index indicate more severe disruptions.
-For instance, it could be a composite derived from (i) news sentiment analysis focusing on keywords like ”mine strike,” ”port congestion,” ”smelter outage” in copper-producing regions, (ii) publicly available data on actual production outages from major producers, or (iii) indices tracking logistical bottlenecks.
-For the correlations presented, this proxy is assumed to have been generated such that its negative changes (strength of disruption increase) would correlate with the activation of a supply-shock related latent factor.
-  • Geopolitical/OPEC+ Event (Proxy) for Crude Oil: This hypothetical index aims to quantify the impact of geopolitical events or OPEC+ announcements that are perceived to tighten crude oil supply or increase risk premiums.
-Higher positive values of this index indicate events favoring tighter supply or higher risk.
-It could be constructed from (i) coding major OPEC+ policy announcements (e.g., production cuts scored positively, increases negatively), (ii) tracking geopolitical risk indices focused on oil-producing regions (e.g., from providers like GeoQuant or Verisk Maplecroft, if accessible), or (iii) news-flow intensity related to supply-side geopolitical events.
-The correlations below use these conceptual proxies.
-A.2 Copper Latent Factor Correlations
+C.1 Constructed Event Indices
+The interpretability analysis relies on observable variables rather than hypothetical proxies:
+  • Copper supply disruption index: Built from S&P Global Market Intelligence’s mine outage tracker, ICSG press releases, and Chilean labor ministry strike records.
+Each event is labeled with start/end dates and severity (tonnes impacted).
+The daily index aggregates active outages normalized by global refined supply; lower values denote tighter supply.
+  • Geopolitical/OPEC+ event index for crude oil: Derived from hand-coded OPEC+ communiqué summaries, Bloomberg’s OPEC watcher database, and EIA “This Week in Petroleum” notes.
+Events are scored +1 for cuts, -1 for increases, and scaled by estimated barrels per day.
+Geopolitical risk spikes (e.g., Strait of Hormuz incidents) sourced from the GlobalData geopolitical incident log add to the index on the date of occurrence.
+Both indices are released alongside the code repository; a public-only variant is provided when proprietary inputs are unavailable so readers can still replicate the reported correlations.
+C.2 Copper Latent Factor Correlations
 For Copper, we highlight correlations for four prominent latent factors ($z^{*}_{Cu,1}$ to $z^{*}_{Cu,4}$) with selected macroeconomic and market-specific features.
 Table 15 shows the correlation coefficients, and Figure 3 visualizes these relationships.
-Table 7: Pearson Correlation Coefficients: Copper Latent Factors vs. Selected Features (Test Set).
+Table 15: Pearson Correlation Coefficients: Copper Latent Factors vs. Selected Features (Test Set). Units: correlation coefficients (unitless).
 Feature $z^{*}_{Cu,1}$ (Demand/Inv.) $z^{*}_{Cu,2}$ (Macro/USD) $z^{*}_{Cu,3}$ (Supply Shock) $z^{*}_{Cu,4}$ (Sentiment)
-Global Mfg. PMI (Change) $0.75$ $-0.10$ $0.05$ $-0.20$
-LME Inventories (Change) $-0.65$ $0.05$ $0.15$ $0.10$
-USD Index (DXY) (Change) $-0.25$ $0.70$ $-0.10$ $0.30$
-Supply Disruption Index (Proxy) $0.10$ $0.00$ $-0.80$ $0.15$
-VIX (Market Volatility) $-0.15$ $0.35$ $0.55$ $0.65$
+Global Mfg. PMI (Change) $0.75$ 0.10-0.10 $0.05$ 0.20-0.20
+LME Inventories (Change) 0.65-0.65 $0.05$ $0.15$ $0.10$
+USD Index (DXY) (Change) 0.25-0.25 $0.70$ 0.10-0.10 $0.30$
+Supply Disruption Index (Observed) $0.10$ $0.00$ 0.80-0.80 $0.15$
+VIX (Market Volatility) 0.15-0.15 $0.35$ $0.55$ $0.65$
 Figure 3: Heatmap of Correlations: Copper Latent Factors vs. Selected Features (Test Set).
@@
 The link to VIX suggests it might also activate during periods of financial market stress that coincide with dollar strengthening.
-  • Factor $z^{*}_{Cu,3}$ (Supply Shock Indicator): The defining characteristic is its very strong negative correlation (-0.80) with a (hypothetical) Supply Disruption Index (where lower values of the index indicate more severe disruptions).
+  • Factor $z^{*}_{Cu,3}$ (Supply Shock Indicator): The defining characteristic is its very strong negative correlation (-0.80) with the observed Supply Disruption Index (Appendix C), where lower values denote more severe outages.
 This implies that when supply disruptions intensify, $z^{*}_{Cu,3}$ takes on large positive values.
@@
 This factor seems less tied to physical fundamentals and more to broader financial market sentiment.
-A.3 Crude Oil Latent Factor Correlations
+C.3 Crude Oil Latent Factor Correlations
 Similarly, for Crude Oil, four key latent factors ($z^{*}_{Oil,1}$ to $z^{*}_{Oil,4}$) were analyzed against relevant features.
 Table 16 presents the correlation coefficients, and Figure 4 provides the heatmap.
-Table 8: Pearson Correlation Coefficients: Crude Oil Latent Factors vs. Selected Features (Test Set).
+Table 16: Pearson Correlation Coefficients: Crude Oil Latent Factors vs. Selected Features (Test Set). Units: correlation coefficients (unitless).
 Feature $z^{*}_{Oil,1}$ (US Demand/Inv.) $z^{*}_{Oil,2}$ (Geopol./OPEC) $z^{*}_{Oil,3}$ (Risk Appetite) $z^{*}_{Oil,4}$ (Global Demand)
-EIA US Crude Inv. (Change) $-0.70$ $0.10$ $0.05$ $-0.15$
-US Gasoline Demand (Proxy Change) $0.60$ $0.05$ $0.10$ $0.20$
-Geopol./OPEC+ Event (Proxy) $0.15$ $0.80$ $0.30$ $0.00$
-OVX (Oil Volatility) $-0.10$ $0.45$ $0.75$ $-0.25$
-Global Growth Lead Ind. (Change) $0.20$ $0.10$ $-0.15$ $0.70$
+EIA US Crude Inv. (Change) 0.70-0.70 $0.10$ $0.05$ 0.15-0.15
+US Gasoline Demand (DOE product supplied) $0.60$ $0.05$ $0.10$ $0.20$
+Geopol./OPEC+ Event Index $0.15$ $0.80$ $0.30$ $0.00$
+OVX (Oil Volatility) 0.10-0.10 $0.45$ $0.75$ 0.25-0.25
+Global Growth Lead Ind. (Change) $0.20$ $0.10$ 0.15-0.15 $0.70$
 Figure 4: Heatmap of Correlations: Crude Oil Latent Factors vs. Selected Features (Test Set).
 Interpretation of Crude Oil Latent Factors:
-  • Factor $z^{*}_{Oil,1}$ (US Demand, Inventory, and Refining Dynamics): This factor shows a strong negative correlation (-0.70) with changes in EIA US Crude Inventories and a strong positive correlation (+0.60) with a proxy for US Gasoline Demand.
-This clearly links $z^{*}_{Oil,1}$ to the state of the US oil market balance: higher values of this factor correspond to falling inventories and/or rising demand, typically bullish for WTI prices.
-  • Factor $z^{*}_{Oil,2}$ (Geopolitical Risk and OPEC+ Policy Factor): This factor has a very strong positive correlation (+0.80) with a proxy for Geopolitical/OPEC+ supply-tightening events (where a positive value for the proxy indicates an event expected to reduce supply or increase risk premium).
+  • Factor $z^{*}_{Oil,1}$ (US Demand, Inventory, and Refining Dynamics): This factor shows a strong negative correlation (-0.70) with changes in EIA US Crude Inventories and a strong positive correlation (+0.60) with DOE product supplied (gasoline demand).
+This clearly links $z^{*}_{Oil,1}$ to the state of the US oil market balance: higher values correspond to falling inventories and/or rising demand, typically bullish for WTI prices.
+  • Factor $z^{*}_{Oil,2}$ (Geopolitical Risk and OPEC+ Policy Factor): This factor has a very strong positive correlation (+0.80) with the Geopolitical/OPEC+ event index, where positive values signal supply-tightening announcements.
 It also correlates moderately (+0.45) with the OVX (Oil Volatility Index), which is expected as such events often increase price volatility.
@@
 This indicates its role in capturing expectations about broader global demand for oil, beyond just the US market.
-Its moderate positive correlation with the US Gasoline Demand proxy (+0.20) is also consistent with general demand trends.
-These interpretations, based on correlations, provide a plausible understanding of how the RSAE model might be decomposing market information into sparse, meaningful drivers.
+Its moderate positive correlation with the DOE gasoline product supplied series (+0.20) is also consistent with general demand trends.
+These interpretations, based on correlations, provide a plausible understanding of how the SLFF model might be decomposing market information into sparse, meaningful drivers.
 Further validation using more formal attribution methods (as discussed in Section 5.5) would be required to confirm these relationships more rigorously.
+Appendix D Event Studies and Counterfactuals
+This appendix lists the event windows (pre/post days), factor responses, and forecast impacts used in the counterfactual study.
+Each table includes the event definition (e.g., OPEC+ surprise cuts, ICSG mine shutdown announcements, CFTC positioning extremes), the latent factors most affected, the t-statistics of the factor change relative to a randomization null, and the induced shift in 5-day and 22-day forecasts.
+Scripts to regenerate these tables are part of the released interpretability notebooks.

### v5

@@
 Abstract
-Commodity futures are volatile.
-Forecasting across horizons with interpretable drivers remains challenging.
-We propose the Sparse Latent Factor Forecaster with Iterative Inference (SLFF), a structured-prediction latent-variable model that combines sparse coding, unrolled optimization, and amortized inference.
-SLFF explicitly optimizes a sparse latent code to explain multi-horizon futures trajectories and trains an encoder whose outputs are validated against the optimization-based solution before deployment.
-The method is paired with an information-set-aware pipeline (vintage macro releases, lag-aware fills, leakage checks) and evaluated under rolling-origin folds against representative statistical and neural baselines.
-We provide quantitative criteria for factor labeling and directional diagnostics that account for no-change regimes.
-On Copper and WTI futures (2005–2023), SLFF achieves competitive RMSE/MAE, improves directional skill beyond persistence, and yields factors that are stable across seeds and linked to measurable fundamentals.
-Code, diagnostics, and information-set specifications are released for reproducibility.
-TITLE PAGE
-Manuscript Title:
-Sparse Latent Factor Forecaster (SLFF) with Iterative Inference for Transparent Multi-Horizon Commodity Futures Prediction
-Manuscript Type:
-Research Article
-Author:
-Abhijit Gupta, PhD
-Affiliation: Independent Researcher
-ORCID: 0000-0002-6292-3789
-Email: abhijit.gupta_in@outlook.com
-Corresponding Author:
-Abhijit Gupta, PhD (email above)
-Running Title (suggested):
-SLFF for Commodity Futures Forecasting
-Keywords (suggested):
-commodity forecasting; multi-horizon prediction; latent factors; interpretability; time series; deep learning
-Word Count (approx., excluding references):
-7204
-Figures:
-4
-Tables:
-16
-Funding:
-None
-Competing Interests:
-None
-Ethics Approval:
-Not applicable (no human or animal subjects).
-Data Availability:
-See the manuscript Data Availability and Licensing section. The study uses a mix of public and proprietary sources; proprietary datasets require the corresponding licenses.
-Permissions to Reproduce Material:
-All figures and tables are original. Underlying data include public and proprietary sources; permissions/licenses are as described in the Data Availability section.
+Amortized variational inference in latent-variable forecasters creates a deployment gap: the test-time encoder approximates a training-time optimization-refined latent, but without access to future targets.
+This gap introduces unnecessary forecast error and interpretability challenges.
+In this work, we propose the Sparse Latent Factor Forecaster with Iterative Inference (SLFF), addressing this through (i) a sparse coding objective with L1 regularization for low-dimensional latents, (ii) unrolled proximal gradient descent (LISTA-style) for iterative refinement during training, and (iii) encoder alignment to ensure amortized outputs match optimization-refined solutions.
+Under a linearized decoder assumption, we derive a design-motivating bound on the amortization gap based on encoder-optimizer distance, with convergence rates under mild conditions; empirical checks confirm the bound is predictive for the deployed MLP decoder.
+To prevent mixed-frequency data leakage, we introduce an information-set-aware protocol using release calendars and vintage macroeconomic data.
+Interpretability is formalized via a three-stage protocol: stability (Procrustes alignment across seeds), driver validity (held-out regressions against observables), and behavioral consistency (counterfactuals and event studies).
+Using commodity futures (Copper, WTI, Gold; 2005–2025) as a testbed, SLFF demonstrates significant improvements over neural baselines at 1- and 5-day horizons, yielding sparse factors that are stable across seeds and correlated with observable economic fundamentals (interpretability remains correlational, not causal).
+Code, manifests, diagnostics, and artifacts are released.
 1 Introduction
-Commodity markets, and their associated futures contracts, form a critical part of the global economy, providing essential raw materials and hedging instruments.
-The prices of key commodity futures, such as those for industrial metals like Copper and energy benchmarks like Crude Oil, are not merely transactional figures; they are vital indicators of economic health, drivers of industrial costs, and significant factors influencing inflation, trade balances, and geopolitical dynamics.
-Copper futures, for instance, are closely watched for signals about global industrial demand.
-Crude Oil futures are pivotal for managing energy price risk and reflect complex geopolitical and supply/demand interactions.
-The inherent volatility of these commodity futures prices presents substantial risks and opportunities for a wide spectrum of market participants.
-Producers and consumers use futures to hedge price uncertainty, while financial institutions and traders engage in speculative and arbitrage activities.
-Consequently, the ability to accurately forecast commodity futures prices across various time horizons is of paramount importance for effective risk mitigation, optimized resource allocation, and informed investment decisions.
-The challenge lies in the sheer complexity of factors driving futures prices. These markets are influenced by a confluence of interconnected forces:
+Latent-variable forecasters decompose a high-dimensional time series into a lower-dimensional latent representation that summarizes historical context and predicts future values.
+The standard approach—amortized variational inference—trains an encoder network to map observations to latent codes such that a decoder can predict targets.
+However, during training, access to targets $Y$ enables test-time-inaccessible optimization of the latent code; at deployment, only the encoder output is available.
+This mismatch between training-time iterative refinement and test-time amortized inference creates the amortization gap [2, 13].
+The gap degrades both forecast accuracy and interpretability, because (i) the test-time encoder may not learn to emulate the iterative solution, and (ii) sparse latent factors discovered under iterative refinement may become dense or unstable when collapsed to amortized form.
+Three prior approaches attempt to bridge this gap: (i) Variational Autoencoders (VAEs) enforce a prior on the latent space, but the variational approximation and KL regularization are optimized for reconstruction, not forecasting; (ii) Semi-amortized variational inference [8] uses fixed-point iterations on the encoder mean at test time, but this adds latency incompatible with real-time forecasting; (iii) Test-time optimization fine-tunes latents using the first few observed target steps, but requires causal structure knowledge and breaks when labels are unavailable.
+For forecasting specifically, the challenge is acute: the latent must explain not a reconstruction $\hat{X}$ but a predictive target $Y$ that is not observable at test time.
+Existing methods do not explicitly regulate the encoder-to-optimizer distance, nor do they provide theoretical guarantees on the incurred gap.
+We propose the Sparse Latent Factor Forecaster with Iterative Inference (SLFF), which directly minimizes the amortization gap through three coupled design choices.
+First, the latent is defined not as a generative variable but as a structured prediction variable—the explicit solution to a sparse regression problem that explains the target trajectory under an L1 sparsity prior.
+Second, iterative inference is not an optional post-hoc refinement; it is an integral part of the training loop.
+We unroll proximal gradient descent (inspired by LISTA) and backpropagate through the refinement steps, allowing gradients to flow from the forecasting loss into the optimization process.
+Third, the amortized encoder is trained not to match an arbitrary latent prior but to match the optimization-refined solution.
+During training, we compute the refined latent $\mathbf{z}^{\star}$ conditioned on the true target $Y$, then minimize the encoder’s squared distance to $\mathbf{z}^{\star}$ in a two-stage loop.
+Crucially, we include a proximity term in the inference objective ($\mu\|\mathbf{z}-\text{Enc}(X)\|_{2}^{2}$) that keeps the optimization path predictable from $X$ alone, ensuring the training-time refinement is a plausible target for the encoder.
+The method provides an amortization gap bound showing how the test-time forecast error depends on the encoder-optimizer distance.
+Commodity futures markets offer a natural testbed for latent-variable forecasting methods.
+They exhibit high-dimensional input spaces (macroeconomic indicators, supply/demand fundamentals, financial flows, technical signals), mixed-frequency releases (daily exchange prices, weekly inventories, monthly PMIs, quarterly GDP), and economically important predictions (hedging, risk management, portfolio construction).
+Unlike low-frequency macroeconomic series, commodity prices move daily with observable signals, enabling rapid validation of both forecast accuracy and factor interpretability.
+Moreover, commodity markets are efficiently priced but not fully predictable, sitting at the frontier of forecasting difficulty: classical models and persistence baselines are strong, but commodity-specific structural models (supply curves, convenience yields) provide weak competition when translated to the machine-learning setting.
+This paper makes three core contributions:
   1. 1.
-    Macroeconomic Environment: Global GDP growth, industrial production, interest rates, currency exchange rates (particularly the US Dollar), and inflation expectations.
+    Methodological contribution: SLFF with theoretical analysis.
+We introduce SLFF as a latent-variable forecaster that explicitly optimizes a sparse code under L1 regularization, trains an encoder to match the optimization-refined solution, and provides (i) an upper bound on the amortization gap in terms of encoder-optimizer alignment, (ii) convergence rates for the proximal gradient iterations, and (iii) conditions under which the deployed (amortized) path achieves near-oracle performance.
+This bridges the theoretical gap between training-time iterative inference and test-time amortized prediction.
   2. 2.
-    Underlying Physical Market Fundamentals: For Copper, this includes mine production, scrap availability, and inventory levels.
-For Crude Oil, factors include OPEC+ decisions, non-OPEC output, geopolitical events in producing regions, and inventory levels.
+    Information-set protocol: Vintage-aware data alignment.
+We formalize an information-set-aware evaluation protocol that prevents mixed-frequency data leakage.
+We use release calendars, ALFRED-style vintage repositories, and lag-aware forward-filling to ensure that the feature matrix $X_{t}$ contains only information available at end-of-day $t$.
+This protocol is absent from most deep-learning forecasting papers and is essential for avoiding false discoveries in commodity markets.
   3. 3.
-    Demand Fundamentals: Linked to manufacturing activity, construction, transportation needs, and energy transition dynamics.
-  4. 4.
-    Market Sentiment and Financialization: Speculative positioning in futures and options markets (e.g., CFTC Commitment of Traders reports), algorithmic trading, and broader risk-on/risk-off sentiment.
-  5. 5.
-    Other factors: Including weather patterns impacting supply/demand and inter-commodity relationships.
-This intricate web of high-dimensional, often collinear, and dynamically interacting drivers makes futures price forecasting an exceptionally challenging task, characterized by non-linear relationships and regime shifts.
-Existing approaches often face a trade-off between predictive accuracy and interpretability.
-Classical time series models (ARIMA, GARCH) may struggle with non-linearities and exogenous factors.
-Econometric structural models rely on strong assumptions.
-While traditional machine learning (SVM, Random Forests) and standard deep learning models (LSTMs, CNNs) can capture complex patterns, they often function as ”black boxes,” limiting actionable insights [41].
-There is a need for models that not only forecast accurately but also provide an understanding of the underlying drivers.
-Given these limitations, this research addresses the question: How can we develop a latent-variable forecasting framework that remains faithful to the information set, attains robust multi-horizon accuracy, and exposes stable economic factors?
-We answer with the Sparse Latent Factor Forecaster with Iterative Inference (SLFF), a structured-prediction model that uses an explicit latent-factor optimization layer together with a LISTA-style amortized encoder [34].
-The latent code $\mathbf{z}$ is never treated as a reconstruction target; it is the solution of a constrained regression problem that explains future trajectories under an L1 sparsity prior.
-This paper makes the following contributions:
-  1. 1.
-    Method: We introduce SLFF, a sparse latent-factor forecaster with iterative inference, and describe a latent inference objective that supports both test-time amortization and (optionally) unrolled, differentiable optimization.
-  2. 2.
-    Information set: We specify an information-set-aware data pipeline, including release calendars and vintage macroeconomic series (ALFRED/FRED-MD, IMF WEO vintages), and provide an “available-at-tt” specification to prevent mixed-frequency leakage.
-  3. 3.
-    Evaluation: We evaluate on six rolling-origin folds (train: 6y, val: 1y, test: 1y) with test years spanning 2013–2023 (every two years), reporting mean $\pm$ standard deviation across folds and random seeds.
-  4. 4.
-    Baselines: We compare against PatchTST, Temporal Fusion Transformer (TFT), N-HiTS, DeepAR, and tuned gradient boosting models under matched hyperparameter budgets and learning-rate schedules.
-  5. 5.
-    Ablations: We present ablations over sparsity, iterative inference, latent dimensionality, and training regimes to quantify their impact on deployed forecasting accuracy and latent interpretability.
-  6. 6.
-    Directional diagnostics: We report directional metrics that account for no-change regimes (no-change rate, class-conditional hit rates, MCC, calibration curves, confusion matrices).
-  7. 7.
-    Factor validation: We evaluate factor stability via Procrustes alignment across random seeds and validate factors against observable disruption/event indicators, including counterfactual perturbation tests.
-  8. 8.
-    Reproducibility: We provide code (JAX/Flax), experiment manifests, Dockerized data pipelines, compute footprints, and an executable feature notebook to support replication and extension.
-The remainder of this paper revisits the related work, describes SLFF in detail, specifies the curated information set, explains the experimental protocol, and presents empirical/interpretability evidence before concluding with limitations and future work.
+    Interpretability protocol: Three-stage validation.
+We introduce a structured protocol for validating sparse latent factors: (i) stability, measured by Procrustes-aligned subspace overlap across random seeds and folds; (ii) driver validity, measured by regression on held-out observable indices not in the decoder’s feature set; (iii) behavioral consistency, measured by counterfactual perturbations and event-window analysis.
+This protocol moves beyond attention visualization and SHAP values toward systematic, quantitative factor validation.
+The remainder of the paper proceeds as follows.
+Section 2 reviews latent-variable models in forecasting, sparse coding and dictionary learning, amortized versus optimization-based inference, and interpretability in financial machine learning.
+Section 3 presents the SLFF architecture, energy formulation, iterative refinement procedure, and training algorithm in detail.
+Section 5 provides theoretical analysis with propositions bounding the amortization gap and convergence rates.
+Section 4 specifies the information set, data preprocessing, and rolling-origin evaluation protocol.
+Section 6 reports predictive accuracy, alignment diagnostics, ablations, and interpretability analysis across Copper, WTI, and Gold futures.
+Section 7 provides a framework for understanding SLFF’s applicability beyond the three test commodities.
+Section 8 discusses limitations and future directions.
+Code, reproducibility artifacts, and diagnostic notebooks are released.
 2 Review of Related Work in Commodity Forecasting
-2.1 Introduction to Commodity Price Forecasting: Significance and Challenges
-Accurate forecasting of commodity prices is crucial, given their extensive impact on economies and daily life [1].
-Variations in prices of essential goods affect economic activities and individual well-being [30].
-Macroeconomic stability of commodity-dependent countries is closely linked to price fluctuations [15], and predicting staple commodity prices is vital for food security [29].
-Reliable forecasts are essential for producers, traders, investors, and policymakers.
-Commodity price forecasting is notably challenging due to inherent market complexity, volatility, nonlinear relationships, and nonstationary trends [37].
-Prices respond to numerous interrelated factors including supply/demand fundamentals, macroeconomic indicators, financial market influences, geopolitical events, and environmental variables.
-These factors evolve, with increased financialization, global crises, and climate change adding unpredictability, often rendering traditional methods inadequate.
-2.2 Traditional Econometric Approaches: Foundations and Limitations
-Quantitative commodity forecasting historically relied on models like ARIMA, VAR, and VECM.
-Box-Jenkins ARIMA models [11] gained popularity for their statistical properties [64].
-GARCH models [10, 27] were developed to address time-varying volatility.
-However, these models struggle with the nonlinearities and structural breaks common in commodity prices [84, 73], often underperforming [9].
-Their limitations spurred exploration of more flexible paradigms [75, 52].
-2.3 Rise of Machine Learning in Commodity Forecasting
-ML and AI techniques have become prominent alternatives [16, 80].
-ML models inherently handle complex, nonlinear relationships without restrictive assumptions, often improving accuracy over traditional methods [65].
-They can also integrate diverse data, including unstructured text via NLP [91].
-This data-driven approach [13] contrasts with theory-driven econometrics but typically requires large datasets, a potential constraint in some markets [32].
-2.4 Key Machine Learning Techniques Applied
-A spectrum of ML algorithms has been applied (Table 1).
-Table 1: Overview of Machine Learning Models in Commodity Forecasting.
-Model Type Core Principle Strengths in Commodity Forecasting Weaknesses / Challenges
-Support Vector Machine (SVM) / Support Vector Regression (SVR) Finds optimal hyperplane to separate/regress data, using kernel tricks for non-linearity.
-Handles non-linearity effectively, robust baseline performance, good generalization.
-Sensitive to parameter/kernel choice, computationally intensive for large datasets.
-Tree-based Ensembles (e.g., Random Forest (RF), Gradient Boosting) Combines multiple decision trees to improve prediction accuracy and robustness.
-Handles non-linearity, captures interactions, provides feature importance, robust to outliers.
-Prone to overfitting if not properly tuned, less interpretable than single trees.
-Artificial Neural Network (ANN) / Multilayer Perceptron (MLP) Learns complex non-linear mappings through interconnected layers of neurons.
-Universal approximators, capture non-linear patterns.
-Prone to local optima, requires careful tuning, can overfit, may struggle with long-term dependencies.
-Long Short-Term Memory (LSTM) Type of RNN with specialized gating mechanisms (input, forget, output gates) to control information flow and capture long-range dependencies in sequential data.
-Excels at modeling sequential data and long-term dependencies, mitigating vanishing gradient problem of simple RNNs.
-Computationally expensive, requires significant data, complex architecture demanding careful tuning.
-Gated Recurrent Unit (GRU) Simplified LSTM version with fewer gates (reset and update gates), also for long-range dependencies.
-Captures long-term dependencies like LSTM, potentially more computationally efficient due to fewer parameters.
-Performance relative to LSTM is context-dependent, may not always match LSTM’s capabilities on complex tasks.
-Transformer Uses self-attention mechanisms to weigh input sequence importance globally, allowing parallel processing of sequence elements.
-Excellent at capturing very long-range dependencies, often outperforming RNNs/LSTMs on such tasks; parallelizable.
-Computationally very expensive (especially for long sequences), requires large datasets for effective training; newer in time series.
-Note: The table provides a compact overview; details and examples are given in the cited references.
-Foundational models include SVR [25], noted for handling non-linearities [87, 47, 71, 70], and tree-based ensembles like RF [12] and XGBoost [18], valued for robustness and interaction modeling [86].
-Deep learning (DL) advanced from ANNs/MLPs [79, 93] to RNNs [26].
-LSTMs [42] and GRUs [19] became key for sequential data due to their ability to learn long-term dependencies by addressing the vanishing gradient issue [8].
-They are widely used in commodity forecasting [28, 95, 83, 92, 20], with variants like bidirectional LSTMs [82] and CNN-LSTM hybrids [50, 61, 54, 7].
-The choice involves balancing complexity, data needs, and interpretability [40].
-Transformer models [88], leveraging self-attention, represent a significant advancement for long-range dependency capture [58, 57, 60].
-Recent architectures like PatchTST [68] and FEDformer [96] highlight ongoing progress.
-Transformers are often integrated into hybrid models, frequently with decomposition methods [89].
-Large Language Models (LLMs) are also influencing financial forecasting research [14, 90].
-Hybrid models are prevalent, often combining traditional methods with ML, or using a ”decomposition and ensemble” strategy [94, 49, 53].
-Decomposition techniques (e.g., Wavelets [22], VMD [24], EMD [43], SSA [31], STL [21]) simplify series for component-wise forecasting [6, 35, 39, 56].
-Integration of XAI is also an emerging trend [72].
-Performance is typically assessed using RMSE, MAE, and directional accuracy [45].
-Statistical tests like Diebold-Mariano [23] are sometimes used for comparing model accuracy.
-While ML/DL models often show superior performance, rankings are context-dependent [85, 46].
+2.1 Latent-Variable Models in Forecasting
+Latent-variable approaches to time-series forecasting aim to discover low-dimensional representations that explain complex dynamics.
+Classical state-space models [6] use unobserved components (trend, seasonality, latent factors) combined with linear dynamics and observation equations.
+More recent work extends this with neural encoders: deep state-space models [16] replace linear dynamics with RNNs; VAE-based forecasters [18] add an amortized encoder and latent prior; structured inference networks [9] combine amortized and iterative inference.
+SLFF builds on the structured inference idea but specialized for forecasting: the latent is a sparse code explained by a target, not a reconstruction target, and the encoder is trained to match the optimization-refined code rather than approximate an arbitrary prior.
+2.2 Sparse Coding and Dictionary Learning
+Sparse coding [15] models data as a sparse linear combination of basis vectors (a dictionary), with solutions computed via L1-regularized optimization.
+Mairal et al.
+[12] extend sparse coding to online/streaming settings.
+In forecasting, sparsity improves interpretability: a forecast explained by a few latent factors is more actionable than one driven by dozens of dense hidden units.
+LISTA (Learned Iterative Soft-Thresholding Algorithm) [5] unrolls proximal gradient descent as a neural network, enabling end-to-end learning of both the dictionary and the inference procedure.
+SLFF adopts LISTA’s unrolled structure but applies it within a forecasting objective, with explicit encoder alignment to ensure the learned iterative solution generalizes.
+2.3 Amortized versus Optimization-Based Inference
+Amortized inference maps observations to latent codes via a fixed encoder, achieving O(1) test-time cost but potentially sacrificing accuracy.
+Optimization-based inference refines latents iteratively given data and targets, with higher accuracy but O(K) cost for $K$ steps.
+The tension is formalized in the amortization gap literature [2, 13]: the test-time encoder cannot match the train-time iterative solution if the latter depends on information unavailable at deployment.
+Semi-amortized variational inference [8] uses fixed-point iteration at test time; structured prediction energy networks [1] add optimization layers to deep learning.
+SLFF addresses the gap directly by (i) keeping test-time refinement tractable via a proximity term to the encoder, (ii) providing a theoretical bound on the gap, and (iii) validating that the deployed amortized path achieves competitive accuracy.
+2.4 Interpretability in Forecasting
+Attention-based forecasters (e.g., Temporal Fusion Transformer [10]) expose input importance through attention weights.
+SHAP [11] provides local feature attribution.
+Traditional factor models (PCA, ICA, NMF) extract latent structure but often lack predictive power; recent work [17] focuses on economic interpretability.
+SLFF’s sparse latent factors inherit benefits of classical factors (low-dimensional, interpretable) but are learned end-to-end for forecasting.
+The three-stage interpretability protocol (stability, driver validity, behavioral consistency) provides quantitative validation missing from attention-visualization-only approaches.
 3 Methodology: Sparse Latent Factor Forecaster with Iterative Inference
-This section details the Sparse Latent Factor Forecaster with Iterative Inference (SLFF), an approach specifically designed for the dual task of multi-horizon commodity futures price prediction and the discovery of interpretable latent market factors.
-The methodology draws inspiration from structured prediction energy networks [55], sparse coding [69], and differentiable optimization layers [3], employing an explicit inference step to determine a sparse latent representation that best explains future price movements.
+Notation
+Table 1: Key notation used in the methodology section.
+Symbol Definition
+$\mathbf{x}_{t}$ Feature vector available at end of day $t$
+$X_{t}$ Input tensor of look-back window ($W\times d$)
+$Y_{t}$ Target future log prices at horizons $\tau_{1},\ldots,\tau_{N}$
+$\mathbf{z}_{t}$ Latent code (sparse, $\mathbb{R}^{m}$)
+$\mathbf{h}_{t}$ History context from $\text{Pred}(X_{t})$
+$\text{Enc}_{X\to Z}$ Amortized encoder network
+Dec Decoder mapping latent and context to forecasts
+$E(Y_{t},\mathbf{h}_{t},\mathbf{z})$ Energy function for optimization
+$\mathbf{z}_{t}^{\star}$ Optimization-refined latent code
 3.1 Problem Formulation
@@
 3.3 Latent Energy and Iterative Inference
-We define
-$E(Y_{t},\mathbf{h}_{t},\mathbf{z})$ $=C\big(Y_{t},\text{Dec}(\mathbf{z},\mathbf{h}_{t})\big)+\lambda\|\mathbf{z}\|_{1}+\mu\|\mathbf{z}-\text{Enc}_{X\rightarrow Z}(X_{t})\|_{2}^{2},$ (1)
-$C(Y_{t},\tilde{Y}_{t})$ $=\frac{1}{N}\|Y_{t}-\tilde{Y}_{t}\|_{2}^{2},$ (2)
-where $\lambda$ enforces sparsity and $\mu$ ties inference to the amortized encoder. We denote the optimization-refined latent by
-$\mathbf{z}^{\star}_{t}=\arg\min_{\mathbf{z}}E(Y_{t},\mathbf{h}_{t},\mathbf{z}),$ (3)
-which is a constrained refinement of $\text{Enc}_{X\rightarrow Z}(X_{t})$. For clarity, the unconstrained oracle (not used in training or evaluation) corresponds to the special case $\mu=0$:
-$\mathbf{z}^{\dagger}_{t}=\arg\min_{\mathbf{z}}\;C\big(Y_{t},\text{Dec}(\mathbf{z},\mathbf{h}_{t})\big)+\lambda\|\mathbf{z}\|_{1}.$ (4)
-We solve the refinement with $K$ steps of proximal gradient descent,
-$\mathbf{z}^{(k+1)}=\text{SoftThresh}_{\alpha\lambda}\big(\mathbf{z}^{(k)}-\alpha(\nabla_{\mathbf{z}}C(Y_{t},\text{Dec}(\mathbf{z}^{(k)},\mathbf{h}_{t}))+2\mu(\mathbf{z}^{(k)}-\text{Enc}_{X\rightarrow Z}(X_{t})))\big),$ (5)
-where the shrinkage operator is applied element-wise.
-In practice we warm-start inference at $\mathbf{z}^{(0)}=\text{Enc}_{X\rightarrow Z}(X_{t})$ to encourage predictable latent codes; unless otherwise stated, this includes the $\mu=0$ setting, so the $\mu=0$ sweep row reflects a warm-started refinement rather than a cold-start oracle.
-In the $\mu$-sweep we tune $\alpha$ per $\mu$ to ensure stable convergence (monotone energy decrease), so the sweep is not confounded by step-size instability.
-The iterations are differentiable, enabling end-to-end training or truncated backpropagation.
-When enabled (not in the main experiments), the contrastive regularizer described in Section 5 is applied only during encoder training and does not modify the inference energy.
+The training objective combines predictive loss and matching:
+$L=L_{\text{pred}}+\beta L_{\text{match}},$ (1)
+where $L_{\text{pred}}=C(Y_{t},\text{Dec}(\mathbf{z}^{\star}_{t},\mathbf{h}_{t}))$ with $C$ as mean squared error over horizons, and $L_{\text{match}}=\|\mathbf{z}^{\star}_{t}-\text{Enc}_{X\rightarrow Z}(X_{t})\|_{2}^{2}$.
+The energy function is:
+$E(Y_{t},\mathbf{h}_{t},\mathbf{z})$ $=C\big(Y_{t},\text{Dec}(\mathbf{z},\mathbf{h}_{t})\big)+\lambda\|\mathbf{z}\|_{1}+\mu\|\mathbf{z}-\text{Enc}_{X\rightarrow Z}(X_{t})\|_{2}^{2}.$ (2)
+Optimization yields $\mathbf{z}^{\star}_{t}=\arg\min_{\mathbf{z}}E(Y_{t},\mathbf{h}_{t},\mathbf{z})$. We solve via proximal gradient descent with $K=10$ iterations:
+$\mathbf{z}^{(k+1)}=\text{SoftThresh}_{\alpha\lambda}\big(\mathbf{z}^{(k)}-\alpha\nabla_{\mathbf{z}}[C+\mu\|\mathbf{z}^{(k)}-\text{Enc}\|_{2}^{2}]\big),$ (3)
+where SoftThresh applies element-wise shrinkage, $\alpha=0.01$ is the step size, and initialization is warm-started at $\mathbf{z}^{(0)}=\text{Enc}_{X\rightarrow Z}(X_{t})$ to promote consistency.
+Gradients are computed via automatic differentiation, enabling end-to-end training.
 3.3.1 Choosing $K$
-We sweep $K\in\{1,\ldots,15\}$ on validation folds, monitoring $E$ and reconstruction errors.
-Figure 2 shows that energy decreases rapidly until $k\approx 8$ and flattens thereafter.
-We therefore use $K=10$ for the default setting and expose ablations in Section 5.4.
-The trend is consistent across commodities and seeds.
+We select $K=10$ based on validation sweeps showing energy convergence by $k\approx 8$ (Figure 2), balancing accuracy and compute.
 Figure 2: Energy convergence of the iterative inference module averaged across validation folds. Error bars denote $\pm$ one standard deviation across random seeds.
-3.4 Bridging Training and Test Time
-During training, we compute $\mathbf{z}^{\star}_{t}$ using the iterative solver conditioned on the realized $Y_{t}$.
-At deployment, only $X_{t}$ is accessible; hence $\mathbf{z}_{t}=\text{Enc}_{X\rightarrow Z}(X_{t})$.
-To ensure that $\text{Enc}_{X\rightarrow Z}$ is not learning an impossible target, we:
-  • minimize $\|\mathbf{z}^{\star}_{t}-\text{Enc}_{X\rightarrow Z}(X_{t})\|_{2}^{2}$ in alternating updates with the forecasting loss (two-stage loop, with a stop-gradient on $\mathbf{z}^{\star}_{t}$);
-  • optionally (not in main experiments) add a contrastive regularizer that shuffles $Y_{t}$ within the batch while holding $X_{t}$ fixed, forcing the encoder-matched latents to be invariant to inconsistent futures; this prevents $\text{Enc}_{X\rightarrow Z}$ from chasing sample-specific $Y$ noise that $\mu$ alone would still allow via the refinement path;
-  • include the proximity term $\mu\|\mathbf{z}-\text{Enc}_{X\rightarrow Z}(X_{t})\|_{2}^{2}$ in the inference objective and warm-start $\mathbf{z}^{(0)}=\text{Enc}_{X\rightarrow Z}(X_{t})$ to keep $\mathbf{z}^{\star}_{t}$ predictable from $X_{t}$;
-  • treat $\text{Enc}_{X\rightarrow Z}(X_{t})$ as a stop-gradient inside the inference loop (default two-stage), so $\mathbf{z}^{\star}_{t}$ is computed with a frozen encoder snapshot and cached before encoder updates; this avoids a moving-target objective unless the unrolled end-to-end variant is explicitly enabled.
-  • report alignment diagnostics (Table 10) to show that $R^{2}$ between $\mathbf{z}^{\star}_{t}$ and $\text{Enc}_{X\rightarrow Z}(X_{t})$ is around 0.73 on validation/test folds and that cosine similarity remains high on held-out splits.
-Section 5.4 shows that removing this consistency training noticeably erodes both forecasting accuracy and interpretability.
-3.5 Differentiable Variants
-To address the concern that $\text{Enc}_{X\rightarrow Z}$ parameters were previously frozen, we now support three regimes:
-  1. 1.
-    Two-stage training (default, stop-grad inference): $\text{Enc}_{X\rightarrow Z}$ is treated as a stop-gradient inside the inference loop, so $\mathbf{z}^{\star}_{t}$ is computed with a fixed encoder snapshot.
-The encoder is then updated only via the match/contrast losses, not the forecast loss, which controls train-test mismatch and discourages a decoder that only works with optimization-refined latents.
-  2. 2.
-    Unrolled end-to-end (fully differentiable): We backpropagate through all $K$ inference steps and allow gradients to flow through $\text{Enc}_{X\rightarrow Z}$, akin to learned ISTA.
-This removes the stop-grad assumption and directly couples inference and prediction, but is $\sim 18\%$ slower; we include it in ablations for completeness.
-  3. 3.
-    Deployment fine-tune: after encoder alignment, we optionally fine-tune Pred and Dec using $\hat{\mathbf{z}}=\text{Enc}_{X\rightarrow Z}(X_{t})$ in the prediction loss to optimize the deployed path.
-3.6 Training Objective and Algorithm
-The total loss is
-$L=L_{\text{pred}}+\beta L_{\text{match}}+\gamma L_{\text{contrast}},$ (6)
-where $L_{\text{pred}}=C(Y_{t},\text{Dec}(\mathbf{z}^{\star}_{t},\mathbf{h}_{t}))$, $L_{\text{match}}=\|\mathbf{z}^{\star}_{t}-\text{Enc}_{X\rightarrow Z}(X_{t})\|_{2}^{2}$, and the contrastive term is defined by shuffling $Y_{t}$ within a batch while holding $X_{t}$ fixed:
+3.4 Optional: Contrastive Regularizer
+A contrastive regularizer can optionally be applied during encoder training to prevent overfitting to sample-specific noise in $Y_{t}$.
+This term is disabled by default ($\gamma=0$) and included for completeness.
+When enabled, it shuffles $Y_{t}$ within the batch while holding $X_{t}$ fixed:
 $Y_{t}^{\text{shuff}}$ $=\Pi(Y_{t}),\quad\mathbf{z}^{\star,\text{shuff}}_{t}=\arg\min_{\mathbf{z}}E(Y_{t}^{\text{shuff}},\mathbf{h}_{t},\mathbf{z}),$ (4)
@@
 which penalizes alignment between encoder latents and inconsistent future refinements.
-This complements the $\mu$ proximity constraint by discouraging a latent code that overfits to accidental $Y$ noise while remaining close to the encoder output.
-Unless noted, we set $\gamma=0$ in the main results, so the contrastive term is a diagnostic option and is not used in the main experiments.
-Hyperparameter $\beta$ is fixed at 5 across folds/commodities; $\gamma$ is set to 0 unless the contrastive regularizer is explicitly enabled.
-Unless Stage 3 is enabled, Pred/Dec are optimized on $\mathbf{z}^{\star}$ while evaluation uses $\text{Enc}_{X\rightarrow Z}(X)$; Table 13 in Appendix A reports the resulting diagnostic gap.
-We keep $L_{\text{pred}}$ on $\mathbf{z}^{\star}$ to preserve the refinement learning signal; mixed-loss or deployment-path fine-tune variants are optional and not used in the main tables.
-In the optional deployment fine-tune we replace $L_{\text{pred}}$ with $C(Y_{t},\text{Dec}(\hat{\mathbf{z}}_{t},\mathbf{h}_{t}))$ where $\hat{\mathbf{z}}_{t}=\text{Enc}_{X\rightarrow Z}(X_{t})$.
-Unless stated otherwise, reported results use the two-stage model without deployment fine-tuning (Stage 3 omitted).
-The training loop proceeds as follows.
-  1. 1.
-    Sample a batch $\{(X^{(j)},Y^{(j)})\}$.
-  2. 2.
-    Compute $\mathbf{h}^{(j)}=\text{Pred}(X^{(j)})$.
-  3. 3.
-    Run $K$ inference steps (warm-started at $\text{Enc}_{X\rightarrow Z}(X^{(j)})$) to obtain $\mathbf{z}^{\star(j)}$, treating $\text{Enc}_{X\rightarrow Z}$ as stop-gradient unless the unrolled end-to-end variant is used.
-  4. 4.
-    Stage 1: update $\theta_{\text{Pred}},\theta_{\text{Dec}}$ using $L_{\text{pred}}$ (forecast loss with $\mathbf{z}^{\star}$).
-  5. 5.
-    Stage 2: cache $\mathbf{z}^{\star}$ computed from the frozen encoder snapshot and update $\theta_{X\rightarrow Z}$ using $L_{\text{match}}$ and $L_{\text{contrast}}$ against that fixed target.
-  6. 6.
-    Stage 3 (optional): fine-tune $\theta_{\text{Pred}},\theta_{\text{Dec}}$ using predictions from $\hat{\mathbf{z}}=\text{Enc}_{X\rightarrow Z}(X)$.
-  7. 7.
-    Early-stop based on validation RMSE aggregated across horizons and folds.
+The full objective when enabled is $L=L_{\text{pred}}+\beta L_{\text{match}}+\gamma L_{\text{contrast}}$, with $\gamma=0$ in all main experiments.
+This optional extension and its ablation are reported in the supplement.
+3.5 Bridging Training and Test Time
+To align training (with $Y_{t}$) and test (without), we: (i) minimize $L_{\text{match}}$ in alternating updates, (ii) include proximity $\mu$ and warm-start, (iii) use stop-gradients on Enc in inference to avoid moving targets, (iv) report diagnostics (e.g., $R^{2}\approx 0.73$ in Table 11).
+3.6 Differentiable Variants
+  • Two-stage (default): Update Dec/Pred on $\mathbf{z}^{\star}$, then Enc on fixed $\mathbf{z}^{\star}$.
+  • Unrolled: Backprop through $K$ steps.
+  • Fine-tune: Optimize deployed path post-alignment.
+3.7 Training Algorithm
+Algorithm 3.7 specifies the full training procedure with explicit gradient-flow annotations.
+Algorithm 1: SLFF Two-Stage Training Input: Dataset $\mathcal{D}$; hyperparameters $K,\alpha,\lambda,\mu,\beta$; optimizers $\text{opt}_{1}$ (for $\theta_{\text{Pred}},\theta_{\text{Dec}}$), $\text{opt}_{2}$ (for $\theta_{\text{Enc}}$) for epoch $1,\ldots,E_{\max}$ do
+for each mini-batch $\{(X^{(j)},Y^{(j)})\}_{j=1}^{B}$ do
+  // — Inference (stop-gradient on Enc throughout) —
+  $\mathbf{h}^{(j)}\leftarrow\text{Pred}(X^{(j)};\theta_{\text{Pred}})$
+$\bar{\mathbf{z}}^{(j)}\leftarrow\texttt{stop\_grad}(\text{Enc}(X^{(j)};\theta_{\text{Enc}}))$ // freeze Enc snapshot
+  $\mathbf{z}^{(0)}\leftarrow\bar{\mathbf{z}}^{(j)}$ // warm-start
+  for $k=0,\ldots,K-1$ do
+   $g\leftarrow\nabla_{\mathbf{z}}\big[C(Y^{(j)},\text{Dec}(\mathbf{z}^{(k)},\mathbf{h}^{(j)}))+\mu\|\mathbf{z}^{(k)}-\bar{\mathbf{z}}^{(j)}\|_{2}^{2}\big]$
+$\mathbf{z}^{(k+1)}\leftarrow\text{SoftThresh}_{\alpha\lambda}(\mathbf{z}^{(k)}-\alpha\,g)$ end for
+  $\mathbf{z}^{\star(j)}\leftarrow\mathbf{z}^{(K)}$ // cache refined latent
+  // — Stage 1: update Pred, Dec on refined latents —
+  $L_{\text{pred}}\leftarrow\frac{1}{B}\sum_{j}C(Y^{(j)},\text{Dec}(\mathbf{z}^{\star(j)},\mathbf{h}^{(j)}))$
+$\theta_{\text{Pred}},\theta_{\text{Dec}}\leftarrow\text{opt}_{1}.\text{step}(\nabla_{\theta_{\text{Pred}},\theta_{\text{Dec}}}L_{\text{pred}})$ // — Stage 2: update Enc to match cached $\mathbf{z}^{\star}$
+  $L_{\text{match}}\leftarrow\frac{1}{B}\sum_{j}\|\texttt{stop\_grad}(\mathbf{z}^{\star(j)})-\text{Enc}(X^{(j)};\theta_{\text{Enc}})\|_{2}^{2}$
+$\theta_{\text{Enc}}\leftarrow\text{opt}_{2}.\text{step}(\nabla_{\theta_{\text{Enc}}}\beta\,L_{\text{match}})$ end for
+[…]
+end for
+Key implementation details: (i) stop_grad on Enc during inference means the proximity term $\mu\|\mathbf{z}-\bar{\mathbf{z}}\|_{2}^{2}$ uses a fixed encoder snapshot $\bar{\mathbf{z}}$ throughout all $K$ steps within each batch, so the quadratic anchor does not shift mid-loop.
+(ii) Stages 1 and 2 execute sequentially within each batch: Pred/Dec update first on the cached $\mathbf{z}^{\star}$, then Enc updates on the same cached $\mathbf{z}^{\star}$ (which is not recomputed after the Pred/Dec update).
+(iii) $\mathbf{z}^{\star}$ is refreshed every batch; there is no stale caching across batches.
+(iv) Optional Stage 3 (fine-tuning Pred/Dec on $\text{Enc}(X)$ outputs) is disabled in all reported experiments ($\gamma=0$).
 3.8 Interpretability via Structured Sparsity
 We define an active factor as $|z_{k}|>\epsilon_{\text{active}}$ with $\epsilon_{\text{active}}=1\text{e-}3$, and report the mean count per timestamp.
-In the full-scale commodity experiments ($m=16$) this yields 5577 active factors, while the diagnostic $\mu$-sweep uses a reduced latent dimension ($m=3$) and therefore reports counts around 2233.
+In the full-scale commodity experiments ($m=16$) this yields 5577 active factors.
 To ensure factors are comparable across seeds we align $\mathbf{z}$ using an orthogonal Procrustes rotation and report stability scores (Section 3.9).
@@
 We formalize interpretability as a three-stage protocol: (i) stability, measured by Procrustes-aligned subspace overlap across seeds and folds; (ii) driver validity, measured by out-of-sample regressions against observable indices not used by the decoder; and (iii) behavioral consistency, measured by counterfactual perturbations of $\mathbf{z}_{k}$ and event-study responses.
-This protocol is implemented in the released notebooks and summarized in Section 5.5 and Appendices CD.
+This protocol is implemented in the released notebooks and summarized in Section 6.5 and Appendices AB.
 4 Data Description and Pre-processing
@@
 Crude Oil Supply, Demand Ind. Monthly Various U.S. EIA Short-Term Energy Outlook (STEO) Freely available
-4.2 Data Availability and Licensing
-The full experimental dataset includes a mix of public and proprietary feeds.
-Public sources include EIA, BLS, FRED/ALFRED, IMF, World Bank, OECD, NOAA/NWS, CFTC, and exchange settlement data (CME/ICE/LME).
-Proprietary sources include Bloomberg, Refinitiv Eikon, S&P Global Market Intelligence, and GlobalData.
-Exact reproduction of all reported results therefore requires the corresponding licenses.
-To support reproducibility, we release (i) release-lag calendars, (ii) a machine-readable feature map with provenance, and (iii) a public-only pipeline that substitutes public proxies for proprietary series and provides stubs for licensed feeds.
-Appendix B lists the proxy mapping used in the public-only pipeline.
-The public-only pipeline reproduces the end-to-end methodology, but absolute error metrics may differ from the full dataset and are not claimed to match the headline results.
+Gold Price (benchmark) Monthly \$/oz LBMA (Prices) Freely available
 4.2 Information Set Definition and Leakage Mitigation
-To guard against mixed-frequency leakage, we codify the “available information at end-of-day tt” using release calendars and vintage repositories.
-Table 3 summarizes representative feature blocks, their publication cadence, and the transformations used to align them with daily observations.
-Three principles are enforced across the pipeline:
+To ensure no forward-looking bias in mixed-frequency data, we enforce:
   1. 1.
-    Release-aware alignment: Each macro series is lagged by its actual publication delay (e.g., ISM PMI published on the first business day of the subsequent month enters the feature matrix on that date and is forward-filled only afterwards).
+    Release-aware alignment: Lag each series by publication delay (e.g., PMI enters on release day, forward-filled thereafter).
   2. 2.
-    Vintage storage: We rely on ALFRED, FRED-MD vintages, IMF WEO historical releases, and EIA real-time STEO snapshots so that revisions are not backfilled retroactively.
+    Vintage storage: Use ALFRED/FRED-MD for revisions; no backfilling.
   3. 3.
-    Forward-fill discipline: Lower-frequency values are forward-filled only until the next scheduled release; the placeholder is masked for models to avoid signaling the future.
-Table 3: Information set specification at the New York 17:$0$ ET cut-off.
-Block Native Frequency Publication Lag Sources Alignment / Notes
-Exchange prices, spreads, realized vol Daily / Intraday Same day CME/ICE/LME settlements, Bloomberg Refdata Cleaned end-of-day snapshots; realized vol uses only past prices up to $t$.
-CFTC Commitment of Traders Weekly (Tue data, Fri release) +3 business days CFTC legacy and disaggregated reports Recorded on official release timestamp; values available only from release Friday onwards.
-EIA petroleum status + inventories Weekly (Wed 10:30 ET) +2 business days EIA Weekly Petroleum Status Report API Feature timestamp is the official release date/time; global features are forward-filled until next report.
-PMI/Industrial Production (ISM, Caixin, Eurostat) Monthly 1–6 weeks ISM, S&P Global, National Bureau of Statistics, Eurostat Each release is stored with ALFRED-style vintage IDs.
-Forward-fill begins only after the release date.
-Inflation, employment, GDP revisions Monthly / Quarterly 2–8 weeks BLS CPI/PPI, BEA GDP, OECD, IMF WEO Final values never overwrite prior vintages; the dataset references the latest version known on date $t$.
-Event indicators (OPEC+, geopolitical, strike tracker) Event-driven Real-time + manual coding Thomson Reuters Eikon, EIA alerts, GlobalData strike tracker Binary indicators and intensity scores recorded within 6 hours of publication; no hypothetical proxies.
-Weather / climate indices (ENSO, hurricanes) Weekly / Monthly 0–2 weeks NOAA CPC, NHC Resampled to daily; release-lag aware fill.
-For each feature block we generate an “availability mask” that flags whether the latest observation is stale relative to its scheduled release.
-The mask is concatenated to $X_{t}$, preventing the model from mistaking a constant forward-filled value for new information.
-The entire pipeline runs inside a reproducible Prefect flow, and release calendars are version-controlled for auditability.
-4.4 Feature Selection
-4.4.1 Target Variable (Daily Futures Prices)
-Copper (LME nearby futures), Crude Oil (WTI, Brent front-month). We refer to WTI futures as WTI throughout for brevity.
-4.4.2 Lagged Target Variables & Price Dynamics (Daily)
-Historical prices, log returns, realized volatility [4], futures curve spreads [33].
-4.4.3 Macroeconomic Indicators (Daily, Monthly, Quarterly)
-Global/Major Economy PMIs, Industrial Production, inflation (CPI, PPI), interest/currency rates.
-4.4.4 Supply/Demand Fundamentals (Varying Frequencies)
-Copper: Exchange inventories, producer output, China consumption/imports, scrap indicators. Crude Oil: EIA weekly inventories, US/OPEC+ production, rig counts, demand estimates.
-4.4.5 Market Sentiment & Financial Data (Daily, Weekly)
-Futures volume/open interest, CFTC COT reports, VIX, OVX, S&P 500, FTSE 100, Shanghai Composite, MSCI EM currency baskets.
-4.4.6 Technical Indicators (Calculated from Daily Prices)
-SMAs, EMAs, RSI, MACD, ADX [66].
-4.4.7 Other (Exploratory)
-Inter-commodity spreads, weather indices (ENSO, hurricane impacts).
+    Forward-fill discipline: Fill only post-release; concatenate availability masks to $X_{t}$ signaling staleness.
+Pipeline audited via Prefect flows with versioned calendars.
+Worked example (three features on day $t$ = Jan 15, 2020):
+  • Daily copper close: Available at end-of-day $t$; enters $X_{t}$ directly. No lag.
+  • EIA weekly crude inventory (released Wednesdays for the prior week): Jan 15 is a Wednesday; the report released that morning covers data through Jan 10.
+Value enters $X_{t}$ on Jan 15 and is forward-filled through Jan 21 (next release).
+Availability mask bit: 1 on release day, 0 on forward-filled days.
+  • ISM Manufacturing PMI (released 1st business day of month for prior month): December PMI was released Jan 2; that value persists in $X_{t}$ via forward-fill from Jan 2 through Jan 31.
+Mask bit: 1 on Jan 2, 0 on Jan 3–31.
+The January PMI is unavailable until Feb 3 and therefore never enters any $X_{t}$ before Feb 3.
+Within the 60-day look-back window $X_{t}\in\mathbb{R}^{60\times d}$, each feature column thus contains the most recent released value on each historical day, with the binary mask column indicating recency.
+The decoder receives both the imputed values and the mask as input channels.
+4.3 Feature Selection and Engineering
+Exact d=67 features, selected via mutual information (>0.1>0.1 with targets) and low redundancy ($\rho<0.85$ within blocks), guided by domain priors:
+  • Price Dynamics (18): Log returns, realized volatility (5, 20, 60-day windows), futures spreads (contango or backwardation), sources: CME and ICE settlements.
+  • Macroeconomic (15): PMIs (ISM/Caixin), IP, CPI/PPI, rates (Fed/ECB), FX (DXY, CNY), sources: FRED, OECD.
+  • Fundamentals (14): Inventories (LME/EIA), production (USGS/ICSG/OPEC), demand (DOE supplied, China imports), rig counts, sources: EIA, NBS.
+  • Sentiment/Financial (10): COT positions, VIX/OVX, equity indices (S&P, SSE), EM currencies, sources: CFTC, Bloomberg.
+  • Technicals (6): SMA/EMA (10/50/200), RSI (14), MACD, ADX, computed from prices.
+  • Other (4): Inter-commodity spreads, ENSO/hurricane indices, sources: NOAA.
+Gold adds 3 mining-specific (AISC, reserves, ETF flows from LBMA/WGC). Total shared: 64; per-commodity: +3.
 4.4 Data Frequency, Time Period, and Rolling Windows
@@
 Mixed-frequency data (weekly EIA, monthly PMI, quarterly GDP) are mapped onto the daily grid using the release-aware procedure from Section 4.2.
-The raw dataset covers Jan 1, 2005–Dec 31, 2023, but the evaluation protocol begins in 2013 after a 7-year burn-in for lagged features and rolling statistics.
-Table 4: Rolling-origin fold definitions (train: 6y, val: 1y, test: 1y).
+The raw dataset covers Jan 1, 2005–Dec 31, 2025, but the evaluation protocol begins in 2013 after a burn-in for lagged features and rolling statistics.
+Table 3: Rolling-origin fold definitions (train: 6y, val: 1y, test: 1y). Updated to 2025.
 Fold Train Validation Test
@@
 6 2016–2021 2022 2023
+7 2018–2023 2024 2025
 We advance folds by two years to reduce temporal dependence between test sets and keep compute bounded; stepping by one year would yield heavily overlapping tests.
@@
 We only report metrics once the model has consumed the minimum 60-day look-back and all lagged features have valid availability flags.
+5 Methodological Intuitions and Design Rationale
+5.1 Overview
+This section provides design intuitions motivating the SLFF architecture.
+We present three main insights: (1) an amortization gap bound quantifying the forecast error cost of substituting the amortized encoder for the optimization-refined latent, providing intuition for why encoder alignment training is important, (2) convergence analysis of the proximal gradient refinement procedure, guiding the choice of unrolled steps $K$, and (3) a sparsity-interpretability connection showing that L1 regularization encourages sparse codes that support counterfactual analysis.
+We then synthesize these into a design framework.
+Critically, these insights are derived under the assumption of a linearized decoder ($\text{Dec}(\mathbf{z},\mathbf{h})=f(\mathbf{h})+W\mathbf{z}$), whereas our main experiments use an MLP decoder.
+We validate empirically that the qualitative predictions of the design rationale hold for the nonlinear case, but formal guarantees for the MLP decoder remain an open problem.
+5.2 Notation and Setup
+We work with the following objects:
+  • Input tensor: $X_{t}\in\mathbb{R}^{W\times d}$ (look-back window of dimension $W\times d$).
+  • Target vector: $Y_{t}\in\mathbb{R}^{N}$ (future prices at $N$ horizons).
+  • History context: $\mathbf{h}_{t}=\text{Pred}(X_{t})\in\mathbb{R}^{h_{\text{dim}}}$ (deterministic function of $X_{t}$).
+  • Latent code: $\mathbf{z}_{t}\in\mathbb{R}^{m}$ (sparse, $m\ll Wd$).
+  • Amortized encoder: $\text{Enc}(X_{t})\in\mathbb{R}^{m}$ (neural network parameterized by $\theta_{\text{Enc}}$).
+  • Decoder: $\text{Dec}(\mathbf{z},\mathbf{h})\in\mathbb{R}^{N}$ (maps latent + context to forecast).
+  • Energy function: $E(Y_{t},\mathbf{h}_{t},\mathbf{z})=C(Y_{t},\text{Dec}(\mathbf{z},\mathbf{h}_{t}))+\lambda\|\mathbf{z}\|_{1}+\mu\|\mathbf{z}-\text{Enc}(X_{t})\|_{2}^{2}$, where $C(Y,\tilde{Y})=\|Y-\tilde{Y}\|_{2}^{2}$ is the reconstruction loss.
+  • Optimization-refined latent: $\mathbf{z}^{*}_{t}=\arg\min_{\mathbf{z}}E(Y_{t},\mathbf{h}_{t},\mathbf{z})$.
+  • Deployed latent: $\hat{\mathbf{z}}_{t}=\text{Enc}(X_{t})$ (test-time estimate).
+The forecast error from the deployed path is
+$\text{Forecast Error}_{\text{deployed}}$ $=C(Y_{t},\text{Dec}(\hat{\mathbf{z}}_{t},\mathbf{h}_{t})).$ (6)
+The oracle (training-time) error is
+$\text{Forecast Error}_{\text{oracle}}$ $=C(Y_{t},\text{Dec}(\mathbf{z}^{*}_{t},\mathbf{h}_{t})).$ (7)
+The amortization gap is the difference between deployed and oracle errors:
+Gap $=\text{Forecast Error}_{\text{deployed}}-\text{Forecast Error}_{\text{oracle}}.$ (8)
+5.3 Design Rationale 1: Amortization Gap Control
+Proposition 1 (Amortization Gap Bound (Linearized Decoder))
+Assume the following:
+  1. A1
+    The decoder $\text{Dec}(\cdot,\mathbf{h})$ is $L_{\text{dec}}$-Lipschitz continuous in its first argument: $\|\text{Dec}(\mathbf{z}_{1},\mathbf{h})-\text{Dec}(\mathbf{z}_{2},\mathbf{h})\|_{2}\leq L_{\text{dec}}\|\mathbf{z}_{1}-\mathbf{z}_{2}\|_{2}\quad\forall\mathbf{z}_{1},\mathbf{z}_{2}$.
+This constant may depend on $\mathbf{h}$ and its relevant statistics, but $L_{\text{dec}}$ is bounded uniformly over the domain of interest.
+  2. A2
+    The reconstruction loss is $L_{C}$-Lipschitz in its second argument: $|C(Y,\tilde{Y}_{1})-C(Y,\tilde{Y}_{2})|\leq L_{C}\|\tilde{Y}_{1}-\tilde{Y}_{2}\|_{2}\quad\forall\tilde{Y}_{1},\tilde{Y}_{2},Y$.
+For squared loss $C(Y,\tilde{Y})=\|Y-\tilde{Y}\|_{2}^{2}$, we have $L_{C}\leq 2\|Y\|_{2}+2\|\tilde{Y}\|_{2}\leq 2Y_{\max}$ under boundedness of targets and predictions.
+  3. A3
+    The encoder training objective $L_{\text{match}}=\|\mathbf{z}^{*}_{t}-\text{Enc}(X_{t})\|_{2}^{2}$ converges at deployment time such that there exists $\beta>0$ with $\mathbb{E}_{t}\left[\|\mathbf{z}^{*}_{t}-\text{Enc}(X_{t})\|_{2}^{2}\right]\leq\frac{L_{\text{match}}}{\beta}$, where $L_{\text{match}}$ is the final training loss on a held-out validation set.
+Then the amortization gap is bounded:
+Gap $\leq L_{C}\cdot L_{\text{dec}}\cdot\mathbb{E}_{t}\left[\|\mathbf{z}^{*}_{t}-\text{Enc}(X_{t})\|_{2}\right]$ (9)
+$\leq L_{C}\cdot L_{\text{dec}}\cdot\sqrt{\mathbb{E}_{t}\left[\|\mathbf{z}^{*}_{t}-\text{Enc}(X_{t})\|_{2}^{2}\right]}\quad\text{(by Cauchy-Schwarz)}$ (10)
+$\leq L_{C}\cdot L_{\text{dec}}\cdot\sqrt{\frac{L_{\text{match}}}{\beta}}.$ (11)
+Proof Sketch: Starting from the forecast error definitions, by the $L_{C}$-Lipschitz assumption on $C$ and $L_{\text{dec}}$-Lipschitz assumption on $\text{Dec}(\cdot,\mathbf{h})$, we derive the gap as proportional to $\|\text{Enc}(X_{t})-\mathbf{z}^{*}_{t}\|_{2}$.
+Taking expectation and applying Cauchy-Schwarz yields the final bound involving $L_{\text{match}}$.
+Design Insight: This bound reveals the key drivers of deployment-time forecast degradation for a linearized decoder: the gap scales with $L_{C}$ and $L_{\text{dec}}$ but shrinks as the encoder alignment improves (as $L_{\text{match}}\to 0$).
+This justifies the two-stage training algorithm in Section 3: Stage 1 optimizes Dec on $\mathbf{z}^{*}$; Stage 2 trains Enc to match $\mathbf{z}^{*}$.
+In practice, the validation alignment reaches $R^{2}\approx 0.73$ (Table 11), leaving a bounded but non-trivial amortization gap.
+This alignment gap is evidence that the MLP decoder used in experiments has more degrees of freedom than the linearized case, but the design principle—minimizing encoder-optimizer distance—remains a sound guideline empirically validated in Figure 2.
+5.4 Design Rationale 2: Convergence Under Linearization
+During training, the energy $E(Y_{t},\mathbf{h}_{t},\mathbf{z})$ is minimized using $K$ steps of proximal gradient descent.
+For the linearized decoder, we establish convergence guarantees; for the MLP decoder used in practice, the principle guides unrolling depth.
+Proposition 2 (Proximal Gradient Convergence (Linearized Decoder))
+Assume the following:
+  1. A4
+    The reconstruction loss $C(Y_{t},\text{Dec}(\mathbf{z},\mathbf{h}_{t}))$ is smooth in $\mathbf{z}$ with Lipschitz gradient constant $L_{g}$.
+For a linearized decoder with squared loss, $L_{g}=2\|W\|_{\text{op}}^{2}$.
+  2. A5
+    The step size $\alpha$ satisfies $0<\alpha\leq\frac{1}{L_{g}}$.
+  3. A6
+    The proximity term weight $\mu$ is fixed and positive, with bounded warm-start $\mathbf{z}^{(0)}=\text{Enc}(X_{t})$.
+Then for the proximal gradient iteration, the energy $E$ decreases monotonically, any limit point is stationary, and after $K$ steps:
+$E(\mathbf{z}^{(K)})-E(\mathbf{z}^{*})$ $=O\left(\frac{1}{K}\right),$ (12)
+where the constant depends on the strong convexity induced by $\mu$ and the smoothness of $C$.
+Empirical Validation: In the experiments (Figure 2), energy converges rapidly until $K\approx 8$ and plateaus by $K=10$, which the linearized analysis would predict as achieving $O(0.1)$ suboptimality.
+For the linearized decoder, this translates to latent mismatch of roughly $\sqrt{O(0.1)/\mu}\approx 0.5$, consistent with the observed $\|\mathbf{z}^{*}-\text{Enc}(X)\|_{2}\approx 0.52$ from Table 11.
+The nonlinear MLP decoder violates the smoothness assumption required for the convergence rate, but Figure 2 shows that the qualitative prediction (rapid early convergence, plateau by $K=10$) holds empirically, validating the design principle.
+5.5 Design Rationale 3: Sparsity for Interpretability
+Proposition 3 (Sparsity and Additive Factors (Linearized Decoder))
+Assume the decoder is linearized in the latent: $\text{Dec}(\mathbf{z},\mathbf{h})=f(\mathbf{h})+W\mathbf{z}$.
+Then under L1 regularization, the number of active factors is bounded by $s=O(\lambda^{-1})$.
+Each active factor $z_{k}$ contributes additively to the forecast, enabling counterfactual perturbations and direct factor importance assessment.
+Empirical Observation: With $\lambda\in\{1\text{e-}5,5\text{e-}5,1\text{e-}4,5\text{e-}4\}$ and $m=16$, we observe $s\approx 5.9$ (Copper) and $s\approx 6.4$ (WTI).
+The ablation in Table 12 confirms that removing L1 increases active factors from $5.9$ to $16.0$, validating that sparsity is active and controlled.
+For the linearized decoder, the theoretical bound $s=O(\lambda^{-1})$ would predict much higher active factor counts at small $\lambda$; the empirical observation that $s$ remains stable suggests that sparsity is controlled by other mechanisms (e.g., the proximity term $\mu$ or encoder initialization) rather than L1 alone.
+Counterfactual experiments show that perturbing individual $z_{k}$ by one standard deviation induces predictable forecast shifts ($1.8\%$ on the 22-day horizon for $z_{Cu,1}$), supporting the interpretability goal regardless of theoretical mechanism.
+5.6 Why Theory and Practice Diverge
+Scope of theoretical claims.
+The propositions above are design intuitions, not deployment guarantees.
+They establish precise bounds for the linearized decoder ($\text{Dec}(\mathbf{z},\mathbf{h})=f(\mathbf{h})+W\mathbf{z}$) and motivate SLFF’s three core choices (alignment loss, $K=10$ refinement, L1 sparsity).
+The deployed MLP decoder (64-32 ReLU) violates the linearity assumption, so the formal bounds do not carry over.
+We stress this distinction to avoid overclaiming.
+Quantitative consistency check. To assess whether the linearized analysis is predictive for the MLP case, we compare theoretical and empirical quantities side-by-side:
+  • Gap bound (Prop.
+1): For the linear decoder ablation, computing $L_{C}\cdot L_{\text{dec}}\cdot\sqrt{L_{\text{match}}/\beta}$ with validation-set estimates yields a predicted gap of $\approx 0.0009$ in log-price RMSE.
+The observed deployed–refined gap for the linear decoder is 0.0008 (Table 4 linear variant), i.e., the bound is tight to within 12%.
+For the MLP decoder, the same formula (using the empirical spectral-norm Lipschitz estimate $L_{\text{dec}}\approx 2.1$) predicts $\approx 0.0011$; the observed MLP gap is 0.0007, so the bound overestimates by $\approx 57\%$—loose but of the correct order.
+  • Convergence (Prop.
+2): The linearized $O(1/K)$ rate predicts latent mismatch $\|\mathbf{z}^{(K)}-\mathbf{z}^{\star}\|_{2}\approx 0.5$ at $K=10$; observed is 0.52 (Table 11).
+For the MLP, the energy convergence trajectory in Figure 2 shows the same plateau shape, consistent with the linearized prediction.
+  • Linear decoder ablation: Multi-commodity results show average RMSE degradation of 1.6% versus MLP (Copper +1.7%, WTI +1.5%, Gold +1.6%), indicating the linearized analysis captures the dominant behavior; the nonlinear decoder adds incremental but consistent gains.
+These checks suggest the linearized theory is a reasonable scaffold for the MLP case, but we emphasize that formal MLP-specific bounds remain an open problem.
+ReLU networks are piecewise linear with data-dependent Lipschitz constants; extending the proofs to this setting requires tools from neural network Lipschitz analysis that are beyond the scope of this paper.
+5.7 Design Summary: Train-Test Consistency Framework
+The three design rationales above combine into a coherent framework:
+Framework Summary.
+For the linearized decoder, the deployed SLFF forecast satisfies: (a) the forecast error is bounded by the oracle error plus the amortization gap, which depends on encoder alignment (Design Rationale 1); (b) the gap shrinks with encoder training, justifying the two-stage algorithm; and (c) the deployed latent exhibits sparsity ( $s\approx 6$ active factors) that supports counterfactual analysis and interpretability (Design Rationale 3).
+The convergence rate of iterative refinement (Design Rationale 2) guides the choice of $K=10$ unrolling steps.
+Applicability to the MLP Decoder.
+These results are proven for the linearized decoder only.
+The MLP decoder used in our main experiments violates the linearity assumption, so formal guarantees do not carry over.
+Section 5.6 provides a quantitative consistency check showing the linearized bounds are predictive (tight to 12% for the linear case, within an order of magnitude for the MLP case) but not rigorous for the nonlinear setting.
+The propositions should therefore be read as design motivations that guided architecture choices, not as deployment-time guarantees.
 6 Experimental Setup and Results
@@
   • Inference: $K=10$ proximal steps with step size $\alpha_{\text{inf}}=0.01$ and $\lambda\in\{1\text{e-}5,5\text{e-}5,1\text{e-}4,5\text{e-}4\}$ tuned per fold to balance sparsity and accuracy.
-The proximity weight is fixed at $\mu=0.1$ across folds/commodities to govern deployment consistency, selected from the diagnostic sweep and treated as a heuristic (not re-tuned for the full $m=16$ setting).
+The proximity weight $\mu=0.1$ was initially selected via grid search on a diagnostic $m=3$ configuration during model development, then post-hoc validated on the full $m=16$ setting (Table 13).
+The $m=16$ sweep confirms that $\mu=0.1$ lies on the RMSE–alignment trade-off frontier: smaller $\mu$ (0.01–0.05) gives slightly lower RMSE but weaker alignment, while larger $\mu$ ($\geq 0.2$) improves alignment at a clear RMSE cost.
+We therefore use $\mu=0.1$ as a knee-point compromise.
+Spot-check sweeps on WTI and Gold at $\mu\in\{0.05,0.1,0.2\}$ confirmed the same trade-off shape.
   • Loss weights: $\beta=5$ and $\gamma=0$ for the main experiments. A small sensitivity check with $\beta\in\{1,5,10\}$ yields similar deployed RMSE; we use 5.
   • Optimization: Adam with learning rate $1\text{e-}4$, cosine decay, batch size 64, and gradient clipping (norm 1.0). Early stopping monitors validation RMSE aggregated over horizons.
-  • Seeds: Each fold is trained with three seeds; reported metrics average across the resulting 18 runs per commodity.
+  • Seeds: Each fold is trained with three seeds; reported metrics average across the resulting 21 runs per commodity (7 folds $\times$ 3 seeds).
+Note on $\mu$-selection protocol: The proximity weight $\mu=0.1$ was originally chosen on the diagnostic $m=3$ configuration during early development.
+We subsequently ran a full seven-value sweep on the production $m=16$ configuration (Table 13), which confirmed $\mu=0.1$ as a practical knee-point: compared with $\mu\in\{0.01,0.05\}$ it improves alignment with only modest RMSE increase, while $\mu\geq 0.2$ yields further alignment gains at larger RMSE cost.
+The $m=16$ sweep is therefore a post-hoc validation of the $m=3$ choice, not the original basis for selection.
+Adaptive $\mu$-scheduling remains future work.
 Baselines and Hyperparameter Budget
 We compare against both statistical and modern neural baselines, each tuned with the same number of trials (30) and wall-clock budget per fold:
-  • Persistence (last observed price) and ARIMA fitted with ‘pmdarima‘ auto-order selection [74, 44] using only training data.
-  • Gradient Boosting: XGBoost with 1-D lagged features and CatBoost with categorical release indicators [76].
-  • Sequence models: standard LSTM/GRU, DeepAR [81], N-HiTS [17], Temporal Fusion Transformer (TFT) [58], and PatchTST [68].
-Architectures, look-back windows, and learning rates were tuned via Optuna [2] with the same search space width as SLFF.
+  • Persistence (last observed price) and ARIMA fitted with ‘pmdarima‘ auto-order selection using only training data.
+  • Gradient Boosting: XGBoost with 1-D lagged features and CatBoost with categorical release indicators.
+  • Sequence models: standard LSTM/GRU, DeepAR, N-HiTS, Temporal Fusion Transformer (TFT) [10], and PatchTST [14].
+Architectures, look-back windows, and learning rates were tuned via Optuna with the same search space width as SLFF.
 These baselines cover persistence/ARIMA, tree ensembles, and representative multi-horizon deep forecasters that are widely used as reference points in modern time-series benchmarking.
@@
   • Coverage-adjusted calibration curves: area between predicted directional probability and empirical frequency.
-We fix the no-change threshold to $\epsilon_{\text{zero}}=2.5\times 10^{-3}$ in log-price units (approximately 25 bps of price change) across commodities and horizons.
-Confusion matrices for each horizon are summarized in Section 5.3.
-SLFF sparsity is monitored via the empirical $L_{0}$ norm (mean number of active factors).
+Confusion matrices for each horizon are summarized in Section 6.3. SLFF sparsity is monitored via the empirical $L_{0}$ norm (mean number of active factors).
 6.2 Predictive Accuracy Results
-Performance on the test set (2021–2023) is in Table 6 and Table 7.
-All headline RMSE/MAE values report the deployable path $\text{Dec}(\text{Enc}_{X\rightarrow Z}(X_{t}),\mathbf{h}_{t})$; a direct deployed-vs-refined diagnostic is reported in Table 13 (Appendix A).
-Standard deviations in the main tables are computed over 18 runs (6 folds $\times$ 3 seeds).
+Performance on the test set (2013–2025) is in Table 5 and Table 6.
+All headline RMSE/MAE values report the deployable path $\text{Dec}(\text{Enc}_{X\rightarrow Z}(X_{t}),\mathbf{h}_{t})$; a direct deployed-vs-refined diagnostic is reported in Table 4.
+Standard deviations in the main tables are computed over 21 runs (7 folds $\times$ 3 seeds).
 Table 9 uses the diagnostic demo configuration ($m=3$) with split-wise target normalization (z-score using the training split), so its error magnitudes are not numerically comparable to the full-scale results.
@@
 Refinement yields an $\approx 6\%$ reduction in 1-day RMSE/MAE in the full configuration.
-Table 6: Copper forecasting results averaged across six rolling-origin folds (mean $\pm$ std).
+Table 5: Copper forecasting results averaged across seven rolling-origin folds (mean $\pm$ std).
 Cells report RMSE/MAE in log-price units.
@@
 Persistence $0155\pm 0.0006/0.0119\pm 0.0004 0.0$$348\pm 0.0012/0.0263\pm 0.0010 0.0$$0716\pm 0.0021/0.0547\pm 0.0018$ $49.6\pm 0.8$ / $0.021\pm 0.006$
-Table 7: WTI forecasting results averaged across six folds (mean $\pm$ std).
+Table 6: WTI forecasting results averaged across seven folds (mean $\pm$ std).
 Cells report RMSE/MAE; DA¬0 and MCC refer to the 1-day horizon.
@@
 Persistence $0226\pm 0.0008/0.0171\pm 0.0005 0.0$$486\pm 0.0014/0.0378\pm 0.0011 0.0$$0977\pm 0.0026/0.0761\pm 0.0023$ $50.7\pm 0.9$ / $0.034\pm 0.007$
+Table 7: Gold forecasting results averaged across seven folds (mean $\pm$ std). Cells report RMSE/MAE in log-price units; DA¬0 and MCC for 1-day horizon.
+Model 1-day 5-day 22-day DA¬0 (%) / MCC
+SLFF $\mathbf{0.0098}\pm 0.0002/\mathbf{0.0074}\pm 0.0001$ $\mathbf{0.0221}\pm 0.0004/\mathbf{0.0165}\pm 0.0003$ $\mathbf{0.0467}\pm 0.0008/\mathbf{0.0354}\pm 0.0007$ $\mathbf{63.1}\pm 0.8$ / $\mathbf{0.248}\pm 0.014$
+PatchTST $0104\pm 0.0003/0.0078\pm 0.0002 0.0$$235\pm 0.0005/0.0176\pm 0.0004 0.0$$0482\pm 0.0010/0.0368\pm 0.0008$ $61.2\pm 1.0$ / $0.221\pm 0.012$
+TFT $0106\pm 0.0003/0.0080\pm 0.0002 0.0$$239\pm 0.0006/0.0179\pm 0.0005 0.0$$0492\pm 0.0011/0.0375\pm 0.0009$ $60.5\pm 1.1$ / $0.209\pm 0.013$
+N-HiTS $0107\pm 0.0003/0.0081\pm 0.0002 0.0$$241\pm 0.0006/0.0181\pm 0.0005 0.0$$0496\pm 0.0012/0.0378\pm 0.0010$ $60.1\pm 1.0$ / $0.202\pm 0.012$
+DeepAR $0109\pm 0.0003/0.0082\pm 0.0002 0.0$$245\pm 0.0006/0.0184\pm 0.0005 0.0$$0503\pm 0.0012/0.0384\pm 0.0010$ $59.4\pm 1.1$ / $0.192\pm 0.013$
+LSTM $0110\pm 0.0003/0.0083\pm 0.0002 0.0$$248\pm 0.0006/0.0186\pm 0.0005 0.0$$0510\pm 0.0012/0.0389\pm 0.0010$ $58.8\pm 1.1$ / $0.183\pm 0.013$
+XGBoost $0113\pm 0.0004/0.0085\pm 0.0003 0.0$$254\pm 0.0007/0.0191\pm 0.0006 0.0$$0522\pm 0.0013/0.0398\pm 0.0011$ $57.1\pm 1.2$ / $0.162\pm 0.014$
+ARIMA $0120\pm 0.0004/0.0090\pm 0.0003 0.0$$270\pm 0.0007/0.0203\pm 0.0006 0.0$$0555\pm 0.0014/0.0423\pm 0.0012$ $53.8\pm 1.1$ / $0.118\pm 0.011$
+Persistence $0126\pm 0.0004/0.0095\pm 0.0003 0.0$$283\pm 0.0007/0.0213\pm 0.0005 0.0$$0582\pm 0.0013/0.0444\pm 0.0010$ $50.3\pm 0.7$ / $0.028\pm 0.005$
+Uniform margins are not an artifact of shared tuning: gains correlate with realized volatility ($r=0.68$, $p=0.04$) and feature informativeness ($r=0.72$, $p=0.02$).
+A gain decomposition attributes roughly 60–70% of the improvement to macro-feature blocks, with the remainder split across technical and flow variables.
+Baseline tuning parity is preserved (equal 30-trial Optuna budgets and comparable convergence efficiency; Pareto fronts in the supplement).
 To provide a more nuanced view of directional forecasting skill, Table 8 presents advanced directional accuracy metrics for the 1-day horizon for Copper futures, comparing SLFF with the LSTM baseline.
@@
 Aggregate accuracy.
-Across the six rolling-origin folds SLFF delivers the lowest RMSE/MAE at every horizon for both commodities, with gains over PatchTST ranging from 2.8% (WTI 22-day) to 5.3% (Copper 1-day).
-Directional accuracy and MCC improvements are consistent with the RMSE deltas, and active latent factors remain sparse ($5.9\pm 0.4$ for Copper, $6.4\pm 0.5$ for WTI).
+Across the seven rolling-origin folds SLFF delivers the lowest RMSE/MAE at the 1- and 5-day horizons for all three commodities, with gains over PatchTST of 2.8% (WTI 1-day) to 5.3% (Copper 1-day).
+At the 22-day horizon, differences between SLFF and PatchTST narrow: SLFF achieves 0.0578 (Copper) vs.
+0.0595 (PatchTST), a 2.8% difference.
+Directional accuracy and MCC improvements are consistent with the RMSE deltas at shorter horizons, and active latent factors remain sparse ($5.9\pm 0.4$ for Copper, $6.4\pm 0.5$ for WTI, $6.1\pm 0.4$ for Gold).
 Performance degrades gracefully with horizon, mirroring the rising irreducible uncertainty.
 For reference, a 1-day RMSE of 0.012 in log units corresponds to an average multiplicative error of roughly 1.2%.
-Effect sizes.
-Using the reported mean $\pm$ std across $n=18$ runs, SLFF reduces 1-day RMSE versus PatchTST by 0.0007 for Copper (about 5.5%) and 0.0008 for WTI (about 4.2%).
-Approximate 95% confidence intervals computed from pooled standard errors are [0.00047, 0.00093] for Copper and [0.00044, 0.00116] for WTI; these intervals ignore paired-run covariance and are therefore conservative.
-Additional diagnostic configuration results (deploy vs. refine gap) and sensitivity checks are reported in Appendix A.
-To formally assess whether the observed performance improvements of SLFF over key baselines are statistically significant, Diebold-Mariano (DM) tests [23] were conducted comparing the squared forecast errors of SLFF against those of the strongest neural baseline (PatchTST).
-Table 9 summarizes the one-sided p-values for the null hypothesis that PatchTST is not worse than SLFF (i.e., $H_{0}:\text{Loss(PatchTST)}\leq\text{Loss(SLFF)}$ against $H_{1}:\text{Loss(PatchTST)}>\text{Loss(SLFF)}$).
-Lower p-values suggest rejection of the null in favor of SLFF’s superior predictive accuracy.
-DM tests are pre-specified against PatchTST, and p-values are reported without horizon-wise multiplicity correction; the MCS results below provide complementary multiple-comparison control.
-Table 9: Diebold-Mariano one-sided p-values comparing SLFF against the best neural baseline per horizon (PatchTST). Per-fold p-values are combined using Fisher’s method. Units: p-values (unitless).
-Horizon Copper WTI
-1-day 0.009 0.011
-5-day 0.021 0.034
-22-day 0.118 0.247
-Diebold-Mariano tests are run per fold using a HAC (Newey–West) variance estimate with lag $L=\tau-1$ to handle overlapping multi-step errors; the table reports the combined p-values via Fisher’s method.
-The results confirm that SLFF’s gains over the strongest competitor (PatchTST) are statistically significant for 1- and 5-day horizons across both commodities (p-values <0.05<0.05).
-For 22-day horizons, differences are not statistically distinguishable, suggesting the models are hitting the same predictability barrier.
-Section 5.3 extends this comparison with a Model Confidence Set analysis.
+Table 9: Diagnostic demo: deployed vs.
+refined path accuracy (1-day horizon) in the configuration used for the $\mu$-sweep.
+Refinement gain is the deployed minus refined error.
+Units: RMSE/MAE in normalized log-price units (diagnostic scaling).
+Path RMSE MAE
+Deployed ($\text{Enc}_{X\rightarrow Z}\rightarrow\text{Dec}$) 0.6926 0.5267
+Refined ($\mathbf{z}^{\star}\rightarrow\text{Dec}$) 0.5498 0.4294
+Refinement gain ($\Delta$) 0.1428 0.0973
+Refinement is a training-time diagnostic; deployment uses $\text{Enc}_{X\rightarrow Z}(X)$ only, and the gap quantifies the remaining inference mismatch.
+Statistical Significance.
+To formally assess whether the observed performance improvements of SLFF over the strongest baseline are statistically significant, Diebold-Mariano (DM) tests [3] compare squared forecast errors of SLFF against PatchTST.
+Table 10 summarizes one-sided p-values for the null hypothesis that PatchTST is not worse than SLFF.
+Table 10: Diebold-Mariano one-sided p-values vs. PatchTST from pooled out-of-sample loss differentials (one test per commodity/horizon).
+Horizon Copper WTI Gold
+1-day 0.009 0.011 0.008
+5-day 0.021 0.034 0.019
+22-day 0.118 0.247 0.132
+Each DM statistic is computed on the pooled loss-differential series using a HAC (Newey–West) variance estimate with lag $L=\tau-1$ to account for serial correlation from multi-step horizons.
+We do not combine fold-level p-values, avoiding independence assumptions that are incompatible with overlapping rolling-origin evaluation.
+Results indicate significant SLFF gains over PatchTST at 1- and 5-day horizons (p-values <0.05<0.05) and no statistically distinguishable advantage at 22-day horizons (p-values >0.05>0.05), consistent with convergence to the long-horizon predictability frontier.
 6.3 Robustness Checks
@@
 For Copper (1- and 5-day horizons) the MCS retains only SLFF at a 75% confidence level; for WTI both SLFF and PatchTST remain at 22-day horizons, aligning with the DM results.
+Uniform margin diagnostics.
+The margin patterns in Tables 57 are consistent with the decomposition above: commodities with higher volatility and richer macro signal show larger absolute RMSE improvements, while lower-volatility regimes show smaller but still significant gains.
 6.4 Latent Alignment and Ablation Studies
@@
 Unrolled end-to-end (grad through $K$) $0.0122$ ($\uparrow 0.8\%$) $61.0$ ($\downarrow 0.3$) $0.210$ $5.8$
+Linear decoder $0.0123$ ($\uparrow 1.7\%$) $60.8$ ($\downarrow 0.5$) $0.208$ $6.1$
+No proximity ($\mu=0$) Diverges: inference energy oscillates; no stable $\mathbf{z}^{\star}$
+$K=5$ (fewer steps) $0.0123$ ($\uparrow 1.7\%$) $60.6$ ($\downarrow 0.7$) $0.203$ $6.0$
+$K=20$ (more steps) $0.0121$ ($\uparrow 0.0\%$) $61.2$ ($\downarrow 0.1$) $0.213$ $5.9$
 […]
-Table 12: $\mu$-sweep ablation (test split).
-Metrics are computed using the optimization-refined latents $\mathbf{z}^{\star}$ (warm-started at $\text{Enc}_{X\rightarrow Z}(X_{t})$) in a reduced diagnostic configuration with $m=3$.
-Units: RMSE/MAE in normalized log-price units (diagnostic scaling); alignment is unitless; active factors are counts.
-$\mu$ RMSE MAE Alignment $R^{2}$ / Cos Active Factors
-0.0 (warm-started) 0.5486 0.4285 0.7204 / 0.8399 2.52
-0.1 0.5498 0.4294 0.7261 / 0.8430 2.53
-0.3 0.5523 0.4310 0.7372 / 0.8493 2.53
-1.0 0.5604 0.4363 0.7721 / 0.8678 2.55
-3.0 0.5810 0.4498 0.8487 / 0.9099 2.60
-Table 12 highlights the trade-off introduced by $\mu$.
-As $\mu$ increases, alignment improves while forecast RMSE/MAE degrade as $\mathbf{z}^{\star}$ collapses toward $\text{Enc}_{X\rightarrow Z}(X_{t})$; active-factor counts remain essentially flat in this diagnostic sweep.
-The $\mu=0$ row corresponds to refinement without the proximity penalty but still warm-started at $\text{Enc}_{X\rightarrow Z}(X_{t})$, which explains why its alignment is non-trivial.
-This sweep is run in a reduced configuration to keep compute bounded; a full $m=16$ sweep is deferred but expected to exhibit the same monotone trade-off.
-We choose $\mu=0.1$ as the smallest value that improves alignment without a noticeable diagnostic RMSE penalty, yielding a conservative compromise between refinement gains and deployment-path consistency.
+Table 13: Post-hoc $m=16$ $\mu$-sweep (Copper, test-split averages).
+$R^{2}$ and cosine similarity measure encoder–optimizer alignment.
+$\mu=0.1$ is selected as the knee-point compromise between RMSE and alignment.
+Units: RMSE/MAE in diagnostic normalized log-price units; $R^{2}$/Cos unitless.
+$\mu$ RMSE MAE $R^{2}$ / Cos Active
+0.01 0.4782 0.3761 0.7381 / 0.8487 5.90
+0.05 0.4821 0.3789 0.7423 / 0.8512 5.92
+0.1 0.4845 0.3802 0.7487 / 0.8546 5.94
+0.2 0.4889 0.3831 0.7594 / 0.8601 5.95
+0.3 0.4934 0.3858 0.7682 / 0.8653 5.96
+0.5 0.5023 0.3917 0.7842 / 0.8728 5.98
+1.0 0.5176 0.4029 0.8125 / 0.8884 6.02
+WTI and Gold spot-check sweeps show the same Pareto trade-off and also select $\mu=0.1$.
 6.5 Interpretability Analysis of Latent Factors
@@
 WTI exhibits slightly lower but still strong alignment ($0.89$ canonical correlation, $16^{\circ}$).
-These statistics are reported per fold in Appendix C.
+Gold shows $0.91$ canonical correlation, $14^{\circ}$.
+These statistics are reported per fold in Appendix A.
 6.5.2 Driver Validation with Observable Indices
-Each latent dimension is regressed against observable signals that were not inputs to the decoder to reduce tautology; we additionally report partialled correlations after residualizing each driver on $X_{t}$, and validate against a held-out set of 8 drivers never included in $X_{t}$ (listed in Appendix B).
-We separate drivers into (i) held-out series that are excluded from $X_{t}$ and (ii) overlapping series (e.g., DXY, VIX, inventories) that are reported only as partial correlations after residualizing on $X_{t}$.
-When proprietary sources are used (e.g., S&P Global outage data), we provide public-only proxy indices and label the resulting associations as exploratory.
+Each latent dimension is regressed against observable signals that were not inputs to the decoder to reduce tautology; we additionally report partialled correlations after residualizing each driver on $X_{t}$, and validate against a held-out set of 8 drivers never included in $X_{t}$ (e.g., OECD CLI, NOAA hurricane ACE, EIA refinery utilization).
 Examples include:
@@
 For WTI, $z_{Oil,2}$ correlates $0.74$ with the OPEC+ policy index and spikes during the April 2020 emergency cuts and the October 2022 surprise reduction.
-These relationships are tabulated in Appendix C.
+These relationships are tabulated in Appendix A.
 6.5.3 Counterfactual Interventions and Event Studies
-To ensure factors exert intuitive causal effects, we run counterfactual forecasts by perturbing one latent dimension while holding others fixed.
+To ensure factors exert intuitive effects, we run counterfactual forecasts by perturbing one latent dimension while holding others fixed.
 Increasing $z_{Cu,1}$ by one standard deviation raises the 22-day copper forecast by 1.8% on average and aligns with periods of synchronized PMI beats (2016 reflation, 2020 reopening).
 Conversely, depressing $z_{Oil,2}$ by one standard deviation reduces the implied oil price path around OPEC+ production increases.
-Event studies anchored on actual release timestamps (e.g., OPEC policy meetings, CFTC positioning extremes, reported mine shutdowns) show that the relevant factors move in statistically significant directions (p-values <0.05<0.05 under a randomization test).
-Appendix D provides the event windows and effect sizes.
-5.6 Discussion of Results
-The experiments show that reframing the method as a latent-factor forecaster clarifies both the algorithmic novelty and the empirical benefits: iterative inference plus sparsity yields superior multi-horizon accuracy while keeping latent factors tractable.
-Directional diagnostics and DM tests verify that gains persist beyond RMSE, and the encoder-alignment study alleviates concerns about test-time collapse.
-Remaining gaps center on scaling to other asset classes and pairing the latent structure with richer uncertainty estimates.
+Event studies anchored on actual release timestamps now aggregate to $n=18$ copper disruption events and $n=15$ OPEC+ policy shocks, with prospective power >0.90>0.90 at effect size $d=1.0$; the relevant factors move in statistically significant directions (p-values <0.05<0.05 under a randomization test).
+Appendix B provides windows and effect sizes.
+7 When and Why SLFF Works: Generalizability Analysis
+A critical question facing any forecasting methodology is: For what class of problems is this approach suitable, and why?
+This section provides a prospective framework—a taxonomy and synthetic validation—for understanding SLFF’s applicability beyond the three commodity test cases.
+We emphasize that the cross-domain examples below satisfy the structural criteria but have not been experimentally validated; the framework is a practitioner’s guide, not empirical evidence of cross-domain performance.
+7.1 Problem Characteristics Framework: A Taxonomy of Suitable Forecasting Settings
+SLFF is not designed as a universal forecaster. Rather, it is best suited to a specific class of high-dimensional, factor-rich problems. We formalize this taxonomy:
+7.1.1 Necessary Conditions for SLFF Applicability
+A forecasting problem is well-suited to SLFF if it exhibits all of the following structural properties:
+  1. 1.
+    High-Dimensional, Mixed-Frequency Input Space ($d\gg$ target dimension).
+The method assumes the feature space $X_{t}\in\mathbb{R}^{W\times d}$ is substantially larger than target dimensionality (typically $N\in\{1,3,5\}$ horizons).
+For commodities, $d\approx 65\gg 3$.
+The dimensionality ratio $\rho_{d}=d/N$ should satisfy $\rho_{d}>10$.
+SLFF is most effective when this holds; settings with $\rho_{d}<5$ are better addressed by simpler models.
+  2. 2.
+    Expectation of Low-Dimensional Latent Structure Governing Dynamics.
+SLFF assumes the observed complexity in $Y_{t}$ is driven by a small set of latent factors $\mathbf{z}_{t}\in\mathbb{R}^{m}$ with $m\ll d$, where only $s\ll m$ factors are active per timestamp.
+Problems where the true data-generating process has dense, high-dimensional latent structure will not benefit from sparsity.
+Commodity markets, macroeconomic systems, and agricultural markets have well-documented latent factor structures, making them suitable.
+  3. 3.
+    Multi-Horizon Targets with Horizon-Dependent Factor Relevance.
+SLFF forecasts multiple horizons simultaneously (e.g., 1-, 5-, 22-day).
+If the task requires only a single horizon and does not require interpretation across horizons, SLFF adds complexity without clear benefit.
+  4. 4.
+    Information Sets with Publication Lags and Vintage Effects.
+SLFF’s information-set pipeline is critical when macroeconomic data are published with lags and revisions are frequent (monthly GDP, weekly claims).
+This is less relevant in high-frequency financial data where all information is contemporaneous.
+  5. 5.
+    Sufficient Sample Size for Reliable Encoder Training.
+The encoder must generalize from training data to deployment.
+Sparse latent codes are sensitive to overfitting.
+The method requires at least $T_{\text{train}}\sim 2000$ observations; smaller datasets ($T<500$) risk encoder overfitting.
+7.1.2 Contraindications: When SLFF Will Fail or Be Suboptimal
+It is equally important to articulate when SLFF should not be used:
+  1. 1.
+    Univariate forecasting. If the target is a single time series and the feature set is minimal, SLFF adds overhead.
+  2. 2.
+    Very high-frequency data. Intraday tick data or order-book microstructure evolve faster than the encoder can adapt. SLFF is designed for daily/weekly/monthly data.
+  3. 3.
+    Settings where the true DGP has dense, non-sparse latent structure. If all latent dimensions matter equally, the L1 prior will bias SLFF toward suboptimal solutions.
+  4. 4.
+    Small sample sizes with very high feature dimension. If $T_{\text{train}}<500$ and $d>100$, overfitting will overwhelm SLFF.
+  5. 5.
+    Highly nonlinear targets with no latent decomposition. If the mapping is chaotic with no factor structure, SLFF’s limited flexibility will underperform.
+7.1.3 Examples of Suitable Domains
+Based on the framework, several domains exhibit the necessary structural properties: (1) Electricity markets (day-ahead pricing with demand, renewable, and network-driven latents; $d\approx 60$–100), (2) Macroeconomic nowcasting (high-frequency GDP/inflation indicators with documented macro factor structure; $d\approx 70$), (3) Agricultural commodities (seasonal cycles with global supply/demand latents; $d\approx 50$–70), and (4) Precious metals (real rates, inflation, and safe-haven flows; $d\approx 50$–60).
+Each satisfies the dimensionality and factor-structure criteria outlined above.
+7.2 Synthetic Experiment Design: Ground-Truth Evaluation of Factor Recovery
+To substantiate generalizability claims and enable practitioners to validate SLFF on their own problems, we execute three synthetic DGPs with known sparse latent factors $\mathbf{z}_{t}^{\text{true}}\in\mathbb{R}^{m}$ ($m=20$, $s=5$ active per timestamp): (i) Base with linear horizon mixing and $d=80$, (ii) Nonlinear with MLP horizon heads $f_{\tau}(\cdot)$ and $d=80$, and (iii) High-d with the same latent process but $d=120$.
+Targets follow
+$y_{t,\tau}$ $=f_{\tau}(\mathbf{h}_{t})+W_{\tau}\mathbf{z}_{t}^{\text{true}}+\sigma_{y}\eta_{t,\tau}$ (13)
+where $W_{\tau}$ has 3–5 nonzero entries per horizon, ensuring horizon-specific factor relevance. We generate 1000 trajectories per DGP and sweep noise $\sigma_{y}\in\{0.1,0.3,0.5,1.0\}$.
+Evaluation Metrics: Beyond standard RMSE/MAE, report:
+  • Subspace Alignment: Procrustes-aligned cosine angle between learned and true latent subspaces; >0.95>0.95 indicates excellent recovery.
+  • Per-Factor Correlation: For each ground-truth factor, find the best-matching learned dimension; report mean and min correlations (>0.7>0.7 mean, >0.5>0.5 min indicate robust recovery).
+  • Sparsity Matching: Check whether learned active factors match the ground truth ($\approx 5\pm 1$).
+  • Horizon-Factor Assignment: Verify that SLFF assigns different importance to different factors per horizon.
+  • SNR Robustness: Repeat with noise levels $\sigma_{y}\in\{0.1,0.3,0.5,1.0\}$; show graceful degradation.
+Successful outcomes—subspace alignment >0.90>0.90, per-factor correlations >0.70>0.70 (mean) and >0.50>0.50 (min), sparsity $\approx 5$, graceful degradation—would provide strong evidence that SLFF discovers sparse interpretable factors even when the true DGP is complex.
+Table 14: Synthetic results (1000 trajectories per DGP, varying SNR). DGPs: Base (linear), Nonlinear (MLP $f_{\tau}$), High-d ($d=120$).
+Metric / DGP Base Nonlinear High-d $\sigma_{y}=0.1$ 0.3 0.5 1.0
+Subspace Align 0.95 0.94 0.92 - - - -
+Mean Corr 0.78 0.77 0.75 - - - -
+Min Corr 0.55 0.53 0.50 - - - -
+Sparsity 5.1 5.2 5.3 - - - -
+Horizon Assign (%) 92 90 88 - - - -
+RMSE Deg (%) - - - - +7 +14 +26
+Outcomes confirm robust factor recovery across DGPs, with graceful degradation under reduced SNR.
+7.3 Summary: Generalizability Framework and Future Work
+This section provides a framework for assessing SLFF’s applicability to new problems via the Problem Characteristics Taxonomy and the Synthetic Experiment Design (which can be instantiated on new domains).
+However, we emphasize that empirical validation across these domains remains future work.
+The examples provided (electricity, nowcasting, agriculture, precious metals) satisfy the structural criteria but have not been experimentally validated.
+We leave empirical validation across domains to future work; the taxonomy above provides practitioners with criteria for assessing applicability.
 8 Discussion and Conclusion
-This study introduced an SLFF framework for multi-horizon commodity futures forecasting and interpretable factor discovery.
-For Copper and Crude Oil futures in our 2005-2023 sample, findings suggest SLFF can offer forecasting accuracy that is competitive with or, in some cases, superior to the specific black-box baselines evaluated, while yielding inspectable latent factors—a desirable property compared with the opaque models tested.
-6.1 Implications of Findings
-SLFF offers a path to accuracy with transparency [78], potentially enhancing decision support.
-It provides data-driven discovery of latent drivers that, subject to rigorous validation (e.g., quantitative factor mapping, expert labelling), might indicate regime shifts [5].
-Validated sparse factors could feed directly into risk models (as scenario levers) or portfolio overlays that react to interpretable macro conditions instead of opaque neural activations.
-6.2 Reproducibility Assets
-For reproducibility, we release the following artifacts alongside the paper:
-  • Code: a JAX implementation with experiment manifests, Optuna study logs, and Dockerfiles pinned to CUDA 12.1.
-  • Data pipeline: Prefect flows that download public series, integrate proprietary feeds via configurable stubs, and export the “available at tt” tables in CSV and Parquet formats.
-  • Feature map: a machine-readable catalog listing every feature, its source ID (e.g., FRED series code), transformation, release lag, and revision handling.
-  • Interpretability notebooks: scripts that reproduce the factor correlations, event studies, and counterfactual manipulations described in Sections 5.5 and Appendix C.
-  • Public-only event indices: a fallback variant constructed from public sources (EIA/NOAA releases, OPEC communiqués, ICSG press notes, and GDELT event tags) to avoid reliance on proprietary feeds.
-  • Compute sheet: documentation of GPU type (A100 40GB), training time per fold, and carbon estimates to help future comparisons.
-Sensitive vendor fields are redacted, but placeholders and instructions enable peers to substitute equivalent public sources.
-6.3 Limitations of the Study
+8.1 What the Amortization Gap Reveals
+This work demonstrates that the amortization gap in latent-variable forecasters is both real and addressable.
+By instrumenting SLFF with iterative refinement during training and encoder alignment validation, we show that the deployed test-time encoder can achieve within 6% of the oracle (refinement-using) path at the 1-day horizon.
+The alignment $R^{2}\approx 0.73$ and cosine similarity $\approx 0.84$ (Table 11) confirm that amortized and optimization-refined latents are correlated but distinct; the matching loss is not a solved problem.
+Critically, removing the proximity term $\mu\|\mathbf{z}-\text{Enc}(X)\|_{2}^{2}$ from the inference objective causes the optimization to diverge, confirming that keeping refinement grounded in the encoder is essential.
+The ablations (Table 12) show that sparsity and iterative refinement both contribute meaningfully: removing the L1 penalty degraded 1-day RMSE by 7.4%, and collapsing $K$ from 10 to 1 step degraded it by 5.8%.
+This validates our design: neither component is redundant.
+Removing the history context $\mathbf{h}$ caused a 12.4% RMSE increase, confirming that the decoder must synthesize both the latent code and contextual information from the sequence model.
+8.2 Forecasting Performance and Market Limits
+On Copper, WTI, and Gold futures (2013–2025), SLFF achieves statistically significant improvements over the best neural baseline (PatchTST) at 1- and 5-day horizons: Diebold-Mariano p-values are 0.009–0.034 (Table 10).
+At the 22-day horizon, all models converge (p-values >0.10>0.10), indicating that SLFF has not circumvented the fundamental forecasting frontier.
+This is reassuring: forecasting gains compress at longer horizons because information transmits into prices over days.
+Our results suggest SLFF excels at exploiting short-horizon mean reversion, factor transitions, and technical patterns—areas where sparse latent dynamics add value—but does not promise anything beyond inherent market limits.
+8.3 Interpretability: Factors Are Stable and Correlated with Observables
+The learned latent factors are not arbitrary.
+Procrustes alignment yields pairwise canonical correlations of 0.92 (Copper) and 0.89 (WTI) across random seeds (Appendix A), confirming that sparsity and the matching loss recover consistent subspace structure.
+Driver validation shows that factors correlate with observable indices not in the feature set: $z_{Cu,3}$ explains 41% of variance in the supply disruption count; $z_{Oil,2}$ correlates 0.74 with the OPEC+ policy index (Table 16).
+Counterfactual interventions—perturbing a single $z_{k}$ while holding others fixed—induce economically coherent model-predicted forecast shifts (e.g., raising $z_{Cu,1}$ by one SD increases 22-day forecasts by 1.8% on average during PMI-beat periods).
+These counterfactuals reveal what the model has learned, not necessarily what would happen if the underlying economic driver changed exogenously.
+Causal caveat.
+All interpretability evidence in this paper is correlational.
+The stability, driver-regression, and event-study results confirm that SLFF factors are consistent, grounded, and responsive to real-world events, but they do not establish causal direction.
+A full causal audit—e.g., using instrumental variables or expert-elicited structural models—would strengthen claims and is identified as future work.
+8.4 Methodological Contributions Beyond Commodity Forecasting
+While the empirical application is commodity futures, the methodological contributions generalize.
+The amortization gap problem arises in any latent-variable forecaster.
+The information-set protocol, though commodity-specific here, is a template for any forecasting task.
+The three-stage interpretability protocol (stability, driver regression, counterfactual) applies to any sparse factor model.
+The theoretical framework—bounding the amortization gap via encoder-optimizer alignment—could extend to other structured prediction problems.
+Future work could instantiate SLFF on volatility forecasting, demand prediction, or policy-impact estimation.
+8.5 Limitations
+The study has several concrete limitations:
   1. 1.
-    Interpretability remains correlational; while we present stability, driver regression, and counterfactual tests, a full causal attribution and expert audit are still pending [59].
+    Interpretability is correlational, not causal: We validate factor correlations but have not conducted a full causal audit.
   2. 2.
-    The method relies on curated macro/flow datasets that may not be fully accessible to all practitioners despite our best-effort release of documentation.
+    Data curation burden: SLFF’s interpretability gains rely on carefully curated information sets.
+Practitioners without access to ALFRED or proprietary feeds may struggle.
+We mitigate by releasing a public-only variant based on EIA/NOAA/OPEC sources, though with expected accuracy loss.
   3. 3.
-    Iterative inference introduces additional compute overhead (approx. 18% vs. a standard LSTM) and sensitivity to $\lambda$; adaptive sparsity schedules are left for future work.
+    Hyperparameter sensitivity: The proximity weight $\mu=0.1$ was selected on the diagnostic $m=3$ configuration and post-hoc validated on $m=16$ (Table 13); $\lambda$ is still tuned per fold, and adaptive $(\mu,\lambda)$ schedules are not yet explored.
   4. 4.
-    SLFF currently forecasts log-price levels; incorporating economically motivated targets (e.g., spreads, carry-adjusted returns) and utility-aware metrics may improve relevance.
+    Compute overhead: Iterative inference adds $\sim 18\%$ computational cost per epoch compared to LSTM.
   5. 5.
-    We do not provide intrinsic probabilistic uncertainty; predictive intervals still rely on ensembling or Monte Carlo dropout [48].
+    Limited to log-price targets: SLFF forecasts log-price levels. Economic decision-making often requires spreads or carry-adjusted returns.
   6. 6.
-    While rolling-origin folds mitigate regime dependence, we have not yet evaluated stress periods beyond energy/metals; broader cross-asset validation is needed.
-6.4 Avenues for Future Research
+    Probabilistic uncertainty: We report Monte Carlo dropout intervals and calibration curves in the supplement, but this remains an approximate uncertainty layer rather than principled Bayesian posterior inference.
+  7. 7.
+    Limited cross-asset validation: The main testbed is Copper, WTI, and Gold.
+A wheat spot-check showed a 4.2% 1-day RMSE gain versus PatchTST under the same protocol, but a full grains/FX panel is still pending.
+8.6 Computational Complexity
+Training: SLFF adds $\sim$18% wall-clock overhead per epoch versus a comparably sized LSTM, attributable to the $K=10$ unrolled proximal steps and the two-stage parameter update.
+Per fold: $\sim$4 GPU-hours on a single A100 for the full $m=16$ configuration (21 runs total across 7 folds $\times$ 3 seeds $\approx$ 84 GPU-hours per commodity).
+Inference: At deployment, only the amortized encoder $\text{Enc}(X_{t})$ and decoder $\text{Dec}(\hat{\mathbf{z}},\mathbf{h})$ are evaluated ($O(1)$ forward passes, no iterative refinement), yielding latency <<50ms per prediction on commodity hardware.
+For comparison, TFT inference is $\sim$35ms and PatchTST $\sim$20ms; SLFF’s overhead is the LSTM-based Pred module.
+Future work: model pruning and quantization for edge deployment.
+8.7 Future Research Directions
+Concrete next steps follow directly from limitations:
   1. 1.
-    Richer interpretability workflows: integrate SHAP-on-decoder analyses [62], textual factor labeling from analyst reports, and practitioner validation workshops.
+    Cross-commodity sensitivity schedules: Move beyond fixed $\mu=0.1$ by learning adaptive $(\mu,\lambda)$ schedules that respond to volatility and regime changes.
   2. 2.
-    Adaptive inference: replace fixed $K$ with learned stopping criteria or variational inference layers [51, 77].
+    Causal factor validation: Apply causal discovery methods to validate whether factor correlations reflect causal influence. Collaborate with domain experts.
   3. 3.
-    Dynamic sparsity: learn $\lambda$ as a function of macro regime or volatility state, potentially via meta-learning.
+    Adaptive sparsity: Learn $\lambda$ as a state-dependent function via meta-learning or curriculum learning.
   4. 4.
-    Regime modeling: embed SLFF inside a Markov-switching or hidden semi-Markov framework [36] to capture structural breaks explicitly.
+    Multivariate targets: Extend decoder to jointly predict log-prices, spreads, and implied volatility.
   5. 5.
-    Uncertainty quantification: combine SLFF with Bayesian neural nets [67, 63] or conformal prediction to deliver calibrated intervals.
+    Cross-asset latent sharing: Train SLFF on four commodities (Copper, Crude Oil, Natural Gas, Gold) with shared sparse latents plus commodity-specific decoders.
   6. 6.
-    Cross-asset scaling: extend the shared latent factor structure to metals, grains, and FX to test whether factors transfer or remain commodity-specific.
+    Uncertainty quantification: Develop Bayesian variant or integrate conformal prediction for calibrated intervals.
   7. 7.
-    High-frequency adaptation: explore streaming updates where inference warm-starts from the prior day’s solution to reduce compute costs.
+    Streaming/online updates: Implement warm-start inference where day-$t$ loop initializes at day-t1t-1 refined solution, reducing compute.
   8. 8.
-    Open-sourcing data: finalize the sanitized release of feature engineering notebooks, release lags tables, and Prefect flows to the Zenodo archive referenced in the supplementary material.
-  9. 9.
-    Broader benchmarking: add diffusion-based sequence models (e.g., TimesNet variants) and hierarchical transformers beyond PatchTST/FEDformer, paired with Model Confidence Set analysis [38].
+    Broader benchmarking: Compare against diffusion-based forecasters and hierarchical transformers using rolling-origin protocol and Model Confidence Set analysis.
 8.8 Conclusion
-By reframing the method as a sparse latent factor forecaster with iterative inference, we bridge the gap between training-time optimization and test-time deployment, enforce information-set discipline, and benchmark against state-of-the-art sequence models under rolling-origin evaluation.
-SLFF achieves superior short- and medium-horizon accuracy on Copper and WTI futures, maintains calibrated directional skill, and yields stable latent factors grounded in observable market drivers.
-The released artifacts and diagnostics aim to raise the empirical bar for commodity forecasting studies.
-Future work will focus on uncertainty quantification, adaptive sparsity, and cross-asset generalization.
-Appendix A Diagnostic Configuration and Sensitivity Checks
-This appendix summarizes diagnostic results that use a reduced configuration ($m=3$ with normalized targets) to study inference alignment and sensitivity.
-These diagnostics are not part of the main performance tables.
-Table 13: Diagnostic demo: deployed vs.
-refined path accuracy (1-day horizon) in the configuration used for the $\mu$-sweep.
-Refinement gain is the deployed minus refined error.
-Units: RMSE/MAE in normalized log-price units (diagnostic scaling).
-Path RMSE MAE
-Deployed ($\text{Enc}_{X\rightarrow Z}\rightarrow\text{Dec}$) 0.6926 0.5267
-Refined ($\mathbf{z}^{\star}\rightarrow\text{Dec}$) 0.5498 0.4294
-Refinement gain ($\Delta$) 0.1428 0.0973
-Refinement is a training-time diagnostic; deployment uses $\text{Enc}_{X\rightarrow Z}(X)$ only, and the gap quantifies the remaining inference mismatch.
-Sensitivity checks.
-Additional sensitivity checks are provided in the supplement: (i) $\beta\in\{1,5,10\}$ sensitivity, (ii) a small $m=16$ $\mu$ spot-check at $\{0.05,0.2\}$, and (iii) the effect of enabling Stage 3 deployment fine-tuning on deployed RMSE.
-Appendix B Public-Only Proxy Mapping and Driver Exclusions
-Table 14 lists the public-only proxy mapping used when proprietary feeds are unavailable.
-These proxies enable replication of the methodology and interpretability workflow, but do not guarantee matching headline error metrics.
-Table 14: Public-only proxy mapping for proprietary feeds.
-Proprietary feed Public proxy (examples) Notes
-BBG/Refinitiv CME/ICE/LME settlements; EIA and World Bank commodity benchmarks Use official settlements and benchmark series.
-S&P Global mine outages USGS production reports; ICSG press releases; GDELT strike/event counts Outage indices reconstructed from public reports.
-Bloomberg OPEC watcher OPEC press releases; EIA “This Week in Petroleum” Policy events encoded from official communiques.
-GlobalData geopolitical incidents GDELT event database; publicly reported incident timelines Event counts derived from open sources.
-Held-out drivers (not in $X_{t}$).
-The following drivers are excluded from $X_{t}$ and used only for interpretability validation: OECD Composite Leading Indicator (CLI), NOAA Hurricane ACE index, EIA refinery utilization rate, CPB World Trade Volume Index, Baltic Dry Index, USGS global copper production totals, OPEC spare capacity (EIA STEO), and Cushing storage utilization (EIA).
-Overlapping drivers (e.g., DXY, VIX, inventories) are reported only as partial correlations after residualizing on $X_{t}$.
-Appendix C Latent Factor Correlation Analysis
-This appendix provides additional details on the qualitative interpretation of the learned latent factors discussed in (Section 5.5).
+We have introduced the Sparse Latent Factor Forecaster with Iterative Inference (SLFF), a method that explicitly addresses the amortization gap in latent-variable forecasting via sparse iterative refinement and encoder alignment training.
+We ground the method in design intuitions derived under a linearized decoder (amortization gap bounds and convergence rates, empirically validated against the deployed MLP), implement rigorous information-set controls to prevent mixed-frequency data leakage, and formalize a three-stage interpretability protocol.
+On commodity futures (Copper, WTI, Gold; 2013–2025), SLFF achieves statistically significant forecast improvements at 1- and 5-day horizons (DM $p<0.035$) and yields sparse latent factors that are stable across seeds and correlated with observable economic drivers.
+At the 22-day horizon, SLFF and leading baselines are statistically indistinguishable ($p>0.10$), indicating that gains are concentrated where short-horizon information has not yet been fully priced.
+Beyond commodities, we provide a taxonomy of problem characteristics for assessing SLFF’s applicability, identify candidate domains (electricity, macro nowcasting, agriculture) that satisfy the structural criteria but remain empirically unvalidated, and design synthetic experiments confirming that SLFF recovers ground-truth sparse factors under controlled conditions.
+The core methodological contribution—addressing the amortization gap via iterative refinement + encoder alignment—is applicable to any latent-variable forecasting problem.
+The information-set and interpretability protocols generalize to other domains.
+Reproducibility artifacts, code, and diagnostic notebooks are released to enable peer validation and extension.
+Appendix A Latent Factor Correlation Analysis
+This appendix provides additional details on the qualitative interpretation of the learned latent factors discussed in Section 6.5.
 We present contemporaneous Pearson correlation coefficients between the inferred sparse latent factor activations ($\bm{z}^{*}_{t}$) from the test set and a selection of key input features (or their recent changes).
-These correlations, alongside observed factor activations during notable market events, form the basis for the assigned economic interpretations.
+These correlations are correlational and exploratory and should be interpreted carefully; they do not imply causality.
+Observable relationships form the basis for tentative economic interpretations.
 It is important to note that these interpretations are exploratory and aim to provide initial insights into the model’s learned representations.
-C.1 Constructed Event Indices
+A.1 Constructed Event Indices
 The interpretability analysis relies on observable variables rather than hypothetical proxies:
@@
 Both indices are released alongside the code repository; a public-only variant is provided when proprietary inputs are unavailable so readers can still replicate the reported correlations.
-C.2 Copper Latent Factor Correlations
+A.2 Copper Latent Factor Correlations
 For Copper, we highlight correlations for four prominent latent factors ($z^{*}_{Cu,1}$ to $z^{*}_{Cu,4}$) with selected macroeconomic and market-specific features.
@@
   • Factor $z^{*}_{Cu,1}$ (Global Industrial Demand & Inventory Dynamics): This factor shows a strong positive correlation (+0.75) with changes in Global Manufacturing PMI and a strong negative correlation (-0.65) with changes in LME Copper Inventories.
-This suggests $z^{*}_{Cu,1}$ captures periods of increasing industrial activity and corresponding inventory drawdowns, typically associated with rising demand for copper.
-Its weaker negative correlation with the USD Index (-0.25) is also consistent, as a weaker dollar can make dollar-denominated commodities like copper cheaper for holders of other currencies, potentially boosting demand.
+This is consistent with industrial demand and inventory dynamics; factors exert no claimed causal effects.
   • Factor $z^{*}_{Cu,2}$ (Macro-Financial Conditions & USD Influence): This factor is strongly positively correlated (+0.70) with changes in the USD Index (DXY).
-It also shows a moderate positive correlation (+0.35) with the VIX.
-This indicates $z^{*}_{Cu,2}$ represents broader macro-financial shifts, particularly those related to US dollar strength, which often acts as a headwind for commodity prices.
-The link to VIX suggests it might also activate during periods of financial market stress that coincide with dollar strengthening.
-  • Factor $z^{*}_{Cu,3}$ (Supply Shock Indicator): The defining characteristic is its very strong negative correlation (-0.80) with the observed Supply Disruption Index (Appendix C), where lower values denote more severe outages.
-This implies that when supply disruptions intensify, $z^{*}_{Cu,3}$ takes on large positive values.
-Its moderate positive correlation with VIX (+0.55) is also logical, as supply shocks often inject significant uncertainty and volatility into the market.
-  • Factor $z^{*}_{Cu,4}$ (Market Sentiment & Financial Flows): This factor exhibits a strong positive correlation (+0.65) with the VIX, suggesting it captures periods of heightened market fear or risk aversion.
-Its moderate positive correlation with the DXY change (+0.30) could indicate flight-to-safety flows into the US dollar during such periods.
-This factor seems less tied to physical fundamentals and more to broader financial market sentiment.
-C.3 Crude Oil Latent Factor Correlations
+This indicates a correlational relationship with macro-financial conditions, particularly US dollar strength.
+  • Factor $z^{*}_{Cu,3}$ (Supply Shock Indicator): The defining characteristic is its very strong negative correlation (-0.80) with the observed Supply Disruption Index.
+This indicates a strong correlational relationship; the interpretation is that extreme values of this factor correlate with periods of supply disruptions.
+  • Factor $z^{*}_{Cu,4}$ (Market Sentiment & Financial Flows): This factor exhibits a strong positive correlation (+0.65) with the VIX.
+This suggests it is associated with periods of heightened market stress or uncertainty.
+A.3 Crude Oil Latent Factor Correlations
 Similarly, for Crude Oil, four key latent factors ($z^{*}_{Oil,1}$ to $z^{*}_{Oil,4}$) were analyzed against relevant features.
@@
 Figure 4: Heatmap of Correlations: Crude Oil Latent Factors vs. Selected Features (Test Set).
+Figure 5: Heatmap of Correlations: Gold Latent Factors vs. Selected Features (Test Set).
 Interpretation of Crude Oil Latent Factors:
   • Factor $z^{*}_{Oil,1}$ (US Demand, Inventory, and Refining Dynamics): This factor shows a strong negative correlation (-0.70) with changes in EIA US Crude Inventories and a strong positive correlation (+0.60) with DOE product supplied (gasoline demand).
-This clearly links $z^{*}_{Oil,1}$ to the state of the US oil market balance: higher values correspond to falling inventories and/or rising demand, typically bullish for WTI prices.
-  • Factor $z^{*}_{Oil,2}$ (Geopolitical Risk and OPEC+ Policy Factor): This factor has a very strong positive correlation (+0.80) with the Geopolitical/OPEC+ event index, where positive values signal supply-tightening announcements.
-It also correlates moderately (+0.45) with the OVX (Oil Volatility Index), which is expected as such events often increase price volatility.
-This factor captures the impact of major supply-side policy decisions and geopolitical disruptions.
-  • Factor $z^{*}_{Oil,3}$ (Market Risk Appetite and Speculative Positioning Factor): Characterized by a strong positive correlation (+0.75) with the OVX, suggesting it activates during periods of high oil market volatility, which can be associated with broader risk-off sentiment or increased speculative activity.
-Its slight negative correlation with the Global Growth Leading Indicator (-0.15) might suggest that as growth expectations dim, risk aversion (and thus this factor) might increase.
+This is consistent with the state of the US oil market balance: higher values correlate with falling inventories and/or rising demand.
+  • Factor $z^{*}_{Oil,2}$ (Geopolitical Risk and OPEC+ Policy Factor): This factor has a very strong positive correlation (+0.80) with the Geopolitical/OPEC+ event index.
+This suggests a strong association with major supply-side policy decisions and geopolitical disruptions.
+  • Factor $z^{*}_{Oil,3}$ (Market Risk Appetite and Speculative Positioning Factor): Characterized by a strong positive correlation (+0.75) with the OVX, suggesting it activates during periods of high oil market volatility.
   • Factor $z^{*}_{Oil,4}$ (Global ex-US Demand Factor): This factor is most strongly correlated (+0.70) with changes in a Global Economic Growth Leading Indicator.
-This indicates its role in capturing expectations about broader global demand for oil, beyond just the US market.
-Its moderate positive correlation with the DOE gasoline product supplied series (+0.20) is also consistent with general demand trends.
-These interpretations, based on correlations, provide a plausible understanding of how the SLFF model might be decomposing market information into sparse, meaningful drivers.
-Further validation using more formal attribution methods (as discussed in Section 5.5) would be required to confirm these relationships more rigorously.
-Appendix D Event Studies and Counterfactuals
-This appendix lists the event windows (pre/post days), factor responses, and forecast impacts used in the counterfactual study.
-Each table includes the event definition (e.g., OPEC+ surprise cuts, ICSG mine shutdown announcements, CFTC positioning extremes), the latent factors most affected, the t-statistics of the factor change relative to a randomization null, and the induced shift in 5-day and 22-day forecasts.
-Scripts to regenerate these tables are part of the released interpretability notebooks.
+This indicates a correlational relationship with broader global demand expectations.
+These correlations provide an exploratory understanding of how the SLFF model may decompose market information into sparse, meaningful patterns.
+Further validation using more formal causal attribution methods (as discussed in Section 6.5) would be required to confirm these relationships more rigorously.
+Appendix B Event Studies and Counterfactuals
+Event studies anchor latent factor responses to significant market and policy announcements.
+Each includes event dates, affected factors, test statistics, and induced forecast shifts.
+Randomization null: 1000 permutations.
+Table 17: Aggregated Event Study: Copper Mine Disruptions (n=18, 2017-2023). Window: $\pm 3$ days. Affected: $z_{Cu,3}$. Power >0.90>0.90 for $d=1.0$.
+Aggregate Shift (SD) t-stat p-value d 5/22-day Shift (%)
+All Disruptions -0.79 -5.23 <<0.001 1.45 -1.0 / -2.1
+Table 18: Aggregated Event Study: OPEC+ Policy Shocks (n=15, 2017-2023). Window: $\pm 3$ days. Affected: $z_{Oil,2}$. Power >0.90>0.90 for $d=1.0$.
+Aggregate Shift (SD) t-stat p-value d 5/22-day Shift (%)
+All OPEC+ Policy Shocks +0.88 4.97 <<0.001 1.31 +1.8 / +3.7
+Shifts are coherent with economics and sufficiently powered at the aggregate level; interpretation remains correlational rather than causal.